### v1

# Frequency Is What You Need: Word-frequency Masking Benefits Vision-Language Model Pre-training

Abstract
Vision Language Models (VLMs) can be trained more efficiently if training sets can be reduced in size. Recent work has shown the benefits of masking text during VLM training using a variety of approaches: truncation, random masking, block masking and syntax masking. In this paper, we show that the best masking strategy changes over training epochs and that, given sufficient training epochs, word frequency information is what you need to achieve the best performance. Experiments on a large range of data sets demonstrate the advantages of our approach, called Contrastive Language-Image Pre-training with word Frequency Masking (CLIPF). The benefits are particularly evident as the number of input tokens decreases. We analyze the impact of CLIPF vs. other masking approaches on word frequency balance and discuss the apparently critical contribution of CLIPF in maintaining word frequency balance across POS categories.
Keywords Vision-Language Model  $\cdot$ Multimodal Data  $\cdot$ Word Frequency Masking.
1 Introduction
Vision-Language Models (VLMs) learn embeddings that represent both visual and textual modalities simultaneously, and are known for the ease with which their strong performance transfers from one domain to another [1, 2, 3, 4, 5, 6, 7, 8]. CLIP (Contrastive Language-Image Pre-training) [3, 9, 10] is a VLM that is trained on a very large number of image-text pairs collected online, where the text is considered a “natural language” annotation for the image.
This paper contributes to the recent trend of text-masking approaches that improve VLMs by removing words from the text data during training, speeding up training in the process [11, 12]. We introduce CLIPF (Contrastive Language-Image Pre-training with word Frequency Masking), which selects words to be masked on the basis of word frequency. We carry out a comparison with CLIPA masking [12], which selects words on the basis of four strategies that do not use frequency information: truncation, random word selection, random block selection, and part-of-speech (POS) information and is currently considered the state-of-the-art text-masking.
The comparison of the learning curves of our CLIPF with CLIPA masking in Fig. 1 illustrates two key contributions of this paper. First, we demonstrate that the number of training epochs is crucial to which masking strategy yields the highest performance. The finding is an important improvement on the state of the art of text masking for CLIP training, represented by [12]. In [12], CLIP is trained on a limited number of epochs and syntax masking is found to outperform truncation, random, and block masking. In contrast, we show that random and block masking outperform syntax masking after a sufficient number of epochs. Second, we show that when masking text for CLIP training, frequency is what you need. As evident in Fig. 1, our frequency-based CLIPF approach outperforms all CLIPA masking approaches with sufficient epochs of training, leading it to be the best performing text masking approach.
Table 1: Illustration of different text masking methods and words masking probabilities. Bold percentages or check marks provide an example of the words that would be retained by each text masking strategy. For the Truncation and Syntax methods, four words per text are retained as the input. For the Random, Block, and CLIPF methods, we present the word masking probabilities. The second row shows the part-of-speech (POS) category of each word.
Models Methods Words
a small dog sleeping on a couch next to a remote .
POS Other Adjective Noun Verb Other Other Noun Other Other Other Noun Other
CLIPA Truncation
Random 50% 50% 50% 50% 50% 50% 50% 50% 50% 50% 50% 50%
Block 50% 50% 50% 50% 50% 50% 50% 50% 50% 50% 50% 50%
Syntax
CLIPF Frequency 99.37% 96.28% 96.13% 88.35% 98.96% 99.37% 87.03% 93.69% 98.89% 99.37% 81.11% 99.37%
Figure 1: Zero-shot classification accuracy on ImageNet-1K over training epochs for CLIPF and CLIPA strategies. The text masking ratio is 75%. The backbone of the image encoder is ViT-B/16. The model is pre-trained on CC12M for 30 epochs. We use image masking (75%) to speed up training. All approaches are subsequently fine-tuned with one epoch on the full, unmasked data.
The history of text masking in language model training dates to [13], where subsampling of frequent words in the training text was used to improve the performance of skip-gram language models. The study of text masking for VLMs got off to a slow start, since initially [8] reported that text masking, although it promotes efficiency, deteriorated performance. Subsequently, [12] expanded the types of masking studied and demonstrated that masking can indeed improve performance. Simultaneously, frequency-masking was proposed at a workshop [11], but fell short of the potential of frequency realized by CLIPF due to the use of a cut-off threshold. To our knowledge, our work reported in this paper is the first to attempt to discover the reasons that certain forms of text masking outperform others.
An initial illustration of the differences between task-masking strategies is provided in Table 1. Truncation consistently retains the beginning of the text. With the random (random word selection) and block (random block selection) methods, each word has an equal probability of being masked, with block masking maintaining word neighborhoods. Syntax masking prioritizes the retention of nouns, followed by other part-of-speech (POS) categories. Masking based on frequency (CLIPF) has the potential to address the shortcomings of the other approaches. Unlike truncation, it does not prioritize words that tend to occur at the beginning of the sentence. In contrast to random and block masking, it represses words with the POS category “Other”, which are often stop words and for this reason high frequency. However, it does not give an overly high priority to nouns and adjectives, as with syntax masking.
In sum, the contributions of this paper are as follows:
  • We point out, to our knowledge for the first time, the importance of the number of training epochs for comparing text masking approaches.
  • We introduce CLIPF, a frequency-based masking approach that outperforms previously proposed approaches.
  • We carry out an analysis that provides insight into why CLIPF outperforms other approaches.
Our work is reproducible because it uses training data sets that are openly available and small enough to allow researchers with limited computational resources to repeat our experiments and because we release our code [^1] https://github.com/MingliangLiang3/CLIPF
2 Related work
2.1 Vision-Language Models
Vision-language models (VLMs), such as CLIP, ALIGN, LiT, and FLIP [3, 4, 10, 8], learn image embeddings from large-scale natural language supervision using contrastive learning [14] to grants the models remarkable transferability. Because CLIP does not release its pre-training dataset, OpenCLIP pre-trained models on several large-scale public datasets [15, 16, 17, 18] to reproduce CLIP. This implementation adheres to the scaling law observed in Neural Language Models [19] and Vision-Language Models [3], where loss decreases according to a power-law relationship with model size, dataset size, and computational resources utilized during training. While large language models (LLMs) have advanced rapidly, we focus on vision-language models (VLMs) like CLIP, as they are essential tools for bridging visual and textual modalities in current research. Although CLIP and OpenCLIP are pre-trained on very large-scale datasets, recent works like DeCLIP [20] and SLIP [5] demonstrate that Vision-language Models (VLMs) can be trained efficiently with smaller datasets, enabling researchers with limited computational resources to contribute. Moreover, masking image patches or reducing image resolution during VLM pre-training also substantially decreases computational costs while maintaining or enhancing performance, as shown by techniques like FLIP, RECLIP, CLIPA [8, 21, 12].
2.2 Text masking for VLM training
Masking text presents another viable option to speed up the pre-training of VLMs, and is particularly effective alongside strategies that employ a high ratio of image patch masking. FLIP further explored text masking techniques, namely, random and truncation masking, to enhance efficiency [8]. While random text masking involves masking text arbitrarily, truncation masking in FLIP strategically eliminates padding tokens and retrains the words that preceded them.
Subsequently, CLIPA [12] introduces block masking and syntax masking. As already mentioned in Section 1, CLIPA [12] shows text masking during training can improve the performance of a VLM. CLIPA syntax masking prioritizes the retention of nouns, followed by adjectives and then other words. In [12], syntax masking outperforms the truncation, random, and block approaches. However, [12] does not vary the number of training epochs. An important contrast with our work, as previously mentioned, is our attention to the number of training epochs. Further, CLIPF, as we will see, counters the danger of giving overly high priority to nouns, leading to overfitting, and also avoids overly suppressing words of the POS category “Other”.
An initial effort to exploit frequency information for text masking during VLM training was made by SW-CLIP [11]. Unfortunately, SW-CLIPF introduces a frequency cutoff threshold, which reduces performance, as we demonstrate later in this paper. In order to exploit frequency information effectively each training text needs to fill as many available input tokens as possible, as is done in CLIPF.
3 Method
In this section, we describe various text masking strategies, including our CLIPF method. All these strategies accelerate pre-training by masking specific words in the text.
3.1 Text Masking
Consider a text $T_{i}$ consisting of words $\{w_{1},w_{2},w_{3},\ldots,w_{n}\}$, where $n$ represents the total number of words in $T_{i}$. Text masking involves selecting a subset of these words to create a new version of the text, denoted as
$T_{new\_i}=f(T_{i}\{w_{1},w_{2},w_{3},\ldots,w_{n}\}).$ (1)
This function $f$ determines which words are masked to get the modified text $T_{new\_i}$. And, we define the number of input tokens as $k$. We will sequentially introduce our masking strategies: random, truncation, block, syntax, and frequency. It is important to note that the number of padding tokens remains consistent across all these strategies.
Random Masking: In the random text masking strategy, the function $f$ operates to randomly select words from the text $T_{i}$. Each word $w_{i}$ has an equal probability of being masked, independently of its position or frequency in the text.
Truncation Masking: In truncation masking, the function $f$ reduces the text by masking words from the end part of the text. The number of text tokens retained depends on the input length of the text encoder.
Block Masking: Block masking is similar to truncation masking but starts at a random index within the text. If the text tokens are longer than the input tokens, we choose a random starting index $j$ that is less than the length of the text tokens minus the length of the input tokens $k$.
Syntax Masking: To implement syntax masking [12], each word $w_{i}$ in the text $T_{i}$ is first tagged using a natural language processing tool like NLTK [22] to identify its part of speech. The words are then sorted according to their syntax role. This ordering strategy $O(w_{i})$ reflects the importance of nouns, adjectives, and verbs in the text. Words are masked based on their assigned order, with priority given to retaining nouns, followed by adjectives, verbs, and then other parts of speech. This method ensures that the core semantic information of the text is less likely to be masked, especially when learning the relationship between the image and the text, as nouns are more relevant to the object of the image.
3.2 Word Frequency Masking
As in previous work, subsampling of frequent words was introduced by [13] in the natural language processing (NLP) domain to accelerate the learning process for word representations. The goals are to accelerate training and to address the imbalance between frequent and infrequent words in the training corpus. In our work, we use the formula (Formula 2) introduced in [13] to determine the probability of masking a word during training. Thus, the complete formula to calculate the probability $P(w_{i})$ that accounts for all scenarios is:
$P(w_{i})=\begin{cases}1&\text{if}N<5\\ 0&\text{if}f(w_{i})<t\\ 1-\sqrt{\frac{t}{f(w_{i})}}&\text{otherwise}\end{cases}$ (2)
Here, the masking probability of the word $w_{i}$ is $P(w_{i})$, $t$ is the threshold, and $f(w_{i})$ is the frequency of the word. Additionally, if the total number of words ($N$) is fewer than 5, deemed too infrequent for effective learning, the probability that the word will be masked is 1.
Following FLIP [8], we fine-tune the model with an additional epoch to eliminate the distribution gap between the masking and unmasking.
4 Experimental Setup
Our implementation follows CLIP [3], OpenCLIP [9], FLIP [8] and CLIPA [12]. In this section, we present the details of the training and testing.
4.1 Implementation
Dataset We pre-train the models on Conceptual Captions 3M (CC3M) [23] and Conceptual 12M (CC12M) [15]. These datasets were selected due to their diverse range of real-world concepts and scenes. Popular VLMs such as OpenCLIP, DeCLIP, and SLIP [9, 20, 5] also employ these datasets for training, leveraging their extensive and diverse content. Due to the expiration and inaccessibility of some URLs in the training set, we have downloaded 2.7 million image-text pairs for CC3M and 9.3 million image-text pairs from CC12M. The characteristics of the dataset are presented in Table 2. The CC12M dataset is 3.42 times larger than the CC3M dataset, with captions that are on average 2.15 times longer. Since the number of input tokens is fixed, the impact of different text masking ratios varies across datasets due to differences related to caption length.
Table 2: Dataset Statistics for Pre-training datasets. Caption length refers to the number of words in the text.
Dataset Samples Total words Caption length
CC3M 2721877 28022082 10.30 $\pm$ 4.72
CC12M 9295444 205716854 22.15 $\pm$ 17.20
Architecture For the image encoder, we use ViT-B/16 [24] architectures with global average pooling and learnable positional embeddings. The input image size is 224. For the text encoder, we use a Transformer-based model [25] and byte-pair encoding with a 49K token vocabulary. We chose these encoders because they are the largest encoders that we can use given our resources. When using full text for training, the maximum context length is 32. Zero-padding is applied to input text that is shorter than the maximum token length of the model.
We trained the models using 8 RTX A5000 GPUs with the same settings to ensure consistent conditions across all models. This allows us to effectively measure and compare the performance of each model in processing text tokens during training. We experiment with different text token input sizes, namely, 32, 16, 8, 6, and 4 text tokens. As the input size gets smaller, we can increase the batch size, maximizing computational memory usage. We used batch sizes of 664, 832, 896, 928, and 944, respectively. Note that CLIP (and therefore CLIPA and CLIPF) is trained using contrastive learning, which benefits from larger batch sizes. Note also that although we used different text masking with the same settings and batch sizes, syntax masking was slower than the other text masking strategies when conducting the experiments because POS processing for each word is time-consuming.
We pre-train the model using the InfoNCE loss [14] with a learnable temperature parameter $\tau$ [26, 3]. To classify images, we calculate the cosine similarity between the image and text embeddings.
Table 3: Details of the pre-training and fine-tuning setups.
Config Pre-training Fine-tuning
optimizer AdamW [27] AdamW [27]
learning rate 1e-3 1e-5
weight decay 0.2 0.2
optimizer momentum $\beta_{1},\beta_{2}=0.9,0.98$ [28] $\beta_{1},\beta_{2}=0.9,0.98$ [28]
learning rate schedule cosine decay [29] cosine decay [29]
warmup steps 10k 10%
epoch 30 1
$t$ $10^{-6}$
$\tau$ 0.07
numerical precision amp amp
augmentation RandomResizedCrop RandomResizedCrop
Table 4: During pre-training, each model processes a certain number of image and text tokens. The first and second columns of the table display the image and text masking ratios, respectively. The third and fourth columns show the number of image and text tokens processed after masking. The fifth column presents the total number of tokens (image + text) after masking. The last column provides the percentage of total tokens relative to the unmasked case.
Image masking Text masking Image tokens Text tokens Total Percentage
0% 0.00% 196 32 228 100%
75% 0.00% 49 32 81 35.53%
50.00% 49 16 65 28.51%
75.00% 49 8 57 25.00%
81.25% 49 6 55 24.12%
87.50% 49 4 53 23.25%
Training We follow the setup of OpenCLIP [9], FLIP [8] and CLIPA to pre-train our model. Following OpenCLIP [9], we pre-train our model for 30 epochs on the CC3M and CC12M datasets. Details of the pre-training configuration are given in Table  3.
Similar to FLIP, to speed up training, we apply 75% image masking to the image encoder as the baseline model. As a result, the speedup is about 4$\times$ compared to training without image masking, while the reduction in the zero-shot performance of ImageNet-1K classification remains within reasonable bounds.
During text masking, we reduce the number of tokens from 32 to 16, 8, 6, and 4. To measure the training speed of CLIPF, CLIP, and CLIPA, we compare the number of text tokens processed by each model. As shown in Table 4, the number of text tokens processed by each model during pre-training is calculated based on different image and text masking ratios. When we pre-trained the model using image masking, the total number of tokens is 81 and the percentage of text tokens is 39.5%. When we apply 50% text masking, the total number of tokens is 65. Compared to training without text masking, this results in a speed increase of approximately 20%. Moreover, when we apply 87.5% text masking, the total number of tokens is 53, resulting in a speed increase of approximately 34% compared to training without text masking.
Fine-tuning Following FLIP, and CLIPA, in order to bridge the gap between pre-training and evaluation, we fine-tuned the model without images and text masking. The details of the fine-tuning configuration are provided in Table 3.
Table 5: Comparison of CLIPF, CLIP, and CLIPA for zero-shot classification on ImageNet-1K. reported as top-1 accuracy. All models use a ViT-B/16 image encoder backbone pre-trained on CC3M and CC12M for 30 epochs with 75% image masking for speedup (pre-train) and fine tuned for an additional epoch on all data (fine-tune). “Image tokens” and “text tokens” reflect the amount of image and text processing during pre-training.
Models Masking Image Tokens Text Tokens CC3M CC12M
pre-train fine-tune pre-train fine-tune
CLIP 197 32 18.6 36.6
FLIP 49 32 14.1 14.2 32.0 33.7
CLIPA truncation 49 16 13.8 13.8 32.8 32.8
random 13.9 13.9 33.7 34.3
block 13.9 13.9 34.3 34.8
syntax 13.3 12.8 32.2 34.4
CLIPF frequency 14.0 14.0 35.5 36.0
CLIPA truncation 49 8 10.8 12.0 25.4 28.4
random 17.6 17.4 34.5 36.9
block 16.2 16.6 35.5 37.9
syntax 17.2 17.5 28.5 35.0
CLIPF frequency 16.8 17.0 36.6 39.3
CLIPA truncation 49 6 8.4 9.4 15.3 23.2
random 12.8 17.9 26.9 34.6
block 12.9 17.0 28.6 35.9
syntax 12.2 15.7 25.2 32.6
CLIPF frequency 14.4 18.2 30.3 37.8
CLIPA truncation 49 4 3.8 8.2 5.3 19.8
random 5.4 14.6 14.0 27.1
block 7.5 14.5 18.7 26.6
syntax 8.9 13.0 14.2 24.6
CLIPF frequency 10.9 16.0 17.0 30.9
4.2 Evaluation
Evaluation setup Following the CLIP [3], FLIP [8], and CLIPA [12] methods, several classification benchmarks were used to evaluate the zero-shot capability of CLIPF. Among these benchmarks is ImageNet-1K, a widely recognized dataset in computer vision. It is frequently used for image classification and VLMs tasks and comprises 50K validation samples across 1K different classes. We filled each class into the templates provided by CLIP [3] to calculate the average of the text embeddings. We use the same evaluation settings as CLIP [3] to evaluate the other downstream datasets.
5 Experimental Results
5.1 Zero-short performance on ImageNet
Table 5 presents results for zero-shot classification on ImageNet-1K for models trained on both the CC3M and CC12M data. The experiments compare different masking approaches and different masking ratios, indicated in the table with the number of input text tokens (Text Tokens column). Models are pre-trained for 30 epochs (pre-train column) and then fine tuned for one epoch (fine-tune column) with the full, unmasked data. Note that since the average text length in CC3M is less than 16 tokens, the benefits of fine-tuning a model pre-trained with 16 text tokens are limited.
We see that CLIPF outperforms the other CLIPA matching approaches in nearly all cases across text masking ratios (input sizes) as reflected by the bolded accuracy scores. In general, random, syntax, and block strategies exhibit similar performance; however, truncation performs worse. Contrary to the conclusion of the original CLIPA work [12], syntax masking does not yield the best performance. As mentioned in Section 1, we observe that the relative performance of different text masking approaches depends on the number of training epochs. In [12], CLIPA only pre-trained the models for 6.4 epochs. We discuss the learning curve in Fig.1 further in Section 6.2.
It is particularly interesting to observe from the results in Table 5 the fact that models trained with text masking (CLIPA and CLIPF) are competitive with, or superior to, models trained on the original, unmasked text data (CLIP, FLIP). Note that at a text masking ratio of 75% (corresponding to 8 tokens), the masking approaches are 1.3 $\times$ faster than FLIP and 4.0 $\times$ faster than CLIP.
We first consider the comparison with CLIP. On CC3M, CLIPF approaches the performance of the original CLIP approach with a 81% text masking ratio (4 input tokens). On CC12M, CLIPF outperforms the original CLIP (no masking) with a 75% masking ratio (8 input tokens). Fine-tuned CLIPF with a 75% masking ratio achieves an accuracy of 39.3% compared to 36.6% for the original CLIP trained on the full data.
Next we consider the comparison with FLIP, which is the same as CLIP except that it uses a 75% image masking ratio (49 image tokens). On CC3M, fine-tuned CLIPF is competitive with or outperforms fine-tuned FLIP at all masking ratios. On CC12M, fine-tuned CLIPF outperforms fine-tuned FLIP, except for the highest masking ratio (87.5% corresponding to 4 input tokens).
Finally, we consider the pattern of the text masking approaches as the masking ratio increases, i.e., as we decrease the input size from 16 to 4 tokens. Here, we focus on the results of the fine-tuned model. Moving from 16 to 8 input tokens benefits most approaches. Recall that the reduction of the number of input tokens allows for an increase of the batch size. Moving from 8 to 6 to 4 tokens, we see that CLIPF maintains its performance surprisingly well. As the masking ratio becomes higher, the margin by which CLIPF outperforms other masking strategies increases. In short, CLIPF is better positioned to deliver benefit as the number of input tokens decreases, reflecting the strength of frequency information in identifying important words to retain during masking.
5.2 Zero-shot performance on different data sets
Table 6: Zero-shot robustness evaluation. Comparison of the zero-shot accuracy performance of CLIP, FLIP, CLIPA, and CLIPF on various datasets. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
Models Masking Image Tokens Text Tokens ViT-B/16
IN-A IN-O IN-R IN-S IN-V2 ON
CLIP 197 32 8.97 37.85 49.11 25.70 31.48 24.20
FLIP 49 32 6.90 40.30 36.90 17.60 28.70 18.10
CLIPA truncation 49 16 7.10 37.70 36.40 17.50 27.40 17.20
random 8.05 38.60 40.00 20.34 29.61 20.84
block 8.05 39.60 39.17 19.31 29.70 19.72
syntax 8.00 40.70 39.41 18.61 29.17 19.96
CLIPF frequency 8.28 39.10 40.86 20.48 30.78 21.15
CLIPA truncation 49 8 6.71 34.05 37.18 15.99 24.49 15.98
random 9.16 38.15 43.35 22.53 31.24 22.17
block 9.05 39.95 43.32 22.58 32.42 21.58
syntax 8.50 38.40 39.90 19.50 29.80 19.50
CLIPF frequency 9.56 39.90 45.76 23.62 34.16 21.59
CLIPA truncation 49 6 6.70 26.60 33.20 13.30 20.20 14.80
random 8.10 36.10 41.80 21.10 30.20 19.80
block 8.32 38.20 42.99 21.73 30.82 19.76
syntax 7.40 34.10 39.60 18.50 28.60 19.80
CLIPF frequency 8.69 38.55 45.09 23.45 32.59 20.04
CLIPA truncation 49 4 5.30 21.70 25.40 9.00 17.10 11.20
random 6.10 30.30 33.80 14.90 23.50 15.30
block 6.08 30.55 34.69 15.54 22.99 15.10
syntax 6.50 29.10 32.40 13.30 22.00 16.00
CLIPF frequency 7.05 32.40 39.90 18.94 26.72 17.14
Following the methodology outlined in CLIP and FLIP [3, 8], we evaluated various masking strategies across data sets to identify their robustness. Table 6 presents results for zero-shot classification for six commonly used data sets. The performance of these strategies on multiple datasets is similar to the behavior observed on the ImageNet-1K dataset. Across different text masking ratios and across different data sets, CLIPF outperforms the other masking strategies in nearly all cases. Reducing the text tokens in CLIPF from 16 to 8 results in substantial performance improvements compared to the other text masking approaches across almost all datasets: ImageNet-A [30] by 1.28%, ImageNet-O [31] by 0.80%, ImageNet-R[32] by 4.9% , ImageNet-Sketch [33] by 3.14%, ImageNet-V2 [34] by 3.38%, and ObjectNet [35] by 0.44%.
Again, we see that the benefits of CLIPF are particularly evident as the masking number of input tokens decreases. In other words, we observe that the gap in performance widen between CLIP and the other strategies as the masking ratio increases.
We also observe that CLIPF is able to improve over the FLIP baseline without masking in a majority of cases. However, despite being the strongest text-matching strategy, the results reveal that it is challenging for CLIPF to improve over the original CLIP. In only a handful of cases does it outperform the original CLIP. However, it is important to recall that training with full images, rather than using image masking, can also be combined with text masking. We did not test this setting here for computational reasons, but depending on the computational resources in an application setting it could be relevant.
Table 7: Zero-shot Image-Text Retrieval. We evaluated CLIP, FLIP, CLIPA, and CLIPF image-text retrieval performance on COCO and Flickr30k datasets. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
Models Masking Text Tokens Text Retrieval Image Retrieval
Flickr30k COCO Flickr30k COCO
R@1 R@1 R@1 R@1
CLIP 32 62.62 35.54 45.42 24.22
FLIP 32 54.73 29.34 38.62 20.88
CLIPA truncation 16 44.67 25.54 34.99 19.64
random 16 58.48 32.36 43.61 22.81
block 16 56.51 30.82 44.06 22.66
syntax 16 54.54 29.60 41.16 21.40
CLIPF frequency 16 57.89 31.52 42.72 22.57
CLIPA truncation 8 30.47 16.92 23.96 12.78
random 8 58.58 36.24 43.79 23.16
block 8 60.06 35.88 45.98 24.65
syntax 8 50.30 29.64 38.46 20.28
CLIPF frequency 8 60.36 35.28 44.81 23.66
CLIPA truncation 6 32.25 16.18 22.39 11.93
random 6 55.23 32.58 42.47 21.34
block 6 54.44 33.12 41.68 21.73
syntax 6 46.15 26.78 34.08 17.99
CLIPF frequency 6 56.02 32.32 41.05 21.28
CLIPA truncation 4 30.97 17.44 21.09 11.07
random 4 41.62 24.14 30.26 15.63
block 4 40.83 23.56 30.77 14.98
syntax 4 38.66 21.68 26.08 14.14
CLIPF frequency 4 45.07 25.60 31.72 15.96
5.3 Zero-shot Image-Text Retrieval
Table 7 shows image/text retrieval results on Flickr30k [36] and COCO [37]. These results demonstrate the ability of text masking to improve the performance of a VLM. The performance improvement is particularly notable in comparison to FLIP, which like our CLIPA and CLIPF uses a 75% image-masking ration. When using 16 text tokens, all strategies except the truncation method outperform FLIP which is pre-trained on the full text. The performance is also quite strong in comparison to CLIP, which uses no masking. When using 16 text tokens, text masking strategies, except for truncation, are able to approach and in a few cases beat the accuracy of CLIP.
In contrast to the previous experiments, with image-text retrieval, we observe that CLIPF does not clearly dominate the other text masking strategies. The CLIPA random, CLIPA block, and CLIPF (frequency) strategies exhibit comparable performance in image and text retrieval across both datasets, but CLIPF does not consistently outperform the other masking strategies. The performance of the syntax strategy is similar to that of FLIP but is lower than that seen with the random, block, and CLIPF. The notable exception is the performance of CLIPF in the case that the text masking ratio is very high, i.e., only 4 input text tokens are used. Here, the benefit of CLIPF’s frequency-based approach is clearly evident.
6 Analysis and Further Experiments
6.1 Word Distribution Analysis
Figure 2: The figure illustrates the distribution of top-50 words in the text before and after applying various text masking strategies, we set the text length after text masking to 6. The x-axis represents the word index which is sorted by counts of the original data, and the y-axis shows the word frequency. The dataset used is CC12M and the value of $t$ of formular 2 is set to $10^{-6}$. We remove special characters from the vocabulary.
In this section, we provide an analysis of how different text masking strategies affect the distribution of words By looking at the differences in distribution we can gain insight into why some text-masking strategies work well, and some less well, and, in particular, why CLIPF delivers the best performance.
Fig. 2 shows the distribution of words in the text before and after applying various text masking strategies. We see that in general there is quite a large difference between the frequency of the highest-frequency words and the frequency of words with lower frequency. Such a long-tail distribution is characteristic of words occurring in text, which is known to be Zipf distributed. However, we also observe differences between the different text masking approaches.
The truncation strategy leads to frequent words being more frequent than they are with the other strategies. This effect might seem surprising at first, but can be explained if we consider that texts typically start with the same words, e.g., “A person is…” We conjecture that truncation does not perform well as a text-masking strategy because the VLM overfits to these words.
The CLIPA random and block strategies exhibit nearly identical distributions, since each word has a 50% probability of being masked in both cases. They lie somewhat below the CLIPA truncation line, and we describe them as being more “balanced” not in the sense that the distributions are flat, but rather that the differences between the frequencies of very high-frequency words and the of the rest of the words is less extreme. Even more balanced is the CLIPF distribution, which lies below CLIPA random and block.
The distribution of the CLIPA syntax strategy is quite different. Because this strategy retains more nouns, the articles “the” and “a” or less frequent, but the nouns create spikes. The rationale for prioritizing nouns is convincing. Nouns are closely related to the depicted content of images and for this reason are considered to be more helpful. Indeed, the helpfulness of nouns for the VLM is reflected in the fact that Syntax masking learns faster than other text masking strategies. In Fig. 1 we saw that syntax masking outperforms the other masking approaches during the early epochs and CLIPA [12] found that syntax masking was the best approach, while testing on a limited number of epochs. However, we believe that retaining nouns can have two adverse effects. First, the model can overfit to certain high frequency nouns. In Fig. 2, we see that CLIPA syntax retains more occurrences of “person,” “illustration,” and “stock” than the other strategies. Second, retaining nouns means necessarily dropping words from other part-of-speech (POS) categories.
To dive deeper into the importance of the distribution of words over POS categories, we provide statistics for each of the different text masking strategies in Table 8. Consulting the column ‘Other’ we see and interesting pattern. CLIPA syntax masking changes the proportion of words in the “Other” category radically with respect to the data before masking, but CLIPA truncation, random and block do not. CLIPF appears to provide a happy medium. CLIPF retains more nouns than CLIPA truncation, block, and syntax, but not so many that the “Other” category becomes poorly represented. We believe that it is this balance that this balance is an important factor in CLIPF’s strong performance.
Table 8: To examine the distribution of words belonging to different POS categories, we calculate the percentage of words in the training data belonging to each category in before and after applying text masking. The dataset is CC12M, and we retain 6 words for each text after applying text masking.
Masking NN (%) JJ (%) VB (%) OTHER (%)
Before masking 50.30% 4.98% 5.18% 39.55%
Truncation 51.72% 5.17% 5.91% 37.20%
Random 51.21% 4.82% 5.04% 38.93%
Block 49.37% 4.74% 5.24% 40.66%
Syntax 88.53% 2.69% 2.92% 5.87%
CLIPF 60.43% 5.06% 6.20% 28.24%
6.2 Learning Curve
Next, we return to consideration of the learning curve that compares CLIPF with the other text-masking strategies, which is shown in Fig. 1. As already mentioned, syntax is a strong performer in the earlier epochs, until epoch 20 remains the best of the CLIPA approaches. We conjecture that CLIPA block and CLIPA random pull ahead of syntax because they retain “Other” words (i.e., words that are not nouns, adjectives or verbs) and do not overprioritize nouns.
The importance of “Other” words is surprising because the category includes many words, which like articles, are said to lack semantic content and have no direct connection to images. In fact, in many NLP applications, articles and other “stop words” are simply removed. We believe that the transformer that CLIP uses to embed text is actually able to exploit the context of content words, particularly words in “Other” category, given enough epochs. CLIPA block masking retains this context a little better than random masking. Surprisingly, the frequency masking of CLIPF also retains this context. Under CLIPA syntax masking a phrase like “light in the house” would always be reduced to “light house”, but with CLIPF frequency masking the reduction would not happen in every case. In short, CLIPF achieves a productive balance in POS categories.
6.3 Comparison with SW-CLIP
In this section, we provide a comparison of CLIPF with SW-CLIP [11], which also uses frequency-based sampling but imposes a threshold on the frequency score. The effect of this threshold is that input tokens will go unused if not enough words in a training text surpass the threshold. In contrast, CLIPF calculates the word frequency by the threshold to prioritize words and maximize the input slots. For instance, using CC3M as an example, the average length of text before masking is 10.31 ± 4.7, and after SW-CLIPF masking, it is 4.26 ± 2.6 [11]. It is clear that the use of the input slots is not maximized when the input length of text tokens is set to 6 or longer. For a fair comparison with CLIPF, we pre-train SW-CLIP using the same setup: 75% image masking and 81.25% text masking. Subsequently, we fine-tuned both models with an additional epoch without any image or text masking. The results in Table 9 show that after fine-tuning CLIPF outperforms SW-CLIP.
Table 9: Comparison of the zero-shot accuracy performance of SW-CLIP and CLIPF on ImageNet-1k.
Models Datasets Image Tokens Text Tokens ViT-B/16
pre-train fine-tune
SW-CLIP CC3M 49 6 14.9 16.8
CLIPF 14.4 18.2
SW-CLIP CC12M 49 8 35.9 38.5
CLIPF 36.6 39.3
Table 10: We pre-train CLIPF on CC12M dataset across different thresholds in Equation 2. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
model Image tokens Text tokens Threshold ViT-B/16
pre-train fine-tune
CLIPF 49 8 1e-5 35.6 38.4
1e-6 36.6 39.3
1e-7 36.1 38.6
6.4 Threshold Analysis
To investigate the impact of the threshold $t$ in Equation 2 on model performance, we pre-trained models with varying thresholds. As shown in Table 10, thresholds of $51e-$ 5, $61e-$ 6, and $71e-$ 7 yield comparable results, all outperforming the other text masking strategies. This indicates that CLIPF is relatively insensitive to small variations in thresholds, which are intended to maintain the word masking probability within the range of 0 to 1. However, using too small a threshold reduces the differences in text masking probabilities; therefore, we do not recommend using an excessively low threshold.
7 Conclusion and Outlook
In this paper, we have introduced CLIPF an approach for text masking based on word frequency that improves the performance of CLIP. Unlike syntax-based masking methods, CLIPF does not require part-of-speech (POS) information, which is time-consuming to process. Unlike truncation, random, and block text masking strategies, which treat all words the same, CLIPF assigns different probabilities based on word frequency.
In the zero-shot transfer evaluation, CLIPF demonstrates a notable ability to rapidly learn image and text representations when compared to a model that pre-trains on the full text and the other text masking strategies. Finally, we analyze the word distribution before and after applying text masking to provide insight into why CLIPF is effective. We establish that CLIPF balances retention of nouns, similar to syntax masking, with retention of other parts of speech, similar to block and random masking, effectively providing the best of both worlds.
Future should investigate two directions. First, CLIPF should be tested on larger data sets, that are beyond the reach of our own computational resources. The fact that CLIPF is effective both for CC3M as well as for the larger CC12M data set suggests that CLIPF scales well. We believe that CLIP could be adapted to any larger-scale image-text dataset; however, this should be shown empirically. We see no intrinsic reason why the advantages of CLIPF should not extend to large data sets, in which the frequency of the most frequent words can be expected to be even larger than for small datasets. Second, now that we have gained insights into the distributional properties of text data that appear important for successful text masking, future work can attempt to optimize masking directly to achieve these properties. In the meantime, frequency is what you need.
Appendix A APPENDIX
Appendix B Text Masking Analysis
B.1 Text Masking Cases
Table  11 presents the potential text resulting from the text masking technique. Since truncation is fixed in each epoch, it may result in the loss of important information at the end of the text. Certain words tend to appear in specific positions within the text; for example, “the” and “a” are most likely to be the first words of a sentence. As shown in Figure 2, truncation retains more occurrences of “the” and “a” than other text masking strategies. In contrast, random and block strategies can generate different texts in each epoch for text data augmentation, but they may retain some words with a high frequency. Syntax masking retains the nouns in the text and remains the same in each epoch. However, this strategy may cause the model to overfit on the frequency of noun words. In contrast, CLIPF varies the text in each epoch, as words may be retained or removed according to their frequency. This strategy serves two primary purposes: it enhances text diversity and reduces the risk of overfitting to frequent words. Another advantage of CLIPF is that it can remove frequent prepositions that are less directly relevant to the objects in the image, such as “a,” “in,” “of,” and others. This helps the model focus on the most helpful aspects of the content. Table 12 shows the masking probabilities for certain words. High-frequency words such as “of" “the" “and" and “." are highly likely to be masked from the text.
Table 11: The example provided showcases from the CC12M dataset before and after text masking is applied, reducing the text from 32 tokens to 8 tokens. The first row displays the original text, while the subsequent rows show potential outcomes after applying text masking. The value of $t$ of formular 2 is set to $10^{-6}$.
Text Walk of the happy young couple and Siberian dog. The handsome man is hugging the smiling red head girl
Truncation walk of the happy young couple
Random the happy and the is the
Block couple and siberian dog . the
Syntax walk dog man smiling head girl
CLIPF siberian handsome man hugging smiling red
Table 12: The probability of masking words in the text is calculated using the formula provided in Equation 2. The example is selected from CC12M [15]. The value of $t$ of formular 2 is set to $10^{-6}$.
Word Probability
walk 0.926171
of 0.992838
the 0.995064
happy 0.951695
young 0.957311
couple 0.941174
and 0.991695
siberian 0.75092
dog 0.960531
. 0.993678
B.2 Word Distribution Analysis for CLIPF
In addition, we analyzed frequency text masking strategies that varied according to the number of words used, as shown in Fig. 3. As the number of words decreased from 16 to 8 and then to 6, more frequent words were masked. However, reducing the number to 4 words led to a smaller vocabulary size, resulting in the loss of some important information. Consequently, the performance of the model pre-trained with 4 words was substantially lower compared to the model pre-trained with 6 words. We recommend setting the number of text words in a frequency-based text masking strategy to strike a balance between frequent and infrequent words and to maintain a larger vocabulary. Based on our experiments, the optimal number of text masking was found to be approximately 40-60% of the average length of the original text. This configuration helps achieve a balanced word distribution, which is beneficial for pre-training VLMs.
Figure 3: Frequency-based text masking strategies vary according to the number of text words used during pre-training. The dataset used is CC12M and the value of $t$ of formular 2 is set to $10^{-6}$.
B.3 Word Distribution of CLIPF and SW-CLIP
We also compare the word distributions between CLIPF and SW-CLIP. As shown in Figure 4, SW-CLIP exhibits a flatter word distribution than CLIPF.
Figure 4: The figure illustrates the distribution of top-50 words in the text for SW-CLIP and CLIPF.
B.4 Word Category Distribution Across Different Text Masking Strategies
As shown in Table 13, we provide the word type counts of Table 8. After applying text masking strategies such as truncation, random, block, syntax, and CLIPF, the word counts remain similar. However, SW-CLIP does not maximize the use of the input slots, utilizing only 85% of the words compared to other text masking strategies. Additionally, SW-CLIP masks a high percentage of other type words, which may impact the model’s zero-shot ability.
Table 13: Distribution of syntax counts and percentages before and after applying text masking. The dataset is CC12M and we retain 6 words for each text after applying text masking.
Masking NN Count JJ Count VB Count OTHER Count NN (%) JJ (%) VB (%) OTHER (%) Total
Before masking 103,469,117 10,245,828 10,649,943 81,351,966 50.30% 4.98% 5.18% 39.55% 205,716,854
Truncation 28,628,129 2,859,462 3,272,401 20,585,364 51.72% 5.17% 5.91% 37.20% 55,345,356
Random 28,334,735 2,666,847 2,790,600 21,550,517 51.21% 4.82% 5.04% 38.93% 55,342,699
Block 27,314,723 2,624,594 2,897,434 22,510,031 49.37% 4.74% 5.24% 40.66% 55,346,782
Syntax 48,989,317 1,483,666 1,618,088 3,245,892 88.53% 2.69% 2.92% 5.87% 55,336,963
SW-CLIP 34,645,367 3,346,676 4,073,825 6,062,345 71.98% 6.95% 8.47% 12.60% 48,128,213
CLIPF 33,516,439 2,803,409 3,473,119 15,666,310 60.43% 5.06% 6.20% 28.24% 55,459,277
Figure  5 shows the word cloud before and after applying text masking. In these visuals, the font size represents the relative frequency of each word. Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency. After applying frequency word masking, the word cloud in Table 5(b) shows a more balanced distribution of word frequencies compared to before the masking. The word “the" appears smaller than in other word clouds. The reduction in frequent words allows infrequent words to become more prominent. The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full-text. As depicted in Table 5(f), syntax masking tends to retain more nouns such as “person," “stock," and “illustration" that show a different pattern from the others.
Figure 5 shows word clouds before and after applying text masking. In these visuals, the font size represents the relative frequency of each word. Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency. After applying word frequency masking, the word cloud in Figure 5(b) displays a more balanced distribution of word frequencies compared to before masking. The word “the” appears smaller than in other word clouds. The reduction in frequent words allows infrequent words to become more prominent. The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full text. As depicted in Figure 5(f), syntax masking tends to retain more nouns such as “person”, “stock” and “illustration” resulting in a different pattern from the others.
Refer to caption Refer to caption
(a) Full Text (b) Frequency
Refer to caption Refer to caption
(c) Truncation (d) Random
Refer to caption Refer to caption
(e) Block (f) Syntax
Figure 5: The word cloud on the left displays the full text, while the one on the right shows the text after applying text masking, where we retain 6 words. The font size of the word represents the relative word frequency for each figure. The dataset used is CC12M and the value of $t$ of formular 2 is set to $10^{-6}$.
B.5 Bi-gram word Distribution
As shown in Figure 6, we also calculated the bigram distribution of the text after applying different text masking strategies. Truncation and block masking better retain information about the order of the words in the text because the words remain contiguous after masking. Random and CLIPF masking also preserves the text order better than syntax masking. Syntax masking retains more continuous bigrams—such as “stock illustration,” “vector illustration,” and “stock photography”—than other text masking strategies. These bigrams might be domain specific and not contribute to the transferrability of the VLM to other data sets.
Figure 6: The figure illustrates the distribution of top-50 bi-grams in the text with different text masking strategies.
B.6 Impact of Symbolic Characters
We note the high frequency of symbolic characters appearing in the dataset, these symbolic characters are not directly related to the depicted content of the images. Since CLIPF tends to mask these characters with high probability, we removed them during the pre-training of the model to evaluate their significance. As shown in Table 14, removing symbolic characters and stop words does not affect the model’s performance.
Table 14: We remove symbolic characters and stop words during the model’s pre-training to evaluate their zero-shot classification performance on ImageNet-1K. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC3M [23] for 30 epochs with image masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
Models Masking Image Tokens Text Tokens Remove ViT-B/16
pre-train fine-tune
FLIP 49 32 14.0 14.2
FLIP 49 32 13.9 13.7
B.7 Word and Token Analysis
Since CLIP encodes text using byte-pair encoding (BPE), some words are encoded into more than one token. Therefore, in this experiment, we calculate the masking probability based on tokens rather than words. As shown in Table 15, we compare the performance of both approaches.
Table 15: Comparison of CLIPF pre-trained with token masking and word masking on ImageNet-1K for zero-shot classification.
Models Masking Image Tokens Text Tokens ViT-B/16
pre-train fine-tune
CLIPF frequency-token 49 8 36.0 38.0
frequency-word 49 8 36.6 39.3
Table 16: Zero-shot accuracy on more classification datasets. Comparison of the zero-shot accuracy performance of CLIP, FLIP, CLIPA, and CLIPF on various datasets.
Models Method
Text Tokens
Food-101
CIFAR-10
CIFAR-100
CUB200
SUN397
Cars
Aircraft
DTD
OxfordPets
Caltech-101
Kinetics700
Flowers102
MNIST
STL10
EuroSAT
Resisc45
GTSRB
KITTI
Country211
PCAM
UCF101
CLEVR
HatefulMemes
SST2
ImageNet
CLIP 32 45.46 66.42 32.45 7.90 49.27 16.17 2.28 15.48 57.51 71.91 24.84 2.12 8.77 91.29 19.08 37.38 5.09 29.81 4.45 50.05 40.52 21.84 52.66 48.76 36.61
FLIP 32 36.26 58.09 26.89 8.73 38.56 8.16 2.98 19.95 51.85 67.17 21.57 2.94 9.35 81.89 18.02 32.30 9.68 21.39 3.63 50.26 36.51 14.67 51.37 50.08 33.70
CLIPA truncation 16 35.53 61.24 26.54 9.42 39.98 7.28 2.57 15.80 46.85 68.28 20.47 1.94 8.32 81.89 18.48 33.75 2.33 17.98 3.79 58.56 30.85 13.31 53.00 50.08 32.83
random 16 38.48 53.61 29.46 7.97 41.92 8.53 2.79 21.54 56.71 69.67 22.88 2.23 9.73 84.86 17.90 36.41 7.28 32.95 4.77 55.00 35.29 16.38 49.93 50.08 34.26
block 16 38.69 57.51 29.71 8.23 44.54 9.85 2.45 21.33 54.01 70.61 23.00 2.79 10.09 85.04 31.88 29.98 9.07 25.07 4.59 60.68 38.07 18.72 51.24 49.64 34.78
syntax 16 35.94 58.78 31.86 8.91 41.92 8.39 1.53 17.07 47.05 67.59 21.34 3.10 10.45 76.92 26.68 32.65 8.73 39.37 4.23 49.96 32.91 15.55 53.30 49.92 34.41
CLIPF frequency 16 39.47 64.43 32.55 10.15 45.51 9.22 2.43 21.70 55.62 70.69 23.08 2.77 9.56 83.00 21.02 33.30 7.20 27.47 4.12 50.24 37.72 13.70 52.70 49.92 35.96
CLIPA truncation 8 36.28 61.99 32.24 8.65 44.40 5.56 2.04 18.56 45.67 65.93 22.68 2.52 10.51 87.96 34.80 29.37 3.35 46.06 4.59 51.55 32.38 12.41 53.46 50.08 30.30
random 8 45.10 68.54 34.04 9.03 51.54 8.64 3.23 30.53 59.29 70.06 27.29 4.21 9.82 90.46 28.72 32.54 5.99 30.61 7.95 58.00 42.90 12.87 53.40 50.08 37.25
block 8 43.70 66.66 32.62 8.08 51.07 9.27 2.54 28.99 57.88 70.89 26.72 2.65 18.10 90.84 31.88 36.94 9.02 33.36 7.52 61.91 40.10 14.66 51.80 53.27 37.88
syntax 8 37.92 65.25 31.72 8.28 45.74 8.73 2.96 17.07 46.54 68.73 23.52 2.59 11.30 88.30 29.90 30.81 6.47 30.82 5.02 54.62 36.58 12.51 52.74 50.08 34.97
CLIPF frequency 8 45.75 59.00 30.76 10.60 50.06 10.32 3.20 29.89 67.54 73.35 26.96 4.09 10.10 90.54 20.80 36.30 7.16 40.31 7.10 57.23 43.09 12.27 54.54 50.03 39.30
CLIPA truncation 6 23.54 60.27 23.52 5.70 36.54 3.82 2.03 10.32 31.12 61.15 17.77 1.86 9.94 85.25 20.62 23.68 7.13 16.71 3.45 50.23 24.53 12.61 53.62 50.08 23.28
random 6 38.45 71.57 33.89 7.73 51.81 6.67 2.72 29.47 59.92 68.65 25.89 2.42 12.28 89.73 26.90 33.17 8.89 29.75 7.23 56.87 38.62 12.63 51.96 50.08 34.62
block 6 39.87 66.60 35.31 7.58 53.23 7.49 2.89 27.87 56.74 69.40 25.97 2.95 10.06 89.30 27.90 32.40 7.31 33.76 6.73 63.10 39.20 12.65 55.41 52.83 35.89
syntax 6 34.18 63.46 29.83 7.54 43.54 7.92 1.85 17.23 44.97 68.37 22.75 1.99 11.37 89.19 17.56 26.13 3.84 33.16 4.98 55.25 35.40 12.30 53.44 50.80 32.57
CLIPF frequency 6 45.33 67.64 32.83 9.98 50.86 9.61 2.19 31.33 66.03 69.24 26.60 2.94 10.71 90.98 33.64 33.22 8.93 36.70 7.32 59.96 40.60 12.45 54.15 50.08 37.81
CLIPA truncation 4 20.64 51.67 22.10 4.19 37.16 3.11 1.18 15.16 29.36 53.99 16.17 1.04 10.10 86.34 17.16 15.81 7.48 30.15 3.49 61.43 24.95 12.63 50.76 50.08 19.80
random 4 28.82 61.90 28.98 5.35 46.39 4.03 1.42 25.16 43.26 62.87 22.03 2.10 12.24 88.54 23.82 24.46 8.07 39.30 5.59 50.62 33.52 12.65 51.86 50.08 27.13
block 4 28.08 60.63 27.10 5.09 45.61 3.82 2.43 23.24 36.63 64.15 21.80 2.29 8.41 87.04 24.44 23.95 7.55 43.05 5.51 55.80 32.62 9.49 53.55 50.08 26.66
syntax 4 25.59 54.13 27.46 4.73 42.26 6.26 2.66 15.27 32.52 58.93 19.58 2.20 10.42 90.11 26.02 20.62 5.19 37.50 4.32 51.12 26.82 12.93 52.77 50.08 24.61
CLIPF frequency 4 36.09 69.76 31.14 8.41 47.72 7.01 2.24 25.64 54.23 64.55 23.55 2.95 10.18 90.25 20.78 26.52 6.90 36.63 6.40 55.71 35.71 12.65 49.63 50.08 30.93
Appendix C More results
More detailed results are presented in this section.
C.1 Zero-shot Classification on More Datasets
In Table 16, we present the evaluation results of CLIPF following the approach of CLIP [3] on additional datasets. In almost all of the downstream datasets, when using six text tokens, CLIPF surpasses not only the baseline CLIP without image and text masking and FLIP with image masking and full-text tokens but also outperforms other text masking strategies. Similar to observations from the ImageNet datasets, reducing the number of text tokens from 16 to 8 substantially enhances CLIPF’s performance across all datasets. Specifically, performance improvements are noted: SUN397 [38] by 6.39%, DTD [39] by 9.31%, OxfordPets [40] by 4.64%, STL10 [41] by 9.31, EuroSAT [42] by 6.60%, UFC101 [43] by 6.68% and ImageNet-1K [44] by 3.13%. The performance of other text masking strategies, including truncation, random, block, and syntax, also improves with this reduction. However, the performance gains are not as large as those observed with CLIPF. This improvement suggests that employing text masking strategies as a data augmentation technique can enhance VLMs’ performance in zero-shot classification tasks. CLIPF, utilizing a word frequency masking method, demonstrates robust zero-shot classification performance and establishes a benchmark for state-of-the-art results.
C.2 Zero-shot Image-text Retrieval
We report the details of zero-shot image-text retrieval for Flickr30k and COCO datasets, the details are shown in Table 17.
Table 17: Zero-shot Image-Text Retrieval, we evaluated CLIP, FLIP, and CLIPF image-text retrieval performance on COCO and Flickr30k datasets. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
Models Masking Text Tokens Text Retrieval Image Retrieval
Flickr30k COCO Flickr30k COCO
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CLIP 32 62.62 86.00 91.81 35.54 62.38 74.08 45.42 72.56 81.50 24.22 48.42 60.42
FLIP 32 54.73 80.37 87.97 29.34 56.08 67.00 38.62 66.09 75.40 20.88 43.22 54.78
CLIPA truncation 16 44.67 73.08 81.85 25.54 51.90 64.64 34.99 61.05 70.85 19.64 41.91 53.51
random 16 58.48 83.93 90.53 32.36 58.76 69.98 43.61 70.67 80.04 22.81 46.00 57.71
block 16 56.51 81.26 89.05 30.82 58.32 70.38 44.06 71.20 80.10 22.66 46.24 58.06
syntax 16 54.54 81.07 88.36 29.60 56.52 68.54 41.16 68.32 77.51 21.40 44.82 56.56
CLIPF frequency 16 57.89 84.62 90.04 31.52 58.38 70.30 42.72 69.61 78.60 22.57 46.15 57.95
CLIPA truncation 8 30.47 59.47 70.51 16.92 38.62 51.34 23.96 47.29 57.85 12.78 31.66 42.92
random 8 58.58 84.52 91.81 36.24 62.16 72.90 43.79 70.89 80.14 23.16 46.74 58.86
block 8 60.06 85.01 92.11 35.88 62.58 73.74 45.98 73.23 82.49 24.65 48.81 60.70
syntax 8 50.30 78.30 86.98 29.64 55.42 67.18 38.46 66.77 77.36 20.28 43.78 55.40
CLIPF frequency 8 60.36 85.80 90.63 35.28 62.70 73.70 44.81 72.07 81.34 23.66 48.05 59.89
CLIPA truncation 6 32.25 61.05 72.98 16.18 39.52 52.48 22.39 47.46 59.39 11.93 30.03 40.86
random 6 55.23 82.54 89.25 32.58 57.54 68.68 42.47 70.12 79.43 21.34 44.79 56.26
block 6 54.44 79.68 88.66 33.12 58.96 70.64 41.68 69.13 79.33 21.73 45.23 57.29
syntax 6 46.15 76.04 84.81 26.78 52.56 64.62 34.08 61.76 73.25 17.99 40.22 52.05
CLIPF frequency 6 56.02 82.05 88.56 32.32 58.58 70.00 41.05 69.17 78.88 21.28 44.55 56.05
CLIPA truncation 4 30.97 56.51 67.06 17.44 38.48 49.76 21.09 43.77 55.70 11.07 27.98 38.54
random 4 41.62 69.53 80.18 24.14 48.32 60.14 30.26 56.92 68.42 15.63 35.72 47.30
block 4 40.83 69.03 79.09 23.56 47.94 59.20 30.77 57.85 69.29 14.98 35.64 47.35
syntax 4 38.66 65.88 75.74 21.68 44.38 56.60 26.08 52.80 64.32 14.14 33.71 44.98
CLIPF frequency 4 45.07 72.49 81.95 25.60 49.98 61.96 31.72 59.35 71.66 15.96 36.58 48.32
C.3 Learn Curved before fine-tuning
In Figure 1, we show the learning curve after fine-tuning. Moreover, we show the learning curve before fine-tuning, they have a similar performance for every text masking strategy.
Figure 7: Zero-shot classification accuracy on ImageNet-1K over the epochs for the different text masking strategies before fine-tuning.

### v2

@@
 Recent work has shown the benefits of masking text during VLM training using a variety of approaches: truncation, random masking, block masking and syntax masking.
-In this paper, we show that the best masking strategy changes over training epochs and that, given sufficient training epochs, word frequency information is what you need to achieve the best performance.
-Experiments on a large range of data sets demonstrate the advantages of our approach, called Contrastive Language-Image Pre-training with word Frequency Masking (CLIPF).
+In this paper, we show that the best masking strategy changes over training epochs and that, given sufficient training epochs.
+We analyze existing text masking approaches including syntax masking, which is currently the state of the art, and identify the word frequency distribution as important in determining their success.
+Experiments on a large range of data sets demonstrate that syntax masking is outperformed by other approaches, given sufficient epochs, and that our proposed frequency-based approach, called Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF) has numerous advantages.
 The benefits are particularly evident as the number of input tokens decreases.
-We analyze the impact of CLIPF vs.
-other masking approaches on word frequency balance and discuss the apparently critical contribution of CLIPF in maintaining word frequency balance across POS categories.
 Keywords Vision-Language Model  $\cdot$ Multimodal Data  $\cdot$ Word Frequency Masking.
@@
 CLIP (Contrastive Language-Image Pre-training) [3, 9, 10] is a VLM that is trained on a very large number of image-text pairs collected online, where the text is considered a “natural language” annotation for the image.
-This paper contributes to the recent trend of text-masking approaches that improve VLMs by removing words from the text data during training, speeding up training in the process [11, 12].
-We introduce CLIPF (Contrastive Language-Image Pre-training with word Frequency Masking), which selects words to be masked on the basis of word frequency.
-We carry out a comparison with CLIPA masking [12], which selects words on the basis of four strategies that do not use frequency information: truncation, random word selection, random block selection, and part-of-speech (POS) information and is currently considered the state-of-the-art text-masking.
-The comparison of the learning curves of our CLIPF with CLIPA masking in Fig.
-1 illustrates two key contributions of this paper.
-First, we demonstrate that the number of training epochs is crucial to which masking strategy yields the highest performance.
-The finding is an important improvement on the state of the art of text masking for CLIP training, represented by [12].
-In [12], CLIP is trained on a limited number of epochs and syntax masking is found to outperform truncation, random, and block masking.
-In contrast, we show that random and block masking outperform syntax masking after a sufficient number of epochs.
-Second, we show that when masking text for CLIP training, frequency is what you need.
-As evident in Fig.
-1, our frequency-based CLIPF approach outperforms all CLIPA masking approaches with sufficient epochs of training, leading it to be the best performing text masking approach.
-Table 1: Illustration of different text masking methods and words masking probabilities.
-Bold percentages or check marks provide an example of the words that would be retained by each text masking strategy.
-For the Truncation and Syntax methods, four words per text are retained as the input.
-For the Random, Block, and CLIPF methods, we present the word masking probabilities.
-The second row shows the part-of-speech (POS) category of each word.
-Models Methods Words
-a small dog sleeping on a couch next to a remote .
-POS Other Adjective Noun Verb Other Other Noun Other Other Other Noun Other
-CLIPA Truncation
-Random 50% 50% 50% 50% 50% 50% 50% 50% 50% 50% 50% 50%
-Block 50% 50% 50% 50% 50% 50% 50% 50% 50% 50% 50% 50%
-Syntax
-CLIPF Frequency 99.37% 96.28% 96.13% 88.35% 98.96% 99.37% 87.03% 93.69% 98.89% 99.37% 81.11% 99.37%
+However, pre-training CLIP models typically requires thousands of GPU-days [3, 9].
+To accelerate the pre-training process, FLIP, A-CLIP, and CLIPA use masking strategies for image and text tokens when pre-training CLIP models [8, 11, 12].
+In addition to the speed-up, masking makes it possible use large batch sizes, which enhances the benefits of contrastive learning and has been considered as a form of regularization, reducing the risk of overfitting [8].
+For these reasons, despite the loss of some information from the original image and text, masking strategies have the potential to improve model performance.
 Figure 1: Zero-shot classification accuracy on ImageNet-1K over training epochs for CLIPF and CLIPA strategies.
-The text masking ratio is 75%.
 The backbone of the image encoder is ViT-B/16.
-The model is pre-trained on CC12M for 30 epochs.
-We use image masking (75%) to speed up training.
+The models are pre-trained on CC12M for 45 epochs with 75% text masking ratio.
+We also use image masking (75%) to speed up pre-training.
 All approaches are subsequently fine-tuned with one epoch on the full, unmasked data.
+Currently, the state of the art in text masking is represented by CLIPA [12], which compares four masking strategies: truncation, random word selection, random block selection, and syntax-based selection, i.e., masking on the basis of part-of-speech (POS) information.
+The conclusion is that syntax masking is the strategy that, overall, provides the best quality.
+In this paper, we carry out a set of experiments that demonstrate that this conclusion must be revisited.
+Specifically, as shown in Fig.
+1, if enough epochs are used, other forms of masking substantially outperform syntax masking.
+The conclusion of CLIPA [12] misses the weakness of syntax masking because it trains with a limited number of epochs (6.4 epochs).
+The special focus of our paper is a frequency masking approach that we introduce: CLIPF (CLIP Frequency masking).
+In Fig.
+1, we see that with enough epochs CLIPF and CLIPA block masking perform similarly.
+We provide an analysis that demonstrates why CLIPF, block masking, and random masking emerge as clearly superior to syntax masking and truncation.
+The analysis studies the distribution of words by frequency of occurrence in the training data after masking.
+It shows that good masking strategies maintain helpful properties of this distribution and that syntax masking distorts the distribution of words over POS categories.
+These observation leads to the main insight of our paper, namely, when masking text for VLM pre-training, what you need for best results is to pay attention to the impact of masking on word frequency.
+Our CLIPF approach is useful in practice because it performs relatively well in the range of 16-40 epochs.
+This range is a typical number of epochs to train a system in practice.
+For example, Open_CLIP [9] trains on 30 epochs.
+In Fig.
+1, we see that in this range CLIPF outperforms other masking strategies.
+We carry out a set of experiments with 30 epochs to shed further light on the role of frequency in masking by comparing CLIPF vs.
+the different variants of CLIPA masking on zero-shot tasks and on image retrieval.
+We also investigate training on a larger data set.
 The history of text masking in language model training dates to [13], where subsampling of frequent words in the training text was used to improve the performance of skip-gram language models.
 The study of text masking for VLMs got off to a slow start, since initially [8] reported that text masking, although it promotes efficiency, deteriorated performance.
-Subsequently, [12] expanded the types of masking studied and demonstrated that masking can indeed improve performance.
-Simultaneously, frequency-masking was proposed at a workshop [11], but fell short of the potential of frequency realized by CLIPF due to the use of a cut-off threshold.
-To our knowledge, our work reported in this paper is the first to attempt to discover the reasons that certain forms of text masking outperform others.
-An initial illustration of the differences between task-masking strategies is provided in Table 1.
-Truncation consistently retains the beginning of the text.
-With the random (random word selection) and block (random block selection) methods, each word has an equal probability of being masked, with block masking maintaining word neighborhoods.
-Syntax masking prioritizes the retention of nouns, followed by other part-of-speech (POS) categories.
-Masking based on frequency (CLIPF) has the potential to address the shortcomings of the other approaches.
-Unlike truncation, it does not prioritize words that tend to occur at the beginning of the sentence.
-In contrast to random and block masking, it represses words with the POS category “Other”, which are often stop words and for this reason high frequency.
-However, it does not give an overly high priority to nouns and adjectives, as with syntax masking.
-In sum, the contributions of this paper are as follows:
-  • We point out, to our knowledge for the first time, the importance of the number of training epochs for comparing text masking approaches.
-  • We introduce CLIPF, a frequency-based masking approach that outperforms previously proposed approaches.
-  • We carry out an analysis that provides insight into why CLIPF outperforms other approaches.
-Our work is reproducible because it uses training data sets that are openly available and small enough to allow researchers with limited computational resources to repeat our experiments and because we release our code [^1] https://github.com/MingliangLiang3/CLIPF
+The CLIPA [12] work is important because it expanded the types of masking studied and demonstrated that masking can indeed improve VLM performance.
+To our knowledge, no previous work has attempted to discover the reason that certain forms of text masking outperform others or identified the importance of the way in which text masking approaches impact word frequency.
+In sum, our paper makes the following contributions to the study of text masking in pre-training VLMs:
+  • We point out the importance of the number of training epochs when comparing text masking strategies.
+  • We show that word frequency distribution after masking differentiates between successful and less successful text masking strategies.
+  • We carry out an in-depth comparison between existing masking strategies (CLIPA variants) and an approach that carries out masking by using frequency information directly, CLIPF.
+Our experiments show that CLIPF consistently outperforms syntax masking, previously considered the best choice for masking and provide a systematic view of how text masking can accelerate pre-training of VLMs while at the same time maintaining or improving VLM performance.
+Our work is reproducible because it uses training data sets that are openly available and small enough to allow researchers with limited computational resources to repeat our experiments, and because we release our code [^1] https://github.com/MingliangLiang3/CLIPF.git.
 2 Related work
 2.1 Vision-Language Models
-Vision-language models (VLMs), such as CLIP, ALIGN, LiT, and FLIP [3, 4, 10, 8], learn image embeddings from large-scale natural language supervision using contrastive learning [14] to grants the models remarkable transferability.
+Vision-language models (VLMs), such as CLIP, ALIGN, LiT, and FLIP [3, 4, 10, 8], learn image embeddings from large-scale natural language supervision using contrastive learning [14], which is important for making models transferable.
 Because CLIP does not release its pre-training dataset, OpenCLIP pre-trained models on several large-scale public datasets [15, 16, 17, 18] to reproduce CLIP.
@@
 While random text masking involves masking text arbitrarily, truncation masking in FLIP strategically eliminates padding tokens and retrains the words that preceded them.
-Subsequently, CLIPA [12] introduces block masking and syntax masking.
-As already mentioned in Section 1, CLIPA [12] shows text masking during training can improve the performance of a VLM.
+Subsequently, CLIPA [12] introduced block masking and syntax masking.
+As already mentioned in Section 1, CLIPA [12] shows that text masking during training can improve the performance of a VLM.
 CLIPA syntax masking prioritizes the retention of nouns, followed by adjectives and then other words.
@@
 An initial effort to exploit frequency information for text masking during VLM training was made by SW-CLIP [22].
-Unfortunately, SW-CLIPF introduces a frequency cutoff threshold, which reduces performance, as we demonstrate later in this paper.
+Unfortunately, SW-CLIPF introduces a frequency cutoff threshold, which reduces performance.
 In order to exploit frequency information effectively each training text needs to fill as many available input tokens as possible, as is done in CLIPF.
-3 Method
-In this section, we describe various text masking strategies, including our CLIPF method. All these strategies accelerate pre-training by masking specific words in the text.
-3.1 Text Masking
+3 Text Masking
+In this section, we describe the text masking strategies that we study in this paper: the current state of the art, namely, the text masking strategies of CLIPA [12] (Section 3.1) and our frequency masking strategy CLIPF (Section 3.2).
+All these strategies accelerate pre-training by masking specific words in the text.
 Consider a text $T_{i}$ consisting of words $\{w_{1},w_{2},w_{3},\ldots,w_{n}\}$, where $n$ represents the total number of words in $T_{i}$.
@@
 And, we define the number of input tokens as $k$.
-We will sequentially introduce our masking strategies: random, truncation, block, syntax, and frequency.
 It is important to note that the number of padding tokens remains consistent across all these strategies.
+3.1 Existing Text Masking Strategies (CLIPA)
+The four text masking strategies from CLIPA [12], which we also investigate here, are:
 Random Masking: In the random text masking strategy, the function $f$ operates to randomly select words from the text $T_{i}$.
@@
 This method ensures that the core semantic information of the text is less likely to be masked, especially when learning the relationship between the image and the text, as nouns are more relevant to the object of the image.
-3.2 Word Frequency Masking
-As in previous work, subsampling of frequent words was introduced by [13] in the natural language processing (NLP) domain to accelerate the learning process for word representations.
-The goals are to accelerate training and to address the imbalance between frequent and infrequent words in the training corpus.
-In our work, we use the formula (Formula 2) introduced in [13] to determine the probability of masking a word during training.
-Thus, the complete formula to calculate the probability $P(w_{i})$ that accounts for all scenarios is:
-$P(w_{i})=\begin{cases}1&\text{if}N<5\\ 0&\text{if}f(w_{i})<t\\ 1-\sqrt{\frac{t}{f(w_{i})}}&\text{otherwise}\end{cases}$ (2)
+3.2 Word Frequency Masking (CLIPF)
+As mentioned in the introduction, subsampling of frequent words was introduced by [13] in the natural language processing (NLP) domain to accelerate the learning process for word representations.
+The goals are to speed up training and to address the imbalance between frequent and infrequent words in the training corpus.
+Our CLIPF method, uses Formula 2, which was introduced in [13], to determine the probability of masking a word during training.
+Thus, the complete formula to calculate the probability $P(w_{i})$ is:
+$P(w_{i})=1-\sqrt{\frac{t}{f(w_{i})}}$ (2)
 Here, the masking probability of the word $w_{i}$ is $P(w_{i})$, $t$ is the threshold, and $f(w_{i})$ is the frequency of the word.
-Additionally, if the total number of words ($N$) is fewer than 5, deemed too infrequent for effective learning, the probability that the word will be masked is 1.
-Following FLIP [8], we fine-tune the model with an additional epoch to eliminate the distribution gap between the masking and unmasking.
+This formula scales the word frequency $f(w_{i})$, resulting in a much higher masking probability for frequent words than infrequent ones.
+It also preserves the overall ranking of word frequencies while aggressively masking high frequency words, controlled by the threshold $t$ [13].
 4 Experimental Setup
-Our implementation follows CLIP [3], OpenCLIP [9], FLIP [8] and CLIPA [12]. In this section, we present the details of the training and testing.
-4.1 Implementation
-Dataset We pre-train the models on Conceptual Captions 3M (CC3M) [23] and Conceptual 12M (CC12M) [15].
+Our implementation follows CLIP [3], OpenCLIP [9], FLIP [8] and CLIPA [12]. In this section, we present the details of our experimental setup.
+Dataset We pre-train the models on Conceptual Captions 3M (CC3M) [24], Conceptual 12M (CC12M) [15].
 These datasets were selected due to their diverse range of real-world concepts and scenes.
-Popular VLMs such as OpenCLIP, DeCLIP, and SLIP [9, 20, 5] also employ these datasets for training, leveraging their extensive and diverse content.
+Popular VLMs such as OpenCLIP, FLIP, DeCLIP, SLIP, A-CLIP and CLIPA [9, 8, 20, 5, 11, 12] also employ these datasets for training, leveraging their extensive and diverse content.
 Due to the expiration and inaccessibility of some URLs in the training set, we have downloaded 2.7 million image-text pairs for CC3M and 9.3 million image-text pairs from CC12M.
@@
 Since the number of input tokens is fixed, the impact of different text masking ratios varies across datasets due to differences related to caption length.
+In an additional set of experiments, we provide results training on LAION-400M [17], a larger data set with about 300M image-text pairs.
 Table 1: Dataset Statistics for Pre-training datasets. Caption length refers to the number of words in the text.
@@
 Architecture For the image encoder, we use ViT-B/16 [25] architectures with global average pooling and learnable positional embeddings.
-The input image size is 224.
+The input image size is 224 for CC3M and CC12M.
 For the text encoder, we use a Transformer-based model [26] and byte-pair encoding with a 49K token vocabulary.
@@
 We trained the models using 8 RTX A5000 GPUs with the same settings to ensure consistent conditions across all models.
-This allows us to effectively measure and compare the performance of each model in processing text tokens during training.
 We experiment with different text token input sizes, namely, 32, 16, 8, 6, and 4 text tokens.
 As the input size gets smaller, we can increase the batch size, maximizing computational memory usage.
-We used batch sizes of 664, 832, 896, 928, and 944, respectively.
+For CC3M and CC12M, the batch sizes corresponding to the input sizes are: 664, 832, 896, 928, and 944.
 Note that CLIP (and therefore CLIPA and CLIPF) is trained using contrastive learning, which benefits from larger batch sizes.
@@
 numerical precision amp amp
-augmentation RandomResizedCrop RandomResizedCrop
-Table 4: During pre-training, each model processes a certain number of image and text tokens.
-The first and second columns of the table display the image and text masking ratios, respectively.
-The third and fourth columns show the number of image and text tokens processed after masking.
-The fifth column presents the total number of tokens (image + text) after masking.
-The last column provides the percentage of total tokens relative to the unmasked case.
+RandomResizedCrop (50, 100) (50, 100)
+Table 3: The number of image and text tokens that are processed during pre-training (CC3M and CC12M).
 Image masking Text masking Image tokens Text tokens Total Percentage
@@
 87.50% 49 4 53 23.25%
-Training We follow the setup of OpenCLIP [9], FLIP [8] and CLIPA to pre-train our model.
-Following OpenCLIP [9], we pre-train our model for 30 epochs on the CC3M and CC12M datasets.
-Details of the pre-training configuration are given in Table  3.
+Training Following CLIP, OpenCLIP [3, 9], we pre-train our model for 30 epochs on the CC3M and CC12M datasets. Details of the pre-training configuration are given in Table 2.
 Similar to FLIP, to speed up training, we apply 75% image masking to the image encoder as the baseline model.
@@
 During text masking, we reduce the number of tokens from 32 to 16, 8, 6, and 4.
-To measure the training speed of CLIPF, CLIP, and CLIPA, we compare the number of text tokens processed by each model.
+To measure the training speed of CLIP, CLIPA, and CLIPF, we compare the number of text tokens processed by each model.
 As shown in Table 3, the number of text tokens processed by each model during pre-training is calculated based on different image and text masking ratios.
@@
 Moreover, when we apply 87.5% text masking, the total number of tokens is 53, resulting in a speed increase of approximately 34% compared to training without text masking.
-Fine-tuning Following FLIP, and CLIPA, in order to bridge the gap between pre-training and evaluation, we fine-tuned the model without images and text masking.
+Fine-tuning Following FLIP, and CLIPA, in order to bridge the distribution gap between pre-training and evaluation, we fine-tuned the model without images and text masking.
+Note that, fine-tuning the models without masking in masking work is focused on reducing the distribution gap and applied sparingly to avoid undoing the tradeoff.
 The details of the fine-tuning configuration are provided in Table 2.
@@
 Reported as top-1 accuracy.
-All models use a ViT-B/16 image encoder backbone pre-trained on CC3M and CC12M for 30 epochs with 75% image masking for speedup (pre-train) and fine tuned for an additional epoch on all data (fine-tune).
+All models use a ViT-B/16 image encoder backbone pre-trained on CC3M and CC12M for 30 epochs with 75% image masking for speedup (pre-train) and fine-tuned for an additional epoch on all data (fine-tune).
 “Image tokens” and “text tokens” reflect the amount of image and text processing during pre-training.
@@
 CLIPF frequency 10.9 16.0 17.0 30.9
-4.2 Evaluation
-Evaluation setup Following the CLIP [3], FLIP [8], and CLIPA [12] methods, several classification benchmarks were used to evaluate the zero-shot capability of CLIPF.
+CLIPF frequency 98 8 19.6 18.7 39.8 41.0
+Evaluation setup Following CLIP [3], FLIP [8], and CLIPA [12], several classification benchmarks were used.
 Among these benchmarks is ImageNet-1K, a widely recognized dataset in computer vision.
-It is frequently used for image classification and VLMs tasks and comprises 50K validation samples across 1K different classes.
+It is frequently used for image classification and VLM tasks and comprises 50K validation samples across 1K different classes.
 We filled each class into the templates provided by CLIP [3] to calculate the average of the text embeddings.
 We use the same evaluation settings as CLIP [3] to evaluate the other downstream datasets.
+5 The Importance of Frequency Distributions
+Our work is motivated by the patterns that we observe when VLMs are trained for additional epochs, shown in Fig.
+1.
+This experiment motivates revisiting CLIPA [12], which was trained only on 6.4 epochs, in two respects.
+First, as already mentioned in the introduction, the conclusion that syntax masking is the best masking approach does not hold.
+Second, truncation masking, which looks relatively promising in low numbers of epochs, does not perform well given more epochs.
+In this section, we analyze the impact of masking on the training data in order to discover the reason behind this behavior.
+Figure 2: Comparison of word masking probabilities for various methods.
+In this example, we keep four words for truncation and syntax.
+Numbers indicate the probability of a word being masked.
+Truncation keeps the first 4 words; Random and Block mask each word with a 50% probability; Syntax prioritizes retaining nouns, followed by adjectives, then others; Frequency masks words based on the frequency of the words.
+We start with a basic illustration of the differences between task-masking strategies is provided in Figure 2.
+Truncation consistently retains the beginning of the text.
+With the random (random word selection) and block (random block selection) methods, each word has an equal probability of being masked, with block masking maintaining word neighborhoods.
+Syntax masking prioritizes the retention of nouns, followed by other part-of-speech (POS) categories.
+Figure 3: The figure illustrates the distribution of top-50 words in the text before and after applying various text masking strategies.
+We set the text length after text masking to 6.
+The x-axis represents the word index, which is sorted by counts of the original data, and the y-axis shows the word frequency.
+The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$.
+We remove special characters from the vocabulary.
+Next, we look at how the different masking strategies impact the frequency distribution of words in the training data.
+Fig.
+3 shows the distribution of words in the text before and after applying all of the text masking strategies.
+We see that in general, there is quite a large difference between the frequency of the highest-frequency words and the frequency of words with lower frequency.
+Such a long-tail distribution is characteristic of words occurring in text, which is known to be Zipf distributed.
+Most masking approaches preserve the general shape of this distribution: The frequency of a word in the masked training data tracks its frequency in the original data.
+The exception is syntax masking, which shows peaks and pits for specific words.
+Specifically, because syntax masking retains more nouns, the articles “the” and “a” are less frequent, but the nouns create spikes.
+The rationale for prioritizing nouns is solid.
+Nouns are closely related to the depicted content of images and, for this reason, are considered to be more helpful.
+Indeed, the helpfulness of nouns for the VLM is reflected in the fact that syntax masking learns faster than other text masking strategies.
+In Fig.
+1, we saw that syntax masking outperforms the other masking approaches during the early epochs.
+However, retaining too many nouns can have two adverse effects.
+First, the model can overfit to certain high frequency nouns.
+In Fig.
+3, we see that CLIPA syntax masking retains more occurrences of “person,” “illustration,” and “stock” than the other strategies.
+Second, retaining nouns means necessarily dropping words from other parts-of-speech (POS) categories, which may be important, as discussed in more detail below.
+Considering the CLIPA truncation masking strategy, we see in Fig.
+1 that the word frequency distribution of the truncated training data shows a difference with the other masking strategies.
+Specifically, truncation is the top line on the left of the graph, which means truncation is making the most highly frequent words in the training data relatively even more frequent.
+The fact that truncation impacts word frequency might seem surprising at first, but it can be explained if we consider that the probability of a word occurring changes with sentence position.
+With truncation, typical text beginnings, e.g., “A person is…”.
+remain over-represented in the data.
+To dive deeper into the importance of the distribution of words over POS categories, we provide statistics for each of the different text masking strategies in Table 5.
+Consulting the column “Other”, we see an interesting pattern.
+Syntax masking changes the proportion of words in the “Other” category radically with respect to the data before masking, but truncation, random and block do not.
+CLIPF appears to provide a happy medium.
+CLIPF retains more nouns than CLIPA truncation, random, and block, but not so many that the “Other” category becomes poorly represented.
+Table 5: To examine the distribution of words belonging to different POS categories, we calculate the percentage of words in the training data belonging to each category in before and after applying text masking.
+The dataset is CC12M, and we retain 6 words for each text after applying text masking.
+Masking NN (%) JJ (%) VB (%) OTHER (%)
+Before masking 50.30% 4.98% 5.18% 39.55%
+Truncation 51.72% 5.17% 5.91% 37.20%
+Random 51.21% 4.82% 5.04% 38.93%
+Block 49.37% 4.74% 5.24% 40.66%
+Syntax 88.53% 2.69% 2.92% 5.87%
+CLIPF 60.43% 5.06% 6.20% 28.24%
+The importance of “Other” words is surprising because the category includes many words, which like articles, lack semantic content and are sometimes considered to have no direct connection to images.
+In fact, in many NLP applications, articles and other “stop words” are simply removed.
+However, since the CLIP text encoder is a transformer, technically it should be able to exploit the context of content words, particularly words in “Other” category.
+Note that with syntax masking a phrase like “smoking in the room” in the training data would always be reduced to “smoking room”, which has a different meaning.
+However, with the other masking approaches, “in the” would not be consistently deleted.
+Note specifically that block masking retains entire phrases, without always using the beginning of the sentence as truncation masking does.
+Also, frequency masking (CLIPF) retains entire phrases in some cases, more than would be expected with random masking.
+In sum, when trained with enough epochs, overfitting effects with syntax masking occur.
+Our analysis has related this problem to peaks in the word frequency distribution caused by over-retention of nouns.
+Also, with enough epochs, truncation masking overfitting effects occur, which our analysis has related to distortion of the word frequency distribution caused by oversampling of the beginning of sentences.
+In sum, our analysis of word frequency distributions in the training data provides insight into why the performance of syntax and block masking falls off in Fig.
+1.
+Further, if word frequency is indeed important, then masking that controls word-frequency directly, such as our CLIPF, can be considered to be a sensible and transparent approach.
 6 Experimental Results
-5.1 Zero-short performance on ImageNet
+In this section, we carry out a systematic investigation of the existing text masking approaches from CLIPA [12] and our CLIPF approach.
+Our results support our observation that syntax and truncation masking are not productive approaches to masking and highlight the benefits of masking that directly leverages frequency information, i.e., our CLIPF.
+Our experiments use 30 epochs, following Open_CLIP [9].
+We note that VLM training work has shown performance improvements up to 100 epochs [5, 11] and we would expect that with more epochs the performance would also improve.
+However, the improvement beyond 30 is expected to be small, predictable and not necessarily important in an operational setting where conserving compute is of concern.
+6.1 Zero-shot Performance on ImageNet-1k
 Table 4 presents results for zero-shot classification on ImageNet-1K for models trained on both the CC3M and CC12M data.
 The experiments compare different masking approaches and different masking ratios, indicated in the table with the number of input text tokens (Text Tokens column).
-Models are pre-trained for 30 epochs (pre-train column) and then fine tuned for one epoch (fine-tune column) with the full, unmasked data.
+Models are pre-trained for 30 epochs (pre-train column) and then fine-tuned for one epoch (fine-tune column) with the full, unmasked data.
 Note that since the average text length in CC3M is less than 16 tokens, the benefits of fine-tuning a model pre-trained with 16 text tokens are limited.
@@
 Contrary to the conclusion of the original CLIPA work [12], syntax masking does not yield the best performance.
-As mentioned in Section 1, we observe that the relative performance of different text masking approaches depends on the number of training epochs.
-In [12], CLIPA only pre-trained the models for 6.4 epochs.
-We discuss the learning curve in Fig.1 further in Section 6.2.
 It is particularly interesting to observe from the results in Table 4 the fact that models trained with text masking (CLIPA and CLIPF) are competitive with, or superior to, models trained on the original, unmasked text data (CLIP, FLIP).
@@
 Fine-tuned CLIPF with a 75% masking ratio achieves an accuracy of 39.3% compared to 36.6% for the original CLIP trained on the full data.
-Next we consider the comparison with FLIP, which is the same as CLIP except that it uses a 75% image masking ratio (49 image tokens).
+Next, we consider the comparison with FLIP, which is the same as CLIP except that it uses a 75% image masking ratio (49 image tokens).
 On CC3M, fine-tuned CLIPF is competitive with or outperforms fine-tuned FLIP at all masking ratios.
 On CC12M, fine-tuned CLIPF outperforms fine-tuned FLIP, except for the highest masking ratio (87.5% corresponding to 4 input tokens).
+Additionally, as shown in the last row of Table 4, if we double the number of image tokens, then CLIPF substantially beats the original CLIP (41.0% vs 36.6%).
+However, it also notes that more resources are required (Tokens processed from 25% to 46%).
 Finally, we consider the pattern of the text masking approaches as the masking ratio increases, i.e., as we decrease the input size from 16 to 4 tokens.
@@
 Moving from 16 to 8 input tokens benefits most approaches.
-Recall that the reduction of the number of input tokens allows for an increase of the batch size.
-Moving from 8 to 6 to 4 tokens, we see that CLIPF maintains its performance surprisingly well.
-As the masking ratio becomes higher, the margin by which CLIPF outperforms other masking strategies increases.
-In short, CLIPF is better positioned to deliver benefit as the number of input tokens decreases, reflecting the strength of frequency information in identifying important words to retain during masking.
-5.2 Zero-shot performance on different data sets
+Recall that the reduction of the number of input tokens allows for an increase in the batch size.
+As we reduce the number of tokens from 8 to 6 to 4, syntax masking and truncation do not perform as well as the other text masking strategies.
+We see that compared to existing text matching approaches, CLIPF is better positioned to deliver benefits as the number of input tokens decreases.
+6.2 Zero-shot Robustness Evaluation
 Table 6: Zero-shot robustness evaluation.
@@
 CLIPF frequency 7.05 32.40 39.90 18.94 26.72 17.14
-Following the methodology outlined in CLIP and FLIP [3, 8], we evaluated various masking strategies across data sets to identify their robustness.
+Following the evaluation of CLIP and FLIP [3, 8], we test the masking strategies across data sets.
 Table 6 presents results for zero-shot classification for six commonly used data sets.
@@
 Again, we see that the benefits of CLIPF are particularly evident as the masking number of input tokens decreases.
-In other words, we observe that the gap in performance widen between CLIP and the other strategies as the masking ratio increases.
 We also observe that CLIPF is able to improve over the FLIP baseline without masking in a majority of cases.
-However, despite being the strongest text-matching strategy, the results reveal that it is challenging for CLIPF to improve over the original CLIP.
-In only a handful of cases does it outperform the original CLIP.
-However, it is important to recall that training with full images, rather than using image masking, can also be combined with text masking.
-We did not test this setting here for computational reasons, but depending on the computational resources in an application setting it could be relevant.
+However, despite being the strongest text-masking strategy, the results reveal that it is challenging for CLIPF to improve over the original CLIP.
+In only a handful of cases does it outperform the original CLIP, but might do so with more image tokens as we saw in 6.1.
 Table 7: Zero-shot Image-Text Retrieval.
-We evaluated CLIP, FLIP, CLIPA, and CLIPF image-text retrieval performance on COCO and Flickr30k datasets.
+We evaluated CLIP, FLIP, CLIPA, and CLIPF image-text retrieval performance on the COCO and Flickr30k datasets.
 The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
@@
 syntax 8 50.30 29.64 38.46 20.28
-CLIPF frequency 8 60.36 35.28 44.81 23.66
+CLIPF frequency 8 58.68 34.74 44.58 23.31
 CLIPA truncation 6 32.25 16.18 22.39 11.93
@@
 Table 7 shows image/text retrieval results on Flickr30k [37] and COCO [38].
-These results demonstrate the ability of text masking to improve the performance of a VLM.
-The performance improvement is particularly notable in comparison to FLIP, which like our CLIPA and CLIPF uses a 75% image-masking ration.
-When using 16 text tokens, all strategies except the truncation method outperform FLIP which is pre-trained on the full text.
-The performance is also quite strong in comparison to CLIP, which uses no masking.
-When using 16 text tokens, text masking strategies, except for truncation, are able to approach and in a few cases beat the accuracy of CLIP.
+In general, we see that text masking approaches are able to improve over FLIP.
+When using 16 text tokens, all strategies except the truncation method outperform FLIP, which is pre-trained on the full text.
+Here, we do not see consistent improvement over CLIP, although some cases come lose.
+Note again that we are testing here only with 49 image tokens and 98 would fare better, with additional computational cost.
 In contrast to the previous experiments, with image-text retrieval, we observe that CLIPF does not clearly dominate the other text masking strategies.
-The CLIPA random, CLIPA block, and CLIPF (frequency) strategies exhibit comparable performance in image and text retrieval across both datasets, but CLIPF does not consistently outperform the other masking strategies.
-The performance of the syntax strategy is similar to that of FLIP but is lower than that seen with the random, block, and CLIPF.
+The CLIPA random, CLIPA block, and CLIPF (frequency) strategies exhibit comparable performance in image and text retrieval across both datasets.
 The notable exception is the performance of CLIPF in the case that the text masking ratio is very high, i.e., only 4 input text tokens are used.
 Here, the benefit of CLIPF’s frequency-based approach is clearly evident.
-6 Analysis and Further Experiments
-6.1 Word Distribution Analysis
-Figure 2: The figure illustrates the distribution of top-50 words in the text before and after applying various text masking strategies, we set the text length after text masking to 6.
-The x-axis represents the word index which is sorted by counts of the original data, and the y-axis shows the word frequency.
-The dataset used is CC12M and the value of $t$ of formular 2 is set to $10^{-6}$.
-We remove special characters from the vocabulary.
-In this section, we provide an analysis of how different text masking strategies affect the distribution of words By looking at the differences in distribution we can gain insight into why some text-masking strategies work well, and some less well, and, in particular, why CLIPF delivers the best performance.
-Fig.
-2 shows the distribution of words in the text before and after applying various text masking strategies.
-We see that in general there is quite a large difference between the frequency of the highest-frequency words and the frequency of words with lower frequency.
-Such a long-tail distribution is characteristic of words occurring in text, which is known to be Zipf distributed.
-However, we also observe differences between the different text masking approaches.
-The truncation strategy leads to frequent words being more frequent than they are with the other strategies.
-This effect might seem surprising at first, but can be explained if we consider that texts typically start with the same words, e.g., “A person is…” We conjecture that truncation does not perform well as a text-masking strategy because the VLM overfits to these words.
-The CLIPA random and block strategies exhibit nearly identical distributions, since each word has a 50% probability of being masked in both cases.
-They lie somewhat below the CLIPA truncation line, and we describe them as being more “balanced” not in the sense that the distributions are flat, but rather that the differences between the frequencies of very high-frequency words and the of the rest of the words is less extreme.
-Even more balanced is the CLIPF distribution, which lies below CLIPA random and block.
-The distribution of the CLIPA syntax strategy is quite different.
-Because this strategy retains more nouns, the articles “the” and “a” or less frequent, but the nouns create spikes.
-The rationale for prioritizing nouns is convincing.
-Nouns are closely related to the depicted content of images and for this reason are considered to be more helpful.
-Indeed, the helpfulness of nouns for the VLM is reflected in the fact that Syntax masking learns faster than other text masking strategies.
-In Fig.
-1 we saw that syntax masking outperforms the other masking approaches during the early epochs and CLIPA [12] found that syntax masking was the best approach, while testing on a limited number of epochs.
-However, we believe that retaining nouns can have two adverse effects.
-First, the model can overfit to certain high frequency nouns.
-In Fig.
-2, we see that CLIPA syntax retains more occurrences of “person,” “illustration,” and “stock” than the other strategies.
-Second, retaining nouns means necessarily dropping words from other part-of-speech (POS) categories.
-To dive deeper into the importance of the distribution of words over POS categories, we provide statistics for each of the different text masking strategies in Table 8.
-Consulting the column ‘Other’ we see and interesting pattern.
-CLIPA syntax masking changes the proportion of words in the “Other” category radically with respect to the data before masking, but CLIPA truncation, random and block do not.
-CLIPF appears to provide a happy medium.
-CLIPF retains more nouns than CLIPA truncation, block, and syntax, but not so many that the “Other” category becomes poorly represented.
-We believe that it is this balance that this balance is an important factor in CLIPF’s strong performance.
-Table 8: To examine the distribution of words belonging to different POS categories, we calculate the percentage of words in the training data belonging to each category in before and after applying text masking.
-The dataset is CC12M, and we retain 6 words for each text after applying text masking.
-Masking NN (%) JJ (%) VB (%) OTHER (%)
-Before masking 50.30% 4.98% 5.18% 39.55%
-Truncation 51.72% 5.17% 5.91% 37.20%
-Random 51.21% 4.82% 5.04% 38.93%
-Block 49.37% 4.74% 5.24% 40.66%
-Syntax 88.53% 2.69% 2.92% 5.87%
-CLIPF 60.43% 5.06% 6.20% 28.24%
-6.2 Learning Curve
-Next, we return to consideration of the learning curve that compares CLIPF with the other text-masking strategies, which is shown in Fig.
-1.
-As already mentioned, syntax is a strong performer in the earlier epochs, until epoch 20 remains the best of the CLIPA approaches.
-We conjecture that CLIPA block and CLIPA random pull ahead of syntax because they retain “Other” words (i.e., words that are not nouns, adjectives or verbs) and do not overprioritize nouns.
-The importance of “Other” words is surprising because the category includes many words, which like articles, are said to lack semantic content and have no direct connection to images.
-In fact, in many NLP applications, articles and other “stop words” are simply removed.
-We believe that the transformer that CLIP uses to embed text is actually able to exploit the context of content words, particularly words in “Other” category, given enough epochs.
-CLIPA block masking retains this context a little better than random masking.
-Surprisingly, the frequency masking of CLIPF also retains this context.
-Under CLIPA syntax masking a phrase like “light in the house” would always be reduced to “light house”, but with CLIPF frequency masking the reduction would not happen in every case.
-In short, CLIPF achieves a productive balance in POS categories.
-6.3 Comparison with SW-CLIP
-In this section, we provide a comparison of CLIPF with SW-CLIP [11], which also uses frequency-based sampling but imposes a threshold on the frequency score.
-The effect of this threshold is that input tokens will go unused if not enough words in a training text surpass the threshold.
-In contrast, CLIPF calculates the word frequency by the threshold to prioritize words and maximize the input slots.
-For instance, using CC3M as an example, the average length of text before masking is 10.31 ± 4.7, and after SW-CLIPF masking, it is 4.26 ± 2.6 [11].
-It is clear that the use of the input slots is not maximized when the input length of text tokens is set to 6 or longer.
-For a fair comparison with CLIPF, we pre-train SW-CLIP using the same setup: 75% image masking and 81.25% text masking.
-Subsequently, we fine-tuned both models with an additional epoch without any image or text masking.
-The results in Table 9 show that after fine-tuning CLIPF outperforms SW-CLIP.
-Table 9: Comparison of the zero-shot accuracy performance of SW-CLIP and CLIPF on ImageNet-1k.
-Models Datasets Image Tokens Text Tokens ViT-B/16
-pre-train fine-tune
-SW-CLIP CC3M 49 6 14.9 16.8
-CLIPF 14.4 18.2
-SW-CLIP CC12M 49 8 35.9 38.5
-CLIPF 36.6 39.3
-Table 10: We pre-train CLIPF on CC12M dataset across different thresholds in Equation 2.
+We point out that for Image-Text Retrieval the performance of syntax masking is particularly low in comparison to the other forms of masking.
+Specifically, when using 8 and 6 text tokens, the performance gap between syntax masking and block, random, and frequency masking is approximately 9% on COCO and 6% on Flickr30k for image-text retrieval.
+This gap is wider than what was observed with Zero-shot Classification.
+We conjecture that the bad performance is due to the Image-Text Retrieval task requiring understanding the overall meaning and logic of the text rather than just recognizing nouns or objects.
+For this reason, the fact that syntax masking retains a high number of nouns, as shown in Section 5, is particularly harmful.
+Table 8: Zero-shot accuracy on more classification datasets.
+Comparison of the zero-shot accuracy performance of CLIP, FLIP, CLIPA, and CLIPF on various datasets.
 The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
-model Image tokens Text tokens Threshold ViT-B/16
-pre-train fine-tune
-CLIPF 49 8 1e-5 35.6 38.4
-1e-6 36.6 39.3
-1e-7 36.1 38.6
-6.4 Threshold Analysis
-To investigate the impact of the threshold $t$ in Equation 2 on model performance, we pre-trained models with varying thresholds.
-As shown in Table 10, thresholds of $51e-$ 5, $61e-$ 6, and $71e-$ 7 yield comparable results, all outperforming the other text masking strategies.
-This indicates that CLIPF is relatively insensitive to small variations in thresholds, which are intended to maintain the word masking probability within the range of 0 to 1.
-However, using too small a threshold reduces the differences in text masking probabilities; therefore, we do not recommend using an excessively low threshold.
-7 Conclusion and Outlook
-In this paper, we have introduced CLIPF an approach for text masking based on word frequency that improves the performance of CLIP.
-Unlike syntax-based masking methods, CLIPF does not require part-of-speech (POS) information, which is time-consuming to process.
-Unlike truncation, random, and block text masking strategies, which treat all words the same, CLIPF assigns different probabilities based on word frequency.
-In the zero-shot transfer evaluation, CLIPF demonstrates a notable ability to rapidly learn image and text representations when compared to a model that pre-trains on the full text and the other text masking strategies.
-Finally, we analyze the word distribution before and after applying text masking to provide insight into why CLIPF is effective.
-We establish that CLIPF balances retention of nouns, similar to syntax masking, with retention of other parts of speech, similar to block and random masking, effectively providing the best of both worlds.
-Future should investigate two directions.
-First, CLIPF should be tested on larger data sets, that are beyond the reach of our own computational resources.
-The fact that CLIPF is effective both for CC3M as well as for the larger CC12M data set suggests that CLIPF scales well.
-We believe that CLIP could be adapted to any larger-scale image-text dataset; however, this should be shown empirically.
-We see no intrinsic reason why the advantages of CLIPF should not extend to large data sets, in which the frequency of the most frequent words can be expected to be even larger than for small datasets.
-Second, now that we have gained insights into the distributional properties of text data that appear important for successful text masking, future work can attempt to optimize masking directly to achieve these properties.
-In the meantime, frequency is what you need.
-Appendix A APPENDIX
-Appendix B Text Masking Analysis
-B.1 Text Masking Cases
-Table  11 presents the potential text resulting from the text masking technique.
-Since truncation is fixed in each epoch, it may result in the loss of important information at the end of the text.
-Certain words tend to appear in specific positions within the text; for example, “the” and “a” are most likely to be the first words of a sentence.
-As shown in Figure 2, truncation retains more occurrences of “the” and “a” than other text masking strategies.
-In contrast, random and block strategies can generate different texts in each epoch for text data augmentation, but they may retain some words with a high frequency.
-Syntax masking retains the nouns in the text and remains the same in each epoch.
-However, this strategy may cause the model to overfit on the frequency of noun words.
-In contrast, CLIPF varies the text in each epoch, as words may be retained or removed according to their frequency.
-This strategy serves two primary purposes: it enhances text diversity and reduces the risk of overfitting to frequent words.
-Another advantage of CLIPF is that it can remove frequent prepositions that are less directly relevant to the objects in the image, such as “a,” “in,” “of,” and others.
-This helps the model focus on the most helpful aspects of the content.
-Table 12 shows the masking probabilities for certain words.
-High-frequency words such as “of" “the" “and" and “." are highly likely to be masked from the text.
-Table 11: The example provided showcases from the CC12M dataset before and after text masking is applied, reducing the text from 32 tokens to 8 tokens.
-The first row displays the original text, while the subsequent rows show potential outcomes after applying text masking.
-The value of $t$ of formular 2 is set to $10^{-6}$.
-Text Walk of the happy young couple and Siberian dog. The handsome man is hugging the smiling red head girl
-Truncation walk of the happy young couple
-Random the happy and the is the
-Block couple and siberian dog . the
-Syntax walk dog man smiling head girl
-CLIPF siberian handsome man hugging smiling red
-Table 12: The probability of masking words in the text is calculated using the formula provided in Equation 2.
-The example is selected from CC12M [15].
-The value of $t$ of formular 2 is set to $10^{-6}$.
-Word Probability
-walk 0.926171
-of 0.992838
-the 0.995064
-happy 0.951695
-young 0.957311
-couple 0.941174
-and 0.991695
-siberian 0.75092
-dog 0.960531
-. 0.993678
-B.2 Word Distribution Analysis for CLIPF
-In addition, we analyzed frequency text masking strategies that varied according to the number of words used, as shown in Fig.
-3.
-As the number of words decreased from 16 to 8 and then to 6, more frequent words were masked.
-However, reducing the number to 4 words led to a smaller vocabulary size, resulting in the loss of some important information.
-Consequently, the performance of the model pre-trained with 4 words was substantially lower compared to the model pre-trained with 6 words.
-We recommend setting the number of text words in a frequency-based text masking strategy to strike a balance between frequent and infrequent words and to maintain a larger vocabulary.
-Based on our experiments, the optimal number of text masking was found to be approximately 40-60% of the average length of the original text.
-This configuration helps achieve a balanced word distribution, which is beneficial for pre-training VLMs.
-Figure 3: Frequency-based text masking strategies vary according to the number of text words used during pre-training. The dataset used is CC12M and the value of $t$ of formular 2 is set to $10^{-6}$.
-B.3 Word Distribution of CLIPF and SW-CLIP
-We also compare the word distributions between CLIPF and SW-CLIP. As shown in Figure 4, SW-CLIP exhibits a flatter word distribution than CLIPF.
-Figure 4: The figure illustrates the distribution of top-50 words in the text for SW-CLIP and CLIPF.
-B.4 Word Category Distribution Across Different Text Masking Strategies
-As shown in Table 13, we provide the word type counts of Table 8.
-After applying text masking strategies such as truncation, random, block, syntax, and CLIPF, the word counts remain similar.
-However, SW-CLIP does not maximize the use of the input slots, utilizing only 85% of the words compared to other text masking strategies.
-Additionally, SW-CLIP masks a high percentage of other type words, which may impact the model’s zero-shot ability.
-Table 13: Distribution of syntax counts and percentages before and after applying text masking. The dataset is CC12M and we retain 6 words for each text after applying text masking.
-Masking NN Count JJ Count VB Count OTHER Count NN (%) JJ (%) VB (%) OTHER (%) Total
-Before masking 103,469,117 10,245,828 10,649,943 81,351,966 50.30% 4.98% 5.18% 39.55% 205,716,854
-Truncation 28,628,129 2,859,462 3,272,401 20,585,364 51.72% 5.17% 5.91% 37.20% 55,345,356
-Random 28,334,735 2,666,847 2,790,600 21,550,517 51.21% 4.82% 5.04% 38.93% 55,342,699
-Block 27,314,723 2,624,594 2,897,434 22,510,031 49.37% 4.74% 5.24% 40.66% 55,346,782
-Syntax 48,989,317 1,483,666 1,618,088 3,245,892 88.53% 2.69% 2.92% 5.87% 55,336,963
-SW-CLIP 34,645,367 3,346,676 4,073,825 6,062,345 71.98% 6.95% 8.47% 12.60% 48,128,213
-CLIPF 33,516,439 2,803,409 3,473,119 15,666,310 60.43% 5.06% 6.20% 28.24% 55,459,277
-Figure  5 shows the word cloud before and after applying text masking.
-In these visuals, the font size represents the relative frequency of each word.
-Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency.
-After applying frequency word masking, the word cloud in Table 5(b) shows a more balanced distribution of word frequencies compared to before the masking.
-The word “the" appears smaller than in other word clouds.
-The reduction in frequent words allows infrequent words to become more prominent.
-The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full-text.
-As depicted in Table 5(f), syntax masking tends to retain more nouns such as “person," “stock," and “illustration" that show a different pattern from the others.
-Figure 5 shows word clouds before and after applying text masking.
-In these visuals, the font size represents the relative frequency of each word.
-Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency.
-After applying word frequency masking, the word cloud in Figure 5(b) displays a more balanced distribution of word frequencies compared to before masking.
-The word “the” appears smaller than in other word clouds.
-The reduction in frequent words allows infrequent words to become more prominent.
-The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full text.
-As depicted in Figure 5(f), syntax masking tends to retain more nouns such as “person”, “stock” and “illustration” resulting in a different pattern from the others.
-Refer to caption Refer to caption
-(a) Full Text (b) Frequency
-Refer to caption Refer to caption
-(c) Truncation (d) Random
-Refer to caption Refer to caption
-(e) Block (f) Syntax
-Figure 5: The word cloud on the left displays the full text, while the one on the right shows the text after applying text masking, where we retain 6 words.
-The font size of the word represents the relative word frequency for each figure.
-The dataset used is CC12M and the value of $t$ of formular 2 is set to $10^{-6}$.
-B.5 Bi-gram word Distribution
-As shown in Figure 6, we also calculated the bigram distribution of the text after applying different text masking strategies.
-Truncation and block masking better retain information about the order of the words in the text because the words remain contiguous after masking.
-Random and CLIPF masking also preserves the text order better than syntax masking.
-Syntax masking retains more continuous bigrams—such as “stock illustration,” “vector illustration,” and “stock photography”—than other text masking strategies.
-These bigrams might be domain specific and not contribute to the transferrability of the VLM to other data sets.
-Figure 6: The figure illustrates the distribution of top-50 bi-grams in the text with different text masking strategies.
-B.6 Impact of Symbolic Characters
-We note the high frequency of symbolic characters appearing in the dataset, these symbolic characters are not directly related to the depicted content of the images.
-Since CLIPF tends to mask these characters with high probability, we removed them during the pre-training of the model to evaluate their significance.
-As shown in Table 14, removing symbolic characters and stop words does not affect the model’s performance.
-Table 14: We remove symbolic characters and stop words during the model’s pre-training to evaluate their zero-shot classification performance on ImageNet-1K.
-The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC3M [23] for 30 epochs with image masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
-Models Masking Image Tokens Text Tokens Remove ViT-B/16
-pre-train fine-tune
-FLIP 49 32 14.0 14.2
-FLIP 49 32 13.9 13.7
-B.7 Word and Token Analysis
-Since CLIP encodes text using byte-pair encoding (BPE), some words are encoded into more than one token.
-Therefore, in this experiment, we calculate the masking probability based on tokens rather than words.
-As shown in Table 15, we compare the performance of both approaches.
-Table 15: Comparison of CLIPF pre-trained with token masking and word masking on ImageNet-1K for zero-shot classification.
-Models Masking Image Tokens Text Tokens ViT-B/16
-pre-train fine-tune
-CLIPF frequency-token 49 8 36.0 38.0
-frequency-word 49 8 36.6 39.3
-Table 16: Zero-shot accuracy on more classification datasets. Comparison of the zero-shot accuracy performance of CLIP, FLIP, CLIPA, and CLIPF on various datasets.
 Models Method
@@
 CLIPF frequency 4 36.09 69.76 31.14 8.41 47.72 7.01 2.24 25.64 54.23 64.55 23.55 2.95 10.18 90.25 20.78 26.52 6.90 36.63 6.40 55.71 35.71 12.65 49.63 50.08 30.93
-Appendix C More results
-More detailed results are presented in this section.
-C.1 Zero-shot Classification on More Datasets
-In Table 16, we present the evaluation results of CLIPF following the approach of CLIP [3] on additional datasets.
-In almost all of the downstream datasets, when using six text tokens, CLIPF surpasses not only the baseline CLIP without image and text masking and FLIP with image masking and full-text tokens but also outperforms other text masking strategies.
-Similar to observations from the ImageNet datasets, reducing the number of text tokens from 16 to 8 substantially enhances CLIPF’s performance across all datasets.
-Specifically, performance improvements are noted: SUN397 [38] by 6.39%, DTD [39] by 9.31%, OxfordPets [40] by 4.64%, STL10 [41] by 9.31, EuroSAT [42] by 6.60%, UFC101 [43] by 6.68% and ImageNet-1K [44] by 3.13%.
+6.4 Zero-shot Classification on More Datasets
+In Table 8, we present evaluation results on additional datasets.
+With six text tokens, CLIPF consistently outperforms the original CLIP (no masking), FLIP (image masking with full-text), and other text masking strategies.
+Aligned with ImageNet findings, reducing text tokens from 16 to 8 significantly boosts performance.
+Improvements include SUN397 [39] (+6.39%), DTD [40] (+9.31%), OxfordPets [41] (+4.64%), STL10 [42] (+9.31%), EuroSAT [43] (+6.60%), UCF101 [44] (+6.68%).
 The performance of other text masking strategies, including truncation, random, block, and syntax, also improves with this reduction.
-However, the performance gains are not as large as those observed with CLIPF.
-This improvement suggests that employing text masking strategies as a data augmentation technique can enhance VLMs’ performance in zero-shot classification tasks.
-CLIPF, utilizing a word frequency masking method, demonstrates robust zero-shot classification performance and establishes a benchmark for state-of-the-art results.
-C.2 Zero-shot Image-text Retrieval
-We report the details of zero-shot image-text retrieval for Flickr30k and COCO datasets, the details are shown in Table 17.
+However, their improvements are more limited.
+We point out that the Describable Textures Dataset (DTD)[40] contains 47 categories of texture patches, whose POS is mostly verb and that the OxfordPets dataset[41] also contains many class names that are not nouns.
+On these data sets the performance of syntax masking is particularly low.
+Specifically, when using six text tokens, syntax masking underperforms CLIPF by 14.10% on DTD and 21.04% on OxfordPets.
+This low performance reflects that the VLM has overfitted on the large number of nouns retained by syntax masking.
+6.5 Scaling to the larger Dataset.
+Finally, we carry out an exploratory experiment to gain insight into different masking strategies with data sets of different sizes.
+We carry out the largest experiment that is possible given our computational capacity, which was training on the LAION-400M [17] data set, using 4 text tokens, a smaller image size (112x112) and 50% image masking.
+We trained for 16 epochs.
+Previous works have shown that both smaller image sizes and image masking are effective for accelerating CLIP pre-training [21, 8, 45].
+Results are shown in Fig.
+4 for both top-1 accuracy (left) and top-5 accuracy (right).
+We see from these graphs that using $<$ 10 epochs would not provide a complete picture on how the different masking approaches perform relative to each other.
+Overall, CLIPF, the frequency based approach dominates.
+However, syntax is performing better than it did on smaller data sets, possibly because the large amount of data is compensating for the overemphasis of syntax masking on nouns.
+It is important not to conclude on the basis of this experiment that syntax masking will continue to dominate block masking.
+Looking at the top-5 performance, we see that these two approaches near each other already in the 16 epochs we are able to compute.
+In sum, this experiment shows that CLIPF is a good performer as data scales up, but that the number of epochs is important in understanding the strength of different masking strategies.
+Figure 4: Zero-shot classification accuracy on ImageNet-1K over training epochs for CLIPF and CLIPA strategies when pre-training on LAION-400M datasets.
+We pre-trained the model using 25 image tokens and 4 text tokens for 16 epochs.
+7 Conclusion and Outlook
+This paper has provided a systematic overview of text masking approaches for VLM training, which accelerate training and have the potential to also improve performance.
+In this paper, we demonstrate that training epochs have a substantial impact on the performance of different text masking strategies.
+We conclude, contrary to the existing state-of-the-art, CLIPA [12], that syntax masking is not the best performing text masking approach and that truncation masking can be avoided.
+We carry out an analysis of word frequency distributions that shows that frequency is what you need to think about when text masking the training data of VLMs.
+Specifically, both syntax and truncation have visible effects on the word frequency distribution, making clear why they do not perform well.
+We have introduced CLIPF, an approach for text masking based on word frequency that improves the performance of CLIP.
+Unlike syntax-based masking methods, CLIPF does not require part-of-speech (POS) information, which is time-consuming to process.
+CLIPF is also interesting because it controls frequency directly and in doing so seems to balance nouns with words in the “Other” POS categories, in contrast to other approaches.
+We carry out an extensive comparison among existing text matching strategies (CLIPA variants) and CLIPF that revals the strong performance of CLIPF in a variety of zero-shot classification tasks and when scaling to larger data sets.
+CLIPF demonstrates a notable ability to rapidly learn image and text representations when compared to a model that pre-trains on the full text and other text masking strategies.
+It also delivers strong performance in under highly aggressive masking (i.e., using only 4 tokens per training text).
+Moving forward, the insights gained from this paper about the importance of frequency can be used to further optimize CLIPF to directly optimize the word distribution allowing for more tuning control when trading off VLM performance and VLM training speech during text masking.
+Appendix A Text Masking Analysis
+A.1 Text Masking Cases
+As illustrated in Fig.
+5, there is a clear relationship between word masking probability and word frequency.
+Frequent words have a higher masking probability compared to infrequent ones.
+Additionally, a smaller threshold $t$ leads to a smaller difference between the masking probabilities of frequent and infrequent words.
+Therefore, it was necessary to choose a relatively larger threshold to ensure that both frequent and infrequent words are effectively masked.
+Figure 5: The curve of Formula 2. The x-axis is the word frequency $f(w_{i})$, and the y-axis is the $P(w_{i})$. The value of $t$ of formular 2 is set to $10^{-5}$, $10^{-6}$, $10^{-7}$,$10^{-8}$.
+Table  9 presents the potential text resulting from the text masking technique.
+Since truncation is fixed in each epoch, it may result in the loss of important information at the end of the text.
+Certain words tend to appear in specific positions within the text; for example, “the” and “a” are most likely to be the first words of a sentence.
+As shown in Fig.
+3, truncation retains more occurrences of “the” and “a” than other text masking strategies.
+In contrast, random and block strategies can generate different texts in each epoch for text data augmentation, but they may retain some words with a high frequency.
+Syntax masking retains the nouns in the text and remains the same in each epoch.
+However, this strategy may cause the model to overfit on the frequency of noun words.
+In contrast, CLIPF varies the text in each epoch, as words may be retained or removed according to their frequency.
+This strategy serves two primary purposes: it enhances text diversity and reduces the risk of overfitting to frequent words.
+Another advantage of CLIPF is that it can remove frequent prepositions that are less directly relevant to the objects in the image, such as “a,” “in,” “of,” and others.
+This helps the model focus on the most helpful aspects of the content.
+Table 10 shows the masking probabilities for certain words. High-frequency words such as “of" “the" “and" and “." are highly likely to be masked from the text.
+Table 9: Example from the CC12M dataset illustrating the effect of various text masking strategies, reducing text length to 6 words.
+The original caption is in the first row, followed by masked variants.
+Parameter $t$ in Equation 2 is set to $10^{-6}$.
+Method Resulting Text
+Original Walk of the happy young couple and Siberian dog. The handsome man is hugging the smiling red head girl
+Truncation walk of the happy young couple
+Random the happy and the is the
+Block couple and siberian dog . the
+Syntax walk dog man smiling head girl
+CLIPF siberian handsome man hugging smiling red
+Table 10: The probability of masking words in the text is calculated using the formula provided in Equation 2.
+The example is selected from CC12M [15].
+The value of $t$ of formular 2 is set to $10^{-6}$.
+Word Probability
+walk 0.926171
+of 0.992838
+the 0.995064
+happy 0.951695
+young 0.957311
+couple 0.941174
+and 0.991695
+siberian 0.75092
+dog 0.960531
+. 0.993678
+A.2 Word Category Distribution Across Different Text Masking Strategies
+As shown in Table 11, we provide the word type counts of Table 5.
+After applying text masking strategies such as truncation, random, block, syntax, and CLIPF, the word counts remain similar.
+However, SW-CLIP does not maximize the use of the input slots, utilizing only 85% of the words compared to other text masking strategies.
+Additionally, SW-CLIP masks a high percentage of other types of words, which may impact the model’s zero-shot ability.
+Table 11: Distribution of syntax counts and percentages before and after applying text masking. The dataset is CC12M and we retain 6 words for each text after applying text masking.
+Masking NN Count JJ VB OTHER NN (%) JJ (%) VB (%) OTHER (%) Total
+Before masking 103,469,117 10,245,828 10,649,943 81,351,966 50.30% 4.98% 5.18% 39.55% 205,716,854
+Truncation 28,628,129 2,859,462 3,272,401 20,585,364 51.72% 5.17% 5.91% 37.20% 55,345,356
+Random 28,334,735 2,666,847 2,790,600 21,550,517 51.21% 4.82% 5.04% 38.93% 55,342,699
+Block 27,314,723 2,624,594 2,897,434 22,510,031 49.37% 4.74% 5.24% 40.66% 55,346,782
+Syntax 48,989,317 1,483,666 1,618,088 3,245,892 88.53% 2.69% 2.92% 5.87% 55,336,963
+SW-CLIP 34,645,367 3,346,676 4,073,825 6,062,345 71.98% 6.95% 8.47% 12.60% 48,128,213
+CLIPF 33,516,439 2,803,409 3,473,119 15,666,310 60.43% 5.06% 6.20% 28.24% 55,459,277
+Figure  6 shows the word cloud before and after applying text masking.
+In these visuals, the font size represents the relative frequency of each word.
+Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency.
+After applying frequency word masking, the word cloud in Table 6(b) shows a more balanced distribution of word frequencies compared to before the masking.
+The word “the" appears smaller than in other word clouds.
+The reduction in frequent words allows infrequent words to become more prominent.
+The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full-text.
+As depicted in Table 6(f), syntax masking tends to retain more nouns such as “person," “stock," and “illustration" that show a different pattern from the others.
+Fig.
+6 shows word clouds before and after applying text masking.
+In these visuals, the font size represents the relative frequency of each word.
+Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency.
+After applying word frequency masking, the word cloud in Fig.
+6(b) displays a more balanced distribution of word frequencies compared to before masking.
+The word “the” appears smaller than in other word clouds.
+The reduction in frequent words allows infrequent words to become more prominent.
+The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full text.
+As depicted in Fig.
+6(f), syntax masking tends to retain more nouns such as “person”, “stock”, and “illustration”, resulting in a different pattern from the others.
+Refer to caption Refer to caption
+(a) Before masking (b) Frequency
+Refer to caption Refer to caption
+(c) Truncation (d) Random
+Refer to caption Refer to caption
+(e) Block (f) Syntax
+Figure 6: The word cloud on the left displays the full text, while the one on the right shows the text after applying text masking, where we retain 6 words.
+The font size of the word represents the relative word frequency for each figure.
+The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$.
+A.3 Word Distribution Analysis for CLIPF
+In addition, we analyzed frequency text masking strategies that varied according to the number of words used, as shown in Fig.
+7.
+As the number of words decreased from 16 to 8 and then to 6, more frequent words were masked.
+However, reducing the number to 4 words led to a smaller vocabulary size, resulting in the loss of some important information.
+Consequently, the performance of the model pre-trained with 4 words was substantially lower compared to the model pre-trained with 6 words.
+We recommend setting the number of text words in a frequency-based text masking strategy to strike a balance between frequent and infrequent words and to maintain a larger vocabulary.
+Based on our experiments, the optimal number of text masking was found to be approximately 40-60% of the average length of the original text.
+This configuration helps achieve a balanced word distribution, which is beneficial for pre-training VLMs.
+Figure 7: Frequency-based text masking strategies vary according to the number of text words used during pre-training. The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$.
+A.4 Word Distribution of CLIPF and SW-CLIP
+We also compare the word distributions of CLIPF and SW-CLIP after applying text masking.
+As illustrated in Fig.
+8, SW-CLIP masks more frequent words based on its threshold, preventing it from fully utilizing the input slots of the text encoder.
+Consequently, SW-CLIP exhibits a flatter word distribution compared to CLIPF.
+Figure 8: The figure illustrates the distribution of top-50 words in the text for SW-CLIP and CLIPF.
+Appendix B More Results
+More detailed results are presented in this section, including the results and learn curve of image-text retrieval.
+B.1 The details of Image-text Retrieval
+In Table 12, we show more details of Zero-shot Image-Text Retrieval on COCO and Flickr30k datasets.
 Table 12: Zero-shot Image-Text Retrieval, we evaluated CLIP, FLIP, and CLIPF image-text retrieval performance on COCO and Flickr30k datasets.
@@
 CLIPA truncation 16 44.67 73.08 81.85 25.54 51.90 64.64 34.99 61.05 70.85 19.64 41.91 53.51
-random 16 58.48 83.93 90.53 32.36 58.76 69.98 43.61 70.67 80.04 22.81 46.00 57.71
+random 16 58.48 84.62 90.53 32.36 58.76 69.98 43.61 70.67 80.04 22.81 46.00 57.71
 block 16 56.51 81.26 89.05 30.82 58.32 70.38 44.06 71.20 80.10 22.66 46.24 58.06
@@
 syntax 8 50.30 78.30 86.98 29.64 55.42 67.18 38.46 66.77 77.36 20.28 43.78 55.40
-CLIPF frequency 8 60.36 85.80 90.63 35.28 62.70 73.70 44.81 72.07 81.34 23.66 48.05 59.89
+CLIPF frequency 8 58.68 85.10 91.51 34.74 61.38 71.88 44.57 72.58 81.92 23.30 47.50 59.24
 CLIPA truncation 6 32.25 61.05 72.98 16.18 39.52 52.48 22.39 47.46 59.39 11.93 30.03 40.86
@@
 CLIPF frequency 4 45.07 72.49 81.95 25.60 49.98 61.96 31.72 59.35 71.66 15.96 36.58 48.32
-C.3 Learn Curved before fine-tuning
-In Figure 1, we show the learning curve after fine-tuning. Moreover, we show the learning curve before fine-tuning, they have a similar performance for every text masking strategy.
+B.2 Learn Curved for Image-text Retrieval
+We also investigate the zero-shot image-text retrieval performance over training epochs on the COCO and Flickr30k datasets.
+It also shows that the number of training epochs has a substantial impact on the performance, as illustrated in Fig.
+9.
+And the performance gap between syntax masking and the other text masking on the zero-shot image-text retrieval task is larger than zero-shot classification task.
+This is because the zero-shot image-text retrieval task requires the model not only to understand the objects/nouns, but also to understand the logic of the text.
+Moreover, we present the zero-shot image-text retrieval performance across training epochs for models pre-trained on the LAION-400M dataset.
+As illustrated in Fig.
+10, the performance patterns of models trained on LAION-400M are similar to those of previous models that pre-trained on the CC12M dataset.
+Specifically, CLIPF consistently outperforms syntax masking.
+Block masking achieves comparable performance to syntax masking and even surpasses it on the image-to-text retrieval task using the COCO dataset at epoch 16th.
+Figure 9: Zero-shot Image-Text Retrieval accuracy on COCO and Flick30k over training epochs for CLIPF and CLIPA strategies.
+The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model for an additional epoch without image and text masking.
+Figure 10: Zero-shot Image-Text Retrieval accuracy on COCO and Flick30k over training epochs for CLIPF and CLIPA strategies.
+The backbone of the image encoder is ViT-B/16, and the model is pre-trained on LAION-400M for 16 epochs with 25 image tokens and 4 text tokens to speed up training and fine-tune the model for 0.4 epoch with 224x224 image size and 32 text tokens.
+Additionally, Fig.
+11 illustrates the validation loss across pre-training epochs.
+It is evident that the validation loss for truncation and syntax masking methods begins to increase earlier compared to other text masking methods, suggesting that these methods may overfit to the data sooner.
+Figure 11: The validation loss curve of COCO and Flick30k datasets over the pre-training epochs.
+B.3 Learn Curved for DTD and OxfordPets Datasets
+The learning curves for the DTD [40] and OxfordPets [41] datasets are presented because not all class names in these datasets are nouns.
+As shown in Table 13, the performance of syntax masking on the DTD dataset is consistently lower than that of random, block, and frequency masking across all epochs, rather than initially outperforming them and declining at later stages in the ImageNet dataset, as shown in Fig.
+12.
+Moreover, the performance of syntax masking is very poor and almost close to truncation masking.
+OxfordPets is a dataset containing incomplete noun class names.
+In the early training epochs, syntax masking performs similarly to frequency masking and still outperforms truncation, block, and random masking.
+However, in the later stages of training, syntax masking performs much worse than the other text masking strategies.
+Figure 12: Zero-shot classification accuracy on ImageNet-1K dataset over training epochs for CLIPF and CLIPA strategies.
+The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
+Figure 13: Zero-shot classification accuracy on DTD dataset over training epochs for CLIPF and CLIPA strategies.
+The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
+Figure 14: Zero-shot classification accuracy on OxfordPets dataset over training epochs for CLIPF and CLIPA strategies.
+The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
+B.4 Learn Curved before fine-tuning
+In Fig.
+1, we show the learning curve after fine-tuning.
+Moreover, we show the learning curve before fine-tuning in Fig.
+15 (30 epochs) and Fig.
+16 (45 epochs).
+They have a similar pattern to the models before fine-tuning for every text masking strategy.
 Figure 15: Zero-shot classification accuracy on ImageNet-1K over the epochs for the different text masking strategies before fine-tuning.
+The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training.
+Figure 16: Zero-shot classification accuracy on ImageNet-1K over the 45 epochs for the different text masking strategies before fine-tuning.
+The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 45 epochs with image and text masking (75%) to speed up training.
+B.5 Fine-tuning the models with More Epochs
+The purpose of fine-tuning the models without image and text masking is to reduce the distribution gap between pre-training and evaluation.
+As demonstrated in FLIP and CLIPA, a small number of fine-tuning steps is usually sufficient to bridge this gap [8, 12].
+We conducted additional experiments using two fine-tuning epochs, shown in Tab 13, to demonstrate that increasing the number of fine-tuning epochs does not substantially improve performance but instead leads to higher computational costs.
+Table 13: Comparison of CLIPF, CLIP, and CLIPA for zero-shot classification on ImageNet-1K.
+Reported as top-1 accuracy.
+All models use a ViT-B/16 image encoder backbone pre-trained on CC12M for 30 epochs with 75% image masking for speedup (pre-train).
+Then, we fine-tune the model with one epoch (fine-tune-1) and two epochs (fine-tune-2) without image and text masking.
+Models Masking Image Tokens Text Tokens CC12M
+pre-train fine-tune-1 fine-tune-2
+CLIP 197 32 36.6
+FLIP 49 32 32.0 33.7 34.1
+CLIPA truncation 49 8 25.4 28.4 28.5
+random 34.5 36.9 37.1
+block 35.5 37.9 38.0
+syntax 28.5 35.0 35.5
+CLIPF frequency 36.6 39.3 39.4
+B.6 Comparison with SW-CLIP
+In this section, we provide a comparison of CLIPF with SW-CLIP [22], which also uses frequency-based sampling but imposes a threshold on the frequency score.
+The effect of this threshold is that input tokens will go unused if not enough words in a training text surpass the threshold.
+In contrast, CLIPF calculates the word frequency by the threshold to prioritize words and maximize the input slots.
+For instance, using CC3M as an example, the average length of text before masking is 10.31 ± 4.7, and after SW-CLIPF masking, it is 4.26 ± 2.6 [22].
+It is clear that the use of the input slots is not maximized when the input length of text tokens is set to 6 or longer.
+For a fair comparison with CLIPF, we pre-train SW-CLIP using the same setup: 75% image masking and 81.25% text masking.
+Subsequently, we fine-tuned both models with an additional epoch without any image or text masking.
+The results in Table 14 show that after fine-tuning CLIPF outperforms SW-CLIP.
+Table 14: Comparison of the zero-shot accuracy performance of SW-CLIP and CLIPF on ImageNet-1k.
+The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
+Models Datasets Image Tokens Text Tokens ViT-B/16
+pre-train fine-tune
+SW-CLIP CC3M 49 6 14.9 16.8
+CLIPF 14.4 18.2
+SW-CLIP CC12M 49 8 35.9 38.5
+CLIPF 36.6 39.3
+B.7 Applying text masking on SigLIP
+We also applied the text masking strategies to SigLIP; the results are presented in Tab.
+15.
+CLIPF still achieves superior performance compared to other text masking strategies and without text masking.
+This further supports our conclusion that frequency-based masking is the more effective strategy.
+Table 15: We apply text masking (75%) on SigLIP to pre-train on CC12M for 30 epochs with 75% image masking for speedup pre-training. Reporting the zero-shot classification accuracy on ImageNet-1K.
+Models Masking Image Tokens Text Tokens ViT-B/16
+pre-train fine-tune
+SigLIP 197 32 39.3
+SigLIP 49 32 28.4 29.6
+SigLIP truncation 49 8 20.4 27.1
+random 29.7 31.7
+block 30.5 32.7
+syntax 25.9 31.6
+SigLIP frequency 32.4 34.8
+Appendix C Ablation
+Table 16: We pre-train CLIPF on CC12M dataset across different thresholds in Equation 2.
+The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
+model Image tokens Text tokens Threshold ViT-B/16
+pre-train fine-tune
+CLIPF 49 8 1e-5 35.6 38.4
+1e-6 36.6 39.3
+1e-7 36.1 38.6
+C.1 Threshold Analysis
+To investigate the impact of the threshold $t$ in Equation 2 on model performance, we pre-trained models with varying thresholds.
+As shown in Table 16, thresholds of $51e-$ 5, $61e-$ 6, and $71e-$ 7 yield comparable results, all outperforming the other text masking strategies.
+This indicates that CLIPF is relatively insensitive to small variations in thresholds, which are intended to maintain the word masking probability within the range of 0 to 1.
+However, using too small a threshold reduces the differences in text masking probabilities; therefore, we do not recommend using an excessively low threshold.
+Table 17: Comparison of CLIPF pre-trained with token masking and word masking on ImageNet-1K for zero-shot classification.
+The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
+Models Masking Image Tokens Text Tokens ViT-B/16
+pre-train fine-tune
+CLIPF frequency-token 49 8 36.0 38.0
+frequency-word 49 8 36.6 39.3
+C.2 Word and Token Analysis
+Since Open_CLIP encodes text using byte-pair encoding (BPE) [9], some words are represented by multiple tokens.
+Therefore, in this experiment, we calculate the masking probability based on token frequency rather than whole words.
+Table 17 compares the performance of both approaches.
+Although token frequency masking may disrupt the original word structure and thus achieve lower performance compared to word frequency masking, it still slightly outperforms other text masking strategies.

### v3

@@
-Frequency Is What You Need: Word-frequency Masking Benefits Vision-Language Model Pre-training
+Frequency Is What You Need: Considering Word Frequency When Text Masking Benefits Vision-Language Model Pre-training
 Abstract
 Vision Language Models (VLMs) can be trained more efficiently if training sets can be reduced in size.
-Recent work has shown the benefits of masking text during VLM training using a variety of approaches: truncation, random masking, block masking and syntax masking.
-In this paper, we show that the best masking strategy changes over training epochs and that, given sufficient training epochs.
-We analyze existing text masking approaches including syntax masking, which is currently the state of the art, and identify the word frequency distribution as important in determining their success.
-Experiments on a large range of data sets demonstrate that syntax masking is outperformed by other approaches, given sufficient epochs, and that our proposed frequency-based approach, called Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF) has numerous advantages.
-The benefits are particularly evident as the number of input tokens decreases.
-Keywords Vision-Language Model  $\cdot$ Multimodal Data  $\cdot$ Word Frequency Masking.
+Recent work has shown the benefits of masking text during VLM training using a variety of strategies (truncation, random masking, block masking and syntax masking) and has reported syntax masking as the top performer.
+In this paper, we analyze the impact of different text masking strategies on the word frequency in the training data, and show that this impact is connected to model success.
+This finding motivates Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF), our proposed masking approach, which directly leverages word frequency.
+Extensive experiments demonstrate the advantages of CLIPF over syntax masking and other existing approaches, particularly when the number of input tokens decreases.
+We show that not only CLIPF, but also other existing masking strategies, outperform syntax masking when enough epochs are used during training, a finding of practical importance for selecting a text masking method for VLM training.
+Our code is available online.
 1 Introduction
@@
 However, pre-training CLIP models typically requires thousands of GPU-days [32, 15].
-To accelerate the pre-training process, FLIP, A-CLIP, and CLIPA use masking strategies for image and text tokens when pre-training CLIP models [8, 11, 12].
-In addition to the speed-up, masking makes it possible use large batch sizes, which enhances the benefits of contrastive learning and has been considered as a form of regularization, reducing the risk of overfitting [8].
+To accelerate the pre-training process, current work makes use of masking strategies for image and text tokens when pre-training CLIP models [22, 43, 20].
+In addition to the speed-up, masking makes it possible to use large batch sizes, which enhances the benefits of contrastive learning and has been considered as a form of regularization, reducing the risk of overfitting [22].
 For these reasons, despite the loss of some information from the original image and text, masking strategies have the potential to improve model performance.
-Figure 1: Zero-shot classification accuracy on ImageNet-1K over training epochs for CLIPF and CLIPA strategies.
+Figure 1: Zero-shot classification accuracy on ImageNet-1K using models trained on CC12M to which different text masking strategies have been applied.
 The backbone of the image encoder is ViT-B/16.
-The models are pre-trained on CC12M for 45 epochs with 75% text masking ratio.
+The text masking ratio is 75%.
 We also use image masking (75%) to speed up pre-training.
-All approaches are subsequently fine-tuned with one epoch on the full, unmasked data.
-Currently, the state of the art in text masking is represented by CLIPA [12], which compares four masking strategies: truncation, random word selection, random block selection, and syntax-based selection, i.e., masking on the basis of part-of-speech (POS) information.
-The conclusion is that syntax masking is the strategy that, overall, provides the best quality.
-In this paper, we carry out a set of experiments that demonstrate that this conclusion must be revisited.
-Specifically, as shown in Fig.
-1, if enough epochs are used, other forms of masking substantially outperform syntax masking.
-The conclusion of CLIPA [12] misses the weakness of syntax masking because it trains with a limited number of epochs (6.4 epochs).
-The special focus of our paper is a frequency masking approach that we introduce: CLIPF (CLIP Frequency masking).
-In Fig.
-1, we see that with enough epochs CLIPF and CLIPA block masking perform similarly.
-We provide an analysis that demonstrates why CLIPF, block masking, and random masking emerge as clearly superior to syntax masking and truncation.
-The analysis studies the distribution of words by frequency of occurrence in the training data after masking.
-It shows that good masking strategies maintain helpful properties of this distribution and that syntax masking distorts the distribution of words over POS categories.
-These observation leads to the main insight of our paper, namely, when masking text for VLM pre-training, what you need for best results is to pay attention to the impact of masking on word frequency.
-Our CLIPF approach is useful in practice because it performs relatively well in the range of 16-40 epochs.
-This range is a typical number of epochs to train a system in practice.
-For example, Open_CLIP [9] trains on 30 epochs.
-In Fig.
-1, we see that in this range CLIPF outperforms other masking strategies.
-We carry out a set of experiments with 30 epochs to shed further light on the role of frequency in masking by comparing CLIPF vs.
-the different variants of CLIPA masking on zero-shot tasks and on image retrieval.
-We also investigate training on a larger data set.
-The history of text masking in language model training dates to [13], where subsampling of frequent words in the training text was used to improve the performance of skip-gram language models.
-The study of text masking for VLMs got off to a slow start, since initially [8] reported that text masking, although it promotes efficiency, deteriorated performance.
-The CLIPA [12] work is important because it expanded the types of masking studied and demonstrated that masking can indeed improve VLM performance.
-To our knowledge, no previous work has attempted to discover the reason that certain forms of text masking outperform others or identified the importance of the way in which text masking approaches impact word frequency.
-In sum, our paper makes the following contributions to the study of text masking in pre-training VLMs:
+At each point, we apply an additional epoch of training on the full, unmasked data.
+Currently, the state-of-the-art in text masking is represented by CLIPA [20].
+The authors compare two strategies originally introduced by FLIP [22], truncation and random word selection, and two new approaches, random block selection and syntax-based selection, i.e., masking on the basis of part-of-speech (POS) information.
+In [20], syntax masking is found to deliver the highest performance.
+In this paper, we carry out an analysis of word frequency in the training data before and after text masking and discover a connection between the impact of a masking strategy on the training data and the success of a VLM.
+As shown in Fig.
+1, our experiments reveal that syntax masking falls behind other text masking strategies (random and block) as the number of epochs increases.
+Our experiments establish two important conclusions: First, syntax masking is not the best performing text masking approach, updating the state-of-the-art finding in [20].
+Second, considering word frequency allows us to explain why random and block masking outperform truncation and syntax masking.
+This analysis motivates Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF), a new text masking strategy that directly leverages word frequency proposed in this paper.
+Because CLIPF leverages word-frequency directly, it does not need to expend compute on POS tagging, necessary for syntax masking.
+Fig.
+1 shows that CLIPF outperforms syntax masking.
+CLIPF is inspired by earlier training of text models, that has leveraged frequency masking [29].
+The key algorithmic novelty of CLIPF is the use of frequency-based word scores to prioritize words for masking, rather than to impose a hard decision on word removal.
+An important aspect of  Fig.
+1 is the development of performance over epochs.
+In contrast to the experiments in FLIP [22] and CLIPA [20], we continue training until additional epochs yield only minimal gains.
+Fig.
+1 demonstrates that our CLIPF approach is useful because it is the top performer in the range of 16-40 epochs, typically used in practice (e.g., OpenCLIP [15] trains on 30 epochs.)
+In sum, our paper makes the following contributions:
+  • We show that considering the impact of text masking on word frequency in the training data differentiates between successful and less successful text masking strategies.
+  • We introduce Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF), which leverages frequency directly during masking, and perform an in-depth quantitative comparison among existing masking strategies and CLIPF.
   • We point out the importance of the number of training epochs when comparing text masking strategies.
-  • We show that word frequency distribution after masking differentiates between successful and less successful text masking strategies.
-  • We carry out an in-depth comparison between existing masking strategies (CLIPA variants) and an approach that carries out masking by using frequency information directly, CLIPF.
-Our experiments show that CLIPF consistently outperforms syntax masking, previously considered the best choice for masking and provide a systematic view of how text masking can accelerate pre-training of VLMs while at the same time maintaining or improving VLM performance.
-Our work is reproducible because it uses training data sets that are openly available and small enough to allow researchers with limited computational resources to repeat our experiments, and because we release our code [^1] https://github.com/MingliangLiang3/CLIPF.git.
+Our quantified evaluation among existing alternative text masking approaches is carried out at two scales.
+A large-scale experiment on the LAION-400M [35] data set reproduces the CLIPA results and shows CLIPF outperforming syntax masking.
+A large number of smaller scale experiments on the CC3M [36] and CC12M [3] data sets, which are widely used for VLM research [15, 21, 30, 43, 18], demonstrate that consideration of frequency is important in two ways.
+First, the frequency-based approach CLIPF emerges an excellent choice for text masking, especially as the number of input tokens decreases.
+Second, the text masking approach should be selected specifically for the application scenario (i.e., data set, input size, number of training epochs).
+However, in general masking strategies that disproportionately impact word frequency (truncation and syntax) should be avoided and strategies that protect relative word frequencies (random, block and CLIPF) should be preferred.
+Our code is available online.
+[^1] https://github.com/MingliangLiang3/CLIPF
 2 Related work
 2.1 Vision-Language Models
-Vision-language models (VLMs), such as CLIP, ALIGN, LiT, and FLIP [3, 4, 10, 8], learn image embeddings from large-scale natural language supervision using contrastive learning [14], which is important for making models transferable.
+Vision-language models (VLMs) learn image embeddings from large-scale natural language supervision, which is important for making models transferable.
+CLIP [32] is the leading example of a VLM widely-used in practice, and other examples are ALIGN, SigLIP, and FLIP [32, 16, 45, 22], CLIP utilizes an image encoder and a text encoder to generate the representation of the image and text, and learn the joint representation by contrastive learning [32, 39].
 Because CLIP does not release its pre-training dataset, OpenCLIP pre-trained models on several large-scale public datasets [3, 38, 35, 34] to reproduce CLIP.
-This implementation adheres to the scaling law observed in Neural Language Models [19] and Vision-Language Models [3], where loss decreases according to a power-law relationship with model size, dataset size, and computational resources utilized during training.
-While large language models (LLMs) have advanced rapidly, we focus on vision-language models (VLMs) like CLIP, as they are essential tools for bridging visual and textual modalities in current research.
 Although CLIP and OpenCLIP are pre-trained on very large-scale datasets, recent works like DeCLIP [21] and SLIP [30] demonstrate that Vision-language Models (VLMs) can be trained efficiently with smaller datasets, enabling researchers with limited computational resources to contribute.
-Moreover, masking image patches or reducing image resolution during VLM pre-training also substantially decreases computational costs while maintaining or enhancing performance, as shown by techniques like FLIP, RECLIP, CLIPA [8, 21, 12].
+Moreover, masking image patches or reducing image resolution during VLM pre-training also substantially speed up pre-training while maintaining or enhancing performance, as shown by techniques like FLIP, RECLIP, A-CLIP and CLIPA [22, 19, 43, 20].
 2.2 Text masking for VLM training
-Masking text presents another viable option to speed up the pre-training of VLMs, and is particularly effective alongside strategies that employ a high ratio of image patch masking.
-FLIP further explored text masking techniques, namely, random and truncation masking, to enhance efficiency [8].
-While random text masking involves masking text arbitrarily, truncation masking in FLIP strategically eliminates padding tokens and retrains the words that preceded them.
+Masking text presents another viable option to speed up and enhance the pre-training of VLMs, and is particularly effective alongside strategies that employ a high ratio of image patch masking [22, 20].
+FLIP further explored text masking techniques, namely, random and truncation masking (referred to as “prioritized”), to enhance efficiency [22].
+While random text masking involves masking text arbitrarily, truncation masking in FLIP strategically eliminates padding tokens and retains the words that preceded them.
+The study of text masking for VLMs got off to a slow start, since [22] found that text masking promoted efficiency, but deteriorated performance.
 Subsequently, CLIPA [20] introduced block masking and syntax masking.
-As already mentioned in Section 1, CLIPA [12] shows that text masking during training can improve the performance of a VLM.
+As already mentioned in Sec.
+1, CLIPA [20] shows that syntax masking achieves the best performance when applying text masking.
 CLIPA syntax masking prioritizes the retention of nouns, followed by adjectives and then other words.
 In [20], syntax masking outperforms the truncation, random, and block approaches.
-However, [12] does not vary the number of training epochs.
+However, [20] experiments only with a fixed number of epochs, and does not test performance as the number of training epochs increases.
 An important contrast with our work, as previously mentioned, is our attention to the number of training epochs.
 Further, CLIPF, as we will see, counters the danger of giving overly high priority to nouns, leading to overfitting, and also avoids overly suppressing words of the POS category “Other”.
-An initial effort to exploit frequency information for text masking during VLM training was made by SW-CLIP [22].
-Unfortunately, SW-CLIPF introduces a frequency cutoff threshold, which reduces performance.
-In order to exploit frequency information effectively each training text needs to fill as many available input tokens as possible, as is done in CLIPF.
-3 Text Masking
-In this section, we describe the text masking strategies that we study in this paper: the current state of the art, namely, the text masking strategies of CLIPA [12] (Section 3.1) and our frequency masking strategy CLIPF (Section 3.2).
-All these strategies accelerate pre-training by masking specific words in the text.
-Consider a text $T_{i}$ consisting of words $\{w_{1},w_{2},w_{3},\ldots,w_{n}\}$, where $n$ represents the total number of words in $T_{i}$.
+As mentioned in the introduction, CLIPF is inspired by earlier training of text models [29], which selected words by frequency.
+An initial attempt was made to transfer this idea to CLIP training in a short workshop paper [23] published at the same time as CLIPA.
+The algorithmic novelty of our CLIPF strategy compared to the frequency masking in  [29] and [23] is the use of the word masking probability score for prioritization, based on the insight that the hard threshold must be dropped.
+In short, in order to exploit frequency information effectively, each training text needs to fill as many available input tokens as possible, as is done in CLIPF.
+3 Considering Word Frequency For Masking
+In this section, we introduce existing text masking strategies and carry out an analysis on their impact on word frequency in the training data, which reveals why random and block masking outperform syntax and truncation masking.
+3.1 Existing Text Masking Strategies
+Here, we define text masking and provide descriptions of the existing text masking strategies introduced by FLIP [22] and CLIPA [20].
+Text Masking: Consider a text $T_{i}$ consisting of $n$ words $\{w_{1},w_{2},w_{3},\ldots,w_{n}\}$.
 Text masking involves selecting a subset of these words to create a new version of the text, denoted as
-$T_{new\_i}=f(T_{i}\{w_{1},w_{2},w_{3},\ldots,w_{n}\}).$ (1)
-This function $f$ determines which words are masked to get the modified text $T_{new\_i}$.
-And, we define the number of input tokens as $k$.
-It is important to note that the number of padding tokens remains consistent across all these strategies.
-3.1 Existing Text Masking Strategies (CLIPA)
-The four text masking strategies from CLIPA [12], which we also investigate here, are:
-Random Masking: In the random text masking strategy, the function $f$ operates to randomly select words from the text $T_{i}$.
-Each word $w_{i}$ has an equal probability of being masked, independently of its position or frequency in the text.
-Truncation Masking: In truncation masking, the function $f$ reduces the text by masking words from the end part of the text.
-The number of text tokens retained depends on the input length of the text encoder.
-Block Masking: Block masking is similar to truncation masking but starts at a random index within the text.
-If the text tokens are longer than the input tokens, we choose a random starting index $j$ that is less than the length of the text tokens minus the length of the input tokens $k$.
-Syntax Masking: To implement syntax masking [12], each word $w_{i}$ in the text $T_{i}$ is first tagged using a natural language processing tool like NLTK [23] to identify its part of speech.
+$T^{\prime}=f(T\{w_{1},w_{2},w_{3},\ldots,w_{m}\}),$ (1)
+In the context of VLMs, the result of text masking is a string of tokens:
+$T^{\prime}=f(T\{t_{1},t_{2},t_{3},\ldots,t_{k}\}),$ (2)
+where $k$ is the length of the input and is determined by the masking ratio. The text is then tokenized, and $k$ tokens are chosen for use in training.
+Truncation Masking: In truncation masking, the first $k$ tokens in text $T$ are used as the input.
+Random Masking: In random text masking, the function $f$ operates to randomly select tokens from the text $T$.
+Each token $t_{i}$ has an equal probability of being masked, independently of its position or frequency in the text.
+Block Masking: Block masking is similar to truncation masking but starts at a random token index $j$ within the text $T$.
+If $T$ contains more tokens than there are input tokens, we choose a random starting index $j$ that is less than the number of the text tokens minus the number of the input tokens $k$.
+Syntax Masking: To implement syntax masking [20], each word $w_{i}$ in the text $T$ is first tagged using a natural language processing tool like NLTK [2] to identify its part of speech (POS).
 The words are then sorted according to their syntax role.
-This ordering strategy $O(w_{i})$ reflects the importance of nouns, adjectives, and verbs in the text.
-Words are masked based on their assigned order, with priority given to retaining nouns, followed by adjectives, verbs, and then other parts of speech.
-This method ensures that the core semantic information of the text is less likely to be masked, especially when learning the relationship between the image and the text, as nouns are more relevant to the object of the image.
-3.2 Word Frequency Masking (CLIPF)
-As mentioned in the introduction, subsampling of frequent words was introduced by [13] in the natural language processing (NLP) domain to accelerate the learning process for word representations.
-The goals are to speed up training and to address the imbalance between frequent and infrequent words in the training corpus.
-Our CLIPF method, uses Formula 2, which was introduced in [13], to determine the probability of masking a word during training.
-Thus, the complete formula to calculate the probability $P(w_{i})$ is:
-$P(w_{i})=1-\sqrt{\frac{t}{f(w_{i})}}$ (2)
-Here, the masking probability of the word $w_{i}$ is $P(w_{i})$, $t$ is the threshold, and $f(w_{i})$ is the frequency of the word.
-This formula scales the word frequency $f(w_{i})$, resulting in a much higher masking probability for frequent words than infrequent ones.
-It also preserves the overall ranking of word frequencies while aggressively masking high frequency words, controlled by the threshold $t$ [13].
-4 Experimental Setup
-Our implementation follows CLIP [3], OpenCLIP [9], FLIP [8] and CLIPA [12]. In this section, we present the details of our experimental setup.
-Dataset We pre-train the models on Conceptual Captions 3M (CC3M) [24], Conceptual 12M (CC12M) [15].
-These datasets were selected due to their diverse range of real-world concepts and scenes.
-Popular VLMs such as OpenCLIP, FLIP, DeCLIP, SLIP, A-CLIP and CLIPA [9, 8, 20, 5, 11, 12] also employ these datasets for training, leveraging their extensive and diverse content.
-Due to the expiration and inaccessibility of some URLs in the training set, we have downloaded 2.7 million image-text pairs for CC3M and 9.3 million image-text pairs from CC12M.
-The characteristics of the dataset are presented in Table 1.
-The CC12M dataset is 3.42 times larger than the CC3M dataset, with captions that are on average 2.15 times longer.
-Since the number of input tokens is fixed, the impact of different text masking ratios varies across datasets due to differences related to caption length.
-In an additional set of experiments, we provide results training on LAION-400M [17], a larger data set with about 300M image-text pairs.
-Table 1: Dataset Statistics for Pre-training datasets. Caption length refers to the number of words in the text.
-Dataset Samples Total words Caption length
-CC3M 2721877 28022082 10.30 $\pm$ 4.72
-CC12M 9295444 205716854 22.15 $\pm$ 17.20
-Architecture For the image encoder, we use ViT-B/16 [25] architectures with global average pooling and learnable positional embeddings.
-The input image size is 224 for CC3M and CC12M.
-For the text encoder, we use a Transformer-based model [26] and byte-pair encoding with a 49K token vocabulary.
-We chose these encoders because they are the largest encoders that we can use given our resources.
-When using full text for training, the maximum context length is 32.
-Zero-padding is applied to input text that is shorter than the maximum token length of the model.
-We trained the models using 8 RTX A5000 GPUs with the same settings to ensure consistent conditions across all models.
-We experiment with different text token input sizes, namely, 32, 16, 8, 6, and 4 text tokens.
-As the input size gets smaller, we can increase the batch size, maximizing computational memory usage.
-For CC3M and CC12M, the batch sizes corresponding to the input sizes are: 664, 832, 896, 928, and 944.
-Note that CLIP (and therefore CLIPA and CLIPF) is trained using contrastive learning, which benefits from larger batch sizes.
-Note also that although we used different text masking with the same settings and batch sizes, syntax masking was slower than the other text masking strategies when conducting the experiments because POS processing for each word is time-consuming.
-We pre-train the model using the InfoNCE loss [14] with a learnable temperature parameter $\tau$ [27, 3]. To classify images, we calculate the cosine similarity between the image and text embeddings.
-Table 2: Details of the pre-training and fine-tuning setups.
-Config Pre-training Fine-tuning
-optimizer AdamW [28] AdamW [28]
-learning rate 1e-3 1e-5
-weight decay 0.2 0.2
-optimizer momentum $\beta_{1},\beta_{2}=0.9,0.98$ [29] $\beta_{1},\beta_{2}=0.9,0.98$ [29]
-learning rate schedule cosine decay [30] cosine decay [30]
-warmup steps 10k 10%
-epoch 30 1
-$t$ $10^{-6}$
-$\tau$ 0.07
-numerical precision amp amp
-RandomResizedCrop (50, 100) (50, 100)
-Table 3: The number of image and text tokens that are processed during pre-training (CC3M and CC12M).
-Image masking Text masking Image tokens Text tokens Total Percentage
-0% 0.00% 196 32 228 100%
-75% 0.00% 49 32 81 35.53%
-50.00% 49 16 65 28.51%
-75.00% 49 8 57 25.00%
-81.25% 49 6 55 24.12%
-87.50% 49 4 53 23.25%
-Training Following CLIP, OpenCLIP [3, 9], we pre-train our model for 30 epochs on the CC3M and CC12M datasets. Details of the pre-training configuration are given in Table 2.
-Similar to FLIP, to speed up training, we apply 75% image masking to the image encoder as the baseline model.
-As a result, the speedup is about 4$\times$ compared to training without image masking, while the reduction in the zero-shot performance of ImageNet-1K classification remains within reasonable bounds.
-During text masking, we reduce the number of tokens from 32 to 16, 8, 6, and 4.
-To measure the training speed of CLIP, CLIPA, and CLIPF, we compare the number of text tokens processed by each model.
-As shown in Table 3, the number of text tokens processed by each model during pre-training is calculated based on different image and text masking ratios.
-When we pre-trained the model using image masking, the total number of tokens is 81 and the percentage of text tokens is 39.5%.
-When we apply 50% text masking, the total number of tokens is 65.
-Compared to training without text masking, this results in a speed increase of approximately 20%.
-Moreover, when we apply 87.5% text masking, the total number of tokens is 53, resulting in a speed increase of approximately 34% compared to training without text masking.
-Fine-tuning Following FLIP, and CLIPA, in order to bridge the distribution gap between pre-training and evaluation, we fine-tuned the model without images and text masking.
-Note that, fine-tuning the models without masking in masking work is focused on reducing the distribution gap and applied sparingly to avoid undoing the tradeoff.
-The details of the fine-tuning configuration are provided in Table 2.
-Table 4: Comparison of CLIPF, CLIP, and CLIPA for zero-shot classification on ImageNet-1K.
-Reported as top-1 accuracy.
-All models use a ViT-B/16 image encoder backbone pre-trained on CC3M and CC12M for 30 epochs with 75% image masking for speedup (pre-train) and fine-tuned for an additional epoch on all data (fine-tune).
-“Image tokens” and “text tokens” reflect the amount of image and text processing during pre-training.
-Models Masking Image Tokens Text Tokens CC3M CC12M
-pre-train fine-tune pre-train fine-tune
-CLIP 197 32 18.6 36.6
-FLIP 49 32 14.1 14.2 32.0 33.7
-CLIPA truncation 49 16 13.8 13.8 32.8 32.8
-random 13.9 13.9 33.7 34.3
-block 13.9 13.9 34.3 34.8
-syntax 13.3 12.8 32.2 34.4
-CLIPF frequency 14.0 14.0 35.5 36.0
-CLIPA truncation 49 8 10.8 12.0 25.4 28.4
-random 17.6 17.4 34.5 36.9
-block 16.2 16.6 35.5 37.9
-syntax 17.2 17.5 28.5 35.0
-CLIPF frequency 16.8 17.0 36.6 39.3
-CLIPA truncation 49 6 8.4 9.4 15.3 23.2
-random 12.8 17.9 26.9 34.6
-block 12.9 17.0 28.6 35.9
-syntax 12.2 15.7 25.2 32.6
-CLIPF frequency 14.4 18.2 30.3 37.8
-CLIPA truncation 49 4 3.8 8.2 5.3 19.8
-random 5.4 14.6 14.0 27.1
-block 7.5 14.5 18.7 26.6
-syntax 8.9 13.0 14.2 24.6
-CLIPF frequency 10.9 16.0 17.0 30.9
-CLIPF frequency 98 8 19.6 18.7 39.8 41.0
-Evaluation setup Following CLIP [3], FLIP [8], and CLIPA [12], several classification benchmarks were used.
-Among these benchmarks is ImageNet-1K, a widely recognized dataset in computer vision.
-It is frequently used for image classification and VLM tasks and comprises 50K validation samples across 1K different classes.
-We filled each class into the templates provided by CLIP [3] to calculate the average of the text embeddings.
-We use the same evaluation settings as CLIP [3] to evaluate the other downstream datasets.
-5 The Importance of Frequency Distributions
-Our work is motivated by the patterns that we observe when VLMs are trained for additional epochs, shown in Fig.
-1.
-This experiment motivates revisiting CLIPA [12], which was trained only on 6.4 epochs, in two respects.
-First, as already mentioned in the introduction, the conclusion that syntax masking is the best masking approach does not hold.
-Second, truncation masking, which looks relatively promising in low numbers of epochs, does not perform well given more epochs.
-In this section, we analyze the impact of masking on the training data in order to discover the reason behind this behavior.
+Priority is given to nouns, followed by adjectives, verbs and then words belong to other POS categories.
+The first $k$ tokens are used as the input.
+An example that illustrates the way that the different forms of masking impact a text in the training set is provided in Fig. 2.
 Figure 2: Comparison of word masking probabilities for various methods.
@@
 Truncation keeps the first 4 words; Random and Block mask each word with a 50% probability; Syntax prioritizes retaining nouns, followed by adjectives, then others; Frequency masks words based on the frequency of the words.
-We start with a basic illustration of the differences between task-masking strategies is provided in Figure 2.
-Truncation consistently retains the beginning of the text.
-With the random (random word selection) and block (random block selection) methods, each word has an equal probability of being masked, with block masking maintaining word neighborhoods.
-Syntax masking prioritizes the retention of nouns, followed by other part-of-speech (POS) categories.
-Figure 3: The figure illustrates the distribution of top-50 words in the text before and after applying various text masking strategies.
+3.2 Impact of masking on word frequency
+Figure 3: The figure illustrates the distribution of top-25 words in the text before and after applying various text masking strategies.
 We set the text length after text masking to 6.
 The x-axis represents the word index, which is sorted by counts of the original data, and the y-axis shows the word frequency.
-The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$.
+The dataset used is CC12M and the value of $t$ of Eq.
+3 is set to $10^{-6}$.
 We remove special characters from the vocabulary.
-Next, we look at how the different masking strategies impact the frequency distribution of words in the training data.
-Fig.
-3 shows the distribution of words in the text before and after applying all of the text masking strategies.
-We see that in general, there is quite a large difference between the frequency of the highest-frequency words and the frequency of words with lower frequency.
-Such a long-tail distribution is characteristic of words occurring in text, which is known to be Zipf distributed.
-Most masking approaches preserve the general shape of this distribution: The frequency of a word in the masked training data tracks its frequency in the original data.
-The exception is syntax masking, which shows peaks and pits for specific words.
-Specifically, because syntax masking retains more nouns, the articles “the” and “a” are less frequent, but the nouns create spikes.
-The rationale for prioritizing nouns is solid.
-Nouns are closely related to the depicted content of images and, for this reason, are considered to be more helpful.
-Indeed, the helpfulness of nouns for the VLM is reflected in the fact that syntax masking learns faster than other text masking strategies.
-In Fig.
-1, we saw that syntax masking outperforms the other masking approaches during the early epochs.
-However, retaining too many nouns can have two adverse effects.
-First, the model can overfit to certain high frequency nouns.
-In Fig.
-3, we see that CLIPA syntax masking retains more occurrences of “person,” “illustration,” and “stock” than the other strategies.
-Second, retaining nouns means necessarily dropping words from other parts-of-speech (POS) categories, which may be important, as discussed in more detail below.
-Considering the CLIPA truncation masking strategy, we see in Fig.
-1 that the word frequency distribution of the truncated training data shows a difference with the other masking strategies.
-Specifically, truncation is the top line on the left of the graph, which means truncation is making the most highly frequent words in the training data relatively even more frequent.
-The fact that truncation impacts word frequency might seem surprising at first, but it can be explained if we consider that the probability of a word occurring changes with sentence position.
-With truncation, typical text beginnings, e.g., “A person is…”.
-remain over-represented in the data.
-To dive deeper into the importance of the distribution of words over POS categories, we provide statistics for each of the different text masking strategies in Table 5.
-Consulting the column “Other”, we see an interesting pattern.
-Syntax masking changes the proportion of words in the “Other” category radically with respect to the data before masking, but truncation, random and block do not.
-CLIPF appears to provide a happy medium.
-CLIPF retains more nouns than CLIPA truncation, random, and block, but not so many that the “Other” category becomes poorly represented.
-Table 5: To examine the distribution of words belonging to different POS categories, we calculate the percentage of words in the training data belonging to each category in before and after applying text masking.
-The dataset is CC12M, and we retain 6 words for each text after applying text masking.
+The figure with more words is provided in the supplementary material.
+Existing literature reports the impact of text masking in terms of the percentage of words or tokens retained in the training data.
+Seen from the word-count perspective, the impact is the same: for any text masking approach, the proportion of words removed controls the computational savings.
+However, to determine the reasons why some text masking approaches are better than others, it is necessary to look beyond overall word count.
+In this section, we carry out an analysis that goes beyond word count and considers the impact of text masking on the frequency of words occurring in the training set.
+We plot word-frequency curves, used in linguistics, for the four existing approaches to text masking and display the head of the curve (the first 25 most frequent words) in Fig.
+6.
+The curves generally follow the long-tail distribution expected of natural language and predicted by Zipf’s Law [17].
+Further, linguistic theory leads us to expect that the top-frequency words are function words (e.g., articles and prepositions), which is indeed what we observe here.
+However, looking at individual words in Fig.
+6, we see aberrations.
+Two observations are important: First, for truncation masking the four most frequent words are markedly more frequent than with the other forms of masking.
+These are function words (“the”, “a”, and “of”) and one noun (“person”), which reflects the fact that sentences tend to start with these words.
+Second, for syntax masking certain nouns form peaks, meaning that they are more frequently retained with syntax masking than with other forms of masking.
+Given the design of syntax masking, it is unsurprising that the words that are disproportionately retained belong to the POS category Noun.
+We dive deeper into the impact of existing masking approaches by examining the word distribution over POS categories.
+Tab.
+1 illustrates statistics for four major part-of-speech classes, Nouns, Adjectives, Verbs and Other (predominantly function words).
 Masking NN (%) JJ (%) VB (%) OTHER (%)
@@
 CLIPF 60.43% 5.06% 6.20% 28.24%
-The importance of “Other” words is surprising because the category includes many words, which like articles, lack semantic content and are sometimes considered to have no direct connection to images.
-In fact, in many NLP applications, articles and other “stop words” are simply removed.
-However, since the CLIP text encoder is a transformer, technically it should be able to exploit the context of content words, particularly words in “Other” category.
-Note that with syntax masking a phrase like “smoking in the room” in the training data would always be reduced to “smoking room”, which has a different meaning.
-However, with the other masking approaches, “in the” would not be consistently deleted.
-Note specifically that block masking retains entire phrases, without always using the beginning of the sentence as truncation masking does.
-Also, frequency masking (CLIPF) retains entire phrases in some cases, more than would be expected with random masking.
-In sum, when trained with enough epochs, overfitting effects with syntax masking occur.
-Our analysis has related this problem to peaks in the word frequency distribution caused by over-retention of nouns.
-Also, with enough epochs, truncation masking overfitting effects occur, which our analysis has related to distortion of the word frequency distribution caused by oversampling of the beginning of sentences.
-In sum, our analysis of word frequency distributions in the training data provides insight into why the performance of syntax and block masking falls off in Fig.
+Table 1: Part-of-speech (POS) distribution: The percentage of nouns (NN), adjectives ($J$), verbs (VB) and Other POS categories before and after text masking, calculated on CC12M, retaining 6 words per text.
+We observe that the distribution for truncation, random, and block masking is comparable to the distribution before masking.
+However, for syntax masking, the number of words in the Noun category is extremely high and the number of words in the Other category is extremely low.
+The prioritization of the POS category Noun and the deprioritization of Other has intuitive appeal.
+Although not explicitly mentioned, we assume that this appeal motivated CLIPA [20] in their design of syntax masking.
+Nouns are likely to be the name of an object directly depicted in an image, and can form a better bridge between image and text than words in the categories Adjectives or Verb.
+Further, words in the category of Other, being function words, are often classical stop words that are removed in information retrieval applications.
+Our results (Fig.
+1) show that this intuition does not directly lead to good performance and must be revisited.
+In short, our analysis reveals the importance of considering word frequency when text masking.
+We have seen that text masking approaches that do not disproportionately impact the frequency of certain words (i.e., random and block) outperform text masking strategies that do (i.e., truncation and syntax).
+The word frequency changes caused by using sentence-initial words (truncation) or emphasizing/de-emphasizing certain POS categories (syntax) hold back performance as the number of epochs increases.
+This observation motivates the importance of a text masking approach that is directly maintains relative word frequency, which we introduce in the next section.
+4 Word Frequency Masking (CLIPF)
+Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF) removes a word $w_{i}$ from the training data according to its masking probability score, $P(w_{i})$, which is calculated on the basis of the word’s frequency as:
+$P(w_{i})=1-\sqrt{\frac{t}{f(w_{i})}}$ (3)
+Word frequency, $f(w_{i})$, is the number of times the word appears in the training data divided by the total number of words in the training data.
+Eq.
+3 scales $f(w_{i})$, resulting in a much higher masking probability for frequent words than infrequent ones.
+It also preserves the overall ranking of word frequencies.
+The aggressiveness with which high frequency words are masked is controlled in Eq.
+3 by the threshold $t$.
+CLIPF enjoys the benefit of being relatively insensitive to the parameter $t$, meaning that the parameter does not have to be precisely tuned.
+The influence of $P(w_{i})$ on $t$ is illustrated by a plot in the supplemental material, which also includes an experimental demonstration that the CLIPF is relatively insensitive to the choice of $t$.
+The basic form of Eq.
+3 has been validated previously in the literature, having been introduced by [29] for the purpose of text model training, which inspired CLIPF.
+Specifically, [29] and SW-CLIP [23] (which attempted to transfer [29] to CLIP training) discard words whose $P(w_{i})$ exceeds a cutoff.
+If there are not enough low $P(w_{i})$ words in a training text to fill in the input, these approaches use padding.
+CLIPF does away with this cutoff, and instead fills the input with words in order of $P(w_{i})$.
+Because CLIPF uses higher-$P(w_{i})$ words rather than padding it can outperform SW-CLIP, as shown in the supplementary material.
+5 Performance of CLIPF
+In this section, we discuss the performance of CLIPF on the basis of a reproduction of the CLIPA paper ( Sec.
+5.1) with a model trained on LAION-400M [35] as well as the initial experiment in Fig.
+1 with models trained on CC3M [36] and CC12M [3].
+5.1 Revisiting CLIPA
+We carry out a reproduction of the original CLIPA experiment on LAION-400M [35].
+We remain as faithful as possible to the original setup, with two exceptions.
+First, not all 400M URLs from LAION-400M are still available online, and we train on the 298M pairs that are available.
+Second, we reduce the computational expense of the experiment by using a smaller image size (112x112) and 50% image masking.
+We employ a ViT-B/16 image encoder with global average pooling and learnable positional embeddings.
+As with CLIPA, we fine-tune for 0.4 epochs on the full data after training on the masked data.
+More details about the experimental setup are available in the supplementary material.
+Results for zero-shot classification on ImageNet-1K are shown in Fig.
+4 for both top-1 accuracy (bottom) and top-5 accuracy (top) for 4-16 epochs of training on the masked data.
+We observe that CLIPF outperforms syntax masking.
+Figure 4: Zero-shot classification accuracy on ImageNet-1K using models trained on LAION-400M to which different text masking strategies have been applied.
+The backbone of the image encoder is ViT-B/16.
+We pre-trained the model using 25 image tokens and 4 text tokens for 16 epochs.
+At each point, we apply an additional 0.4 epoch of training on the full, unmasked data.
+The original CLIPA paper established that truncation masking was not a good performer, and we do not test it here.
+However, we find that other aspects of the conclusions of CLIPA concerning text masking are limited.
+Specifically, in order to observe this difference it is necessary to continue training until the increase in performance starts to level off.
+The original CLIPA paper was trained with only 6.4 epochs, and here we train to 16.
+The additional epochs allow us to observe block masking and random masking moving towards syntax masking as the number of epochs increases, with block masking on par with syntax masking with respect to top-5 accuracy at 16 epochs.
+Comparing Fig.
+4 with the first part of Fig.
+1, we see that the relative levels of performance of syntax, block, random and CLIPF are the same.
+This comparability gives us confidence that we can study text masking with smaller data sets, which allows us to train for a large number of epochs using different comparative conditions, and we carry out the remaining experiments on data sets that require less computational expenditure.
+5.2 Advantages of CLIPF
+We report on the performance of CLIPF vs.
+the original CLIP with an experiment carried out on CC3M [36] and CC12M [3].
+Not all text-image pairs were available online and our final data sets consist of 2.7M pairs from CC3M and 9.3M pairs from CC12M.
+For CLIPF, we fine-tune for one epoch on the full data after training on the masked data.
+We employ the same ViT-B/16 image encoder, but we use a 224×224 image input size.
+Further details about the experimental setup are available in the supplementary material.
+The results in Tab.
+2 confirm that CLIPF is able to outperform CLIP.
+“Image tokens” and “text tokens” reflect the amount of image and text processing during pre-training and this performance gain comes with the computational advantage of using half the number of image tokens and one quarter of the number of text tokens.
+Models Masking Image Tokens Text Tokens CC3M CC12M
+pre-train fine-tune pre-train fine-tune
+CLIP 197 32 18.6 36.6
+CLIPF frequency 98 8 19.6 18.7 39.8 41.0
+Table 2: Comparison of zero-shot classification on ImageNet-1K between CLIP and CLIPF using 98 image tokens and 8 text tokens.
+Reported as top-1 accuracy.
+All models use a ViT-B/16 image encoder backbone.
+They are pre-trained on CC3M and CC12M for 30 epochs (pre-train).
+CLIPF is fine-tuned for an additional epoch on all (no masking) data (fine-tune).
+When text masking in an application scenario, choices need to be made, as already mentioned in Sec.
 1.
-Further, if word frequency is indeed important, then masking that controls word-frequency directly, such as our CLIPF, can be considered to be a sensible and transparent approach.
-6 Experimental Results
-In this section, we carry out a systematic investigation of the existing text masking approaches from CLIPA [12] and our CLIPF approach.
-Our results support our observation that syntax and truncation masking are not productive approaches to masking and highlight the benefits of masking that directly leverages frequency information, i.e., our CLIPF.
-Our experiments use 30 epochs, following Open_CLIP [9].
-We note that VLM training work has shown performance improvements up to 100 epochs [5, 11] and we would expect that with more epochs the performance would also improve.
-However, the improvement beyond 30 is expected to be small, predictable and not necessarily important in an operational setting where conserving compute is of concern.
+As seen in Fig.
+1, the number of epochs matters.
+Here, we choose to train with 30 epochs, following [15].
+The aggressiveness of the masking used will be determined by the computational resources available in the application scenario.
+Here, we used 50% image masking and in our further experiments, we adopt 25% image masking for greater speed.
+In the next section, we carry out a quantified evaluation of all existing text masking approaches, comparing them to CLIPF.
+This evaluation provides the information necessary to support the choices of text masking approaches, including the number of text input tokens, for use in application scenarios.
+6 Comparison of Text Masking Strategies
+In this section, we carry out a systematic evaluation of the existing text masking approaches from FLIP [22] (random and truncation), CLIPA [20] (block and syntax) and our CLIPF approach (frequency-based).
+In a nutshell, the results support our observation that syntax and truncation masking are not productive approaches to masking and highlight the benefits of masking that directly leverages frequency information, i.e., our CLIPF.
+These experiments are carried out on CC3M [36] and CC12M [3] with the same experimental setup as in Sec.
+5.2, using 30 epochs, as previously mentioned, following OpenCLIP [15].
+Following prior evaluations of CLIP models [32, 22], we evaluate the models’ zero-shot performance on ImageNet-1k (Sec.
+6.1), zero-shot robustness (Sec.
+6.2), zero-shot image-text retrieval (Sec.
+6.3) and zero-shot classification on 25 additional data sets (Sec.
+6.4).
+Further details of the experimental setup are in the supplementary material.
 6.1 Zero-shot Performance on ImageNet-1k
-Table 4 presents results for zero-shot classification on ImageNet-1K for models trained on both the CC3M and CC12M data.
+Tab.
+3 presents results for zero-shot classification on ImageNet-1K, which complement the ImageNet-1K results already presented in Fig.
+1.
 The experiments compare different masking approaches and different masking ratios, indicated in the table with the number of input text tokens (Text Tokens column).
@@
 Note that since the average text length in CC3M is less than 16 tokens, the benefits of fine-tuning a model pre-trained with 16 text tokens are limited.
-We see that CLIPF outperforms the other CLIPA matching approaches in nearly all cases across text masking ratios (input sizes) as reflected by the bolded accuracy scores.
-In general, random, syntax, and block strategies exhibit similar performance; however, truncation performs worse.
-Contrary to the conclusion of the original CLIPA work [12], syntax masking does not yield the best performance.
-It is particularly interesting to observe from the results in Table 4 the fact that models trained with text masking (CLIPA and CLIPF) are competitive with, or superior to, models trained on the original, unmasked text data (CLIP, FLIP).
+From Tab.
+3, we observe the ability of fine-tuned text-masking strategies to outperform FLIP [22] in many cases (image masking alone), while using fewer text input tokens.
+This result is important since [22] states that text masking does not contribute to improvement in model performance.
+We attribute the fact that we find a gain to the limited number of epochs (6.4) with which [22] tested text masking.
+From Tab.
+3, we also observe that CLIPF outperforms the other masking matching approaches in nearly all cases across text masking ratios (input sizes), as reflected by the bolded accuracy scores.
+Fine-tuned CLIPF delivers good performance both on CC3M and also the larger CC12M.
+On CC12M, fine-tuned CLIPF even outperforms the original CLIP (no masking) with a 75% masking ratio (8 input tokens).
 Note that at a text masking ratio of 75% (corresponding to 8 tokens), the masking approaches are 1.3 $\times$ faster than FLIP and 4.0 $\times$ faster than CLIP.
-We first consider the comparison with CLIP.
-On CC3M, CLIPF approaches the performance of the original CLIP approach with a 81% text masking ratio (4 input tokens).
-On CC12M, CLIPF outperforms the original CLIP (no masking) with a 75% masking ratio (8 input tokens).
-Fine-tuned CLIPF with a 75% masking ratio achieves an accuracy of 39.3% compared to 36.6% for the original CLIP trained on the full data.
-Next, we consider the comparison with FLIP, which is the same as CLIP except that it uses a 75% image masking ratio (49 image tokens).
-On CC3M, fine-tuned CLIPF is competitive with or outperforms fine-tuned FLIP at all masking ratios.
-On CC12M, fine-tuned CLIPF outperforms fine-tuned FLIP, except for the highest masking ratio (87.5% corresponding to 4 input tokens).
-Additionally, as shown in the last row of Table 4, if we double the number of image tokens, then CLIPF substantially beats the original CLIP (41.0% vs 36.6%).
-However, it also notes that more resources are required (Tokens processed from 25% to 46%).
+Truncation performs poorly, consistent with the observations of CLIPA [20].
+However, in contrast to [20], we observe that syntax masking is a poor performer.
+In contrast, approaches that do not disproportionately impact the frequencies of words (random and block) perform more strongly.
 Finally, we consider the pattern of the text masking approaches as the masking ratio increases, i.e., as we decrease the input size from 16 to 4 tokens.
@@
 Moving from 16 to 8 input tokens benefits most approaches.
-Recall that the reduction of the number of input tokens allows for an increase in the batch size.
-As we reduce the number of tokens from 8 to 6 to 4, syntax masking and truncation do not perform as well as the other text masking strategies.
+The benefit derives from the fact that the reduction of the number of input tokens allows for an increase in the batch size.
+As we reduce the number of tokens from 8 to 6 to 4, the gap between fine-tuned CLIPF and the next best masking strategy widens.
 We see that compared to existing text matching approaches, CLIPF is better positioned to deliver benefits as the number of input tokens decreases.
+Models Masking Image Tokens Text Tokens CC3M CC12M
+pre-train fine-tune pre-train fine-tune
+CLIP 197 32 18.6 36.6
+FLIP 49 32 14.1 14.2 32.0 33.7
+FLIP truncation 49 16 13.8 13.8 32.8 32.8
+random 13.9 13.9 33.7 34.3
+CLIPA block 13.9 13.9 34.3 34.8
+syntax 13.3 12.8 32.2 34.4
+CLIPF frequency 14.0 14.0 35.5 36.0
+FLIP truncation 49 8 10.8 12.0 25.4 28.4
+random 17.6 17.4 34.5 36.9
+CLIPA block 16.2 16.6 35.5 37.9
+syntax 17.2 17.5 28.5 35.0
+CLIPF frequency 16.8 17.0 36.6 39.3
+FLIP truncation 49 6 8.4 9.4 15.3 23.2
+random 12.8 17.9 26.9 34.6
+CLIPA block 12.9 17.0 28.6 35.9
+syntax 12.2 15.7 25.2 32.6
+CLIPF frequency 14.4 18.2 30.3 37.8
+FLIP truncation 49 4 3.8 8.2 5.3 19.8
+random 5.4 14.6 14.0 27.1
+CLIPA block 7.5 14.5 18.7 26.6
+syntax 8.9 13.0 14.2 24.6
+CLIPF frequency 10.9 16.0 17.0 30.9
+Table 3: Comparison of CLIP, FLIP, CLIPA and CLIPF for zero-shot classification on ImageNet-1K.
+Reported as top-1 accuracy.
+All models use a ViT-B/16 image encoder backbone pre-trained on CC3M and CC12M for 30 epochs with 75% image masking for speedup (pre-train) and fine-tuned for an additional epoch on all data (fine-tune).
+“Image tokens” and “text tokens” reflect the amount of image and text processing during pre-training.
 6.2 Zero-shot Robustness Evaluation
-Table 6: Zero-shot robustness evaluation.
-Comparison of the zero-shot accuracy performance of CLIP, FLIP, CLIPA, and CLIPF on various datasets.
-The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
 Models Masking Image Tokens Text Tokens ViT-B/16
@@
 FLIP 49 32 6.90 40.30 36.90 17.60 28.70 18.10
-CLIPA truncation 49 16 7.10 37.70 36.40 17.50 27.40 17.20
+FLIP truncation 49 16 7.10 37.70 36.40 17.50 27.40 17.20
 random 8.05 38.60 40.00 20.34 29.61 20.84
-block 8.05 39.60 39.17 19.31 29.70 19.72
+CLIPA block 8.05 39.60 39.17 19.31 29.70 19.72
 syntax 8.00 40.70 39.41 18.61 29.17 19.96
 CLIPF frequency 8.28 39.10 40.86 20.48 30.78 21.15
-CLIPA truncation 49 8 6.71 34.05 37.18 15.99 24.49 15.98
+FLIP truncation 49 8 6.71 34.05 37.18 15.99 24.49 15.98
 random 9.16 38.15 43.35 22.53 31.24 22.17
-block 9.05 39.95 43.32 22.58 32.42 21.58
+CLIPA block 9.05 39.95 43.32 22.58 32.42 21.58
 syntax 8.50 38.40 39.90 19.50 29.80 19.50
 CLIPF frequency 9.56 39.90 45.76 23.62 34.16 21.59
-CLIPA truncation 49 6 6.70 26.60 33.20 13.30 20.20 14.80
+FLIP truncation 49 6 6.70 26.60 33.20 13.30 20.20 14.80
 random 8.10 36.10 41.80 21.10 30.20 19.80
-block 8.32 38.20 42.99 21.73 30.82 19.76
+CLIPA block 8.32 38.20 42.99 21.73 30.82 19.76
 syntax 7.40 34.10 39.60 18.50 28.60 19.80
 CLIPF frequency 8.69 38.55 45.09 23.45 32.59 20.04
-CLIPA truncation 49 4 5.30 21.70 25.40 9.00 17.10 11.20
+FLIP truncation 49 4 5.30 21.70 25.40 9.00 17.10 11.20
 random 6.10 30.30 33.80 14.90 23.50 15.30
-block 6.08 30.55 34.69 15.54 22.99 15.10
+CLIPA block 6.08 30.55 34.69 15.54 22.99 15.10
 syntax 6.50 29.10 32.40 13.30 22.00 16.00
 CLIPF frequency 7.05 32.40 39.90 18.94 26.72 17.14
-Following the evaluation of CLIP and FLIP [3, 8], we test the masking strategies across data sets.
-Table 6 presents results for zero-shot classification for six commonly used data sets.
-The performance of these strategies on multiple datasets is similar to the behavior observed on the ImageNet-1K dataset.
+Table 4: Zero-shot robustness evaluation.
+Comparison of the zero-shot accuracy performance of CLIP, FLIP, CLIPA, and CLIPF on various datasets.
+The models are pre-trained on CC12M [3] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
+Tab.
+4 presents results for zero-shot classification for six data sets commonly used for robustness evaluation (e.g., by [32, 22]): ImageNet-A (IN-A) [13], ImageNet-O (IN-O) [14], ImageNet-R (IN-R) [12], ImageNet-Sketch (IN-S) [41], ImageNet-V2 (IN-Vs) [33], and ObjectNet (ON) [1].
+The results lead to the same observations as with the zero-shot performance on ImageNet-1k in Sec.
+6.1.
 Across different text masking ratios and across different data sets, CLIPF outperforms the other masking strategies in nearly all cases.
-Reducing the text tokens in CLIPF from 16 to 8 results in substantial performance improvements compared to the other text masking approaches across almost all datasets: ImageNet-A [31] by 1.28%, ImageNet-O [32] by 0.80%, ImageNet-R[33] by 4.9% , ImageNet-Sketch [34] by 3.14%, ImageNet-V2 [35] by 3.38%, and ObjectNet [36] by 0.44%.
+Truncation and syntax are generally not good performers.
+Reducing the text tokens in CLIPF from 16 to 8 results in substantial performance improvements compared to the other text masking approaches across almost all datasets.
 Again, we see that the benefits of CLIPF are particularly evident as the masking number of input tokens decreases.
-We also observe that CLIPF is able to improve over the FLIP baseline without masking in a majority of cases.
+We also observe that CLIPF is able to improve over the FLIP baseline, without text masking in a majority of cases.
 However, despite being the strongest text-masking strategy, the results reveal that it is challenging for CLIPF to improve over the original CLIP.
-In only a handful of cases does it outperform the original CLIP, but might do so with more image tokens as we saw in 6.1.
-Table 7: Zero-shot Image-Text Retrieval.
-We evaluated CLIP, FLIP, CLIPA, and CLIPF image-text retrieval performance on the COCO and Flickr30k datasets.
-The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
-Models Masking Text Tokens Text Retrieval Image Retrieval
+In only a handful of cases does it outperform the original CLIP, but might do so with more image tokens as we saw in Sec.
+5.2.
+Models Masking Image Tokens Text Tokens Text Retrieval Image Retrieval
 Flickr30k COCO Flickr30k COCO
 R@1 R@1 R@1 R@1
-CLIP 32 62.62 35.54 45.42 24.22
-FLIP 32 54.73 29.34 38.62 20.88
-CLIPA truncation 16 44.67 25.54 34.99 19.64
-random 16 58.48 32.36 43.61 22.81
-block 16 56.51 30.82 44.06 22.66
-syntax 16 54.54 29.60 41.16 21.40
-CLIPF frequency 16 57.89 31.52 42.72 22.57
-CLIPA truncation 8 30.47 16.92 23.96 12.78
-random 8 58.58 36.24 43.79 23.16
-block 8 60.06 35.88 45.98 24.65
-syntax 8 50.30 29.64 38.46 20.28
-CLIPF frequency 8 58.68 34.74 44.58 23.31
-CLIPA truncation 6 32.25 16.18 22.39 11.93
-random 6 55.23 32.58 42.47 21.34
-block 6 54.44 33.12 41.68 21.73
-syntax 6 46.15 26.78 34.08 17.99
-CLIPF frequency 6 56.02 32.32 41.05 21.28
-CLIPA truncation 4 30.97 17.44 21.09 11.07
-random 4 41.62 24.14 30.26 15.63
-block 4 40.83 23.56 30.77 14.98
-syntax 4 38.66 21.68 26.08 14.14
-CLIPF frequency 4 45.07 25.60 31.72 15.96
-6.3 Zero-shot Image-Text Retrieval
-Table 7 shows image/text retrieval results on Flickr30k [37] and COCO [38].
-In general, we see that text masking approaches are able to improve over FLIP.
-When using 16 text tokens, all strategies except the truncation method outperform FLIP, which is pre-trained on the full text.
-Here, we do not see consistent improvement over CLIP, although some cases come lose.
-Note again that we are testing here only with 49 image tokens and 98 would fare better, with additional computational cost.
-In contrast to the previous experiments, with image-text retrieval, we observe that CLIPF does not clearly dominate the other text masking strategies.
-The CLIPA random, CLIPA block, and CLIPF (frequency) strategies exhibit comparable performance in image and text retrieval across both datasets.
-The notable exception is the performance of CLIPF in the case that the text masking ratio is very high, i.e., only 4 input text tokens are used.
-Here, the benefit of CLIPF’s frequency-based approach is clearly evident.
-We point out that for Image-Text Retrieval the performance of syntax masking is particularly low in comparison to the other forms of masking.
-Specifically, when using 8 and 6 text tokens, the performance gap between syntax masking and block, random, and frequency masking is approximately 9% on COCO and 6% on Flickr30k for image-text retrieval.
-This gap is wider than what was observed with Zero-shot Classification.
-We conjecture that the bad performance is due to the Image-Text Retrieval task requiring understanding the overall meaning and logic of the text rather than just recognizing nouns or objects.
-For this reason, the fact that syntax masking retains a high number of nouns, as shown in Section 5, is particularly harmful.
-Table 8: Zero-shot accuracy on more classification datasets.
-Comparison of the zero-shot accuracy performance of CLIP, FLIP, CLIPA, and CLIPF on various datasets.
+CLIP 197 32 62.62 35.54 45.42 24.22
+FLIP 49 32 54.73 29.34 38.62 20.88
+FLIP truncation 49 16 44.67 25.54 34.99 19.64
+random 58.48 32.36 43.61 22.81
+CLIPA block 56.51 30.82 44.06 22.66
+syntax 54.54 29.60 41.16 21.40
+CLIPF frequency 57.89 31.52 42.72 22.57
+FLIP truncation 49 8 30.47 16.92 23.96 12.78
+random 58.58 36.24 43.79 23.16
+CLIPA block 60.06 35.88 45.98 24.65
+syntax 50.30 29.64 38.46 20.28
+CLIPF frequency 58.68 34.74 44.58 23.31
+FLIP truncation 49 6 32.25 16.18 22.39 11.93
+random 55.23 32.58 42.47 21.34
+CLIPA block 54.44 33.12 41.68 21.73
+syntax 46.15 26.78 34.08 17.99
+CLIPF frequency 56.02 32.32 41.05 21.28
+FLIP truncation 49 4 30.97 17.44 21.09 11.07
+random 41.62 24.14 30.26 15.63
+CLIPA block 40.83 23.56 30.77 14.98
+syntax 38.66 21.68 26.08 14.14
+CLIPF frequency 45.07 25.60 31.72 15.96
+Table 5: Zero-shot Image-Text retrieval evaluation.
+We evaluated CLIP, FLIP, CLIPA, and CLIPF image-text retrieval performance on the COCO and Flickr30k datasets.
 The models are pre-trained on CC12M [3] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
@@
 FLIP 32 36.26 58.09 26.89 8.73 38.56 8.16 2.98 19.95 51.85 67.17 21.57 2.94 9.35 81.89 18.02 32.30 9.68 21.39 3.63 50.26 36.51 14.67 51.37 50.08 33.70
-CLIPA truncation 16 35.53 61.24 26.54 9.42 39.98 7.28 2.57 15.80 46.85 68.28 20.47 1.94 8.32 81.89 18.48 33.75 2.33 17.98 3.79 58.56 30.85 13.31 53.00 50.08 32.83
+FLIP truncation 16 35.53 61.24 26.54 9.42 39.98 7.28 2.57 15.80 46.85 68.28 20.47 1.94 8.32 81.89 18.48 33.75 2.33 17.98 3.79 58.56 30.85 13.31 53.00 50.08 32.83
 random 16 38.48 53.61 29.46 7.97 41.92 8.53 2.79 21.54 56.71 69.67 22.88 2.23 9.73 84.86 17.90 36.41 7.28 32.95 4.77 55.00 35.29 16.38 49.93 50.08 34.26
-block 16 38.69 57.51 29.71 8.23 44.54 9.85 2.45 21.33 54.01 70.61 23.00 2.79 10.09 85.04 31.88 29.98 9.07 25.07 4.59 60.68 38.07 18.72 51.24 49.64 34.78
+CLIPA block 16 38.69 57.51 29.71 8.23 44.54 9.85 2.45 21.33 54.01 70.61 23.00 2.79 10.09 85.04 31.88 29.98 9.07 25.07 4.59 60.68 38.07 18.72 51.24 49.64 34.78
 syntax 16 35.94 58.78 31.86 8.91 41.92 8.39 1.53 17.07 47.05 67.59 21.34 3.10 10.45 76.92 26.68 32.65 8.73 39.37 4.23 49.96 32.91 15.55 53.30 49.92 34.41
 CLIPF frequency 16 39.47 64.43 32.55 10.15 45.51 9.22 2.43 21.70 55.62 70.69 23.08 2.77 9.56 83.00 21.02 33.30 7.20 27.47 4.12 50.24 37.72 13.70 52.70 49.92 35.96
-CLIPA truncation 8 36.28 61.99 32.24 8.65 44.40 5.56 2.04 18.56 45.67 65.93 22.68 2.52 10.51 87.96 34.80 29.37 3.35 46.06 4.59 51.55 32.38 12.41 53.46 50.08 30.30
+FLIP truncation 8 36.28 61.99 32.24 8.65 44.40 5.56 2.04 18.56 45.67 65.93 22.68 2.52 10.51 87.96 34.80 29.37 3.35 46.06 4.59 51.55 32.38 12.41 53.46 50.08 30.30
 random 8 45.10 68.54 34.04 9.03 51.54 8.64 3.23 30.53 59.29 70.06 27.29 4.21 9.82 90.46 28.72 32.54 5.99 30.61 7.95 58.00 42.90 12.87 53.40 50.08 37.25
-block 8 43.70 66.66 32.62 8.08 51.07 9.27 2.54 28.99 57.88 70.89 26.72 2.65 18.10 90.84 31.88 36.94 9.02 33.36 7.52 61.91 40.10 14.66 51.80 53.27 37.88
+CLIPA block 8 43.70 66.66 32.62 8.08 51.07 9.27 2.54 28.99 57.88 70.89 26.72 2.65 18.10 90.84 31.88 36.94 9.02 33.36 7.52 61.91 40.10 14.66 51.80 53.27 37.88
 syntax 8 37.92 65.25 31.72 8.28 45.74 8.73 2.96 17.07 46.54 68.73 23.52 2.59 11.30 88.30 29.90 30.81 6.47 30.82 5.02 54.62 36.58 12.51 52.74 50.08 34.97
 CLIPF frequency 8 45.75 59.00 30.76 10.60 50.06 10.32 3.20 29.89 67.54 73.35 26.96 4.09 10.10 90.54 20.80 36.30 7.16 40.31 7.10 57.23 43.09 12.27 54.54 50.03 39.30
-CLIPA truncation 6 23.54 60.27 23.52 5.70 36.54 3.82 2.03 10.32 31.12 61.15 17.77 1.86 9.94 85.25 20.62 23.68 7.13 16.71 3.45 50.23 24.53 12.61 53.62 50.08 23.28
+FLIP truncation 6 23.54 60.27 23.52 5.70 36.54 3.82 2.03 10.32 31.12 61.15 17.77 1.86 9.94 85.25 20.62 23.68 7.13 16.71 3.45 50.23 24.53 12.61 53.62 50.08 23.28
 random 6 38.45 71.57 33.89 7.73 51.81 6.67 2.72 29.47 59.92 68.65 25.89 2.42 12.28 89.73 26.90 33.17 8.89 29.75 7.23 56.87 38.62 12.63 51.96 50.08 34.62
-block 6 39.87 66.60 35.31 7.58 53.23 7.49 2.89 27.87 56.74 69.40 25.97 2.95 10.06 89.30 27.90 32.40 7.31 33.76 6.73 63.10 39.20 12.65 55.41 52.83 35.89
+CLIPA block 6 39.87 66.60 35.31 7.58 53.23 7.49 2.89 27.87 56.74 69.40 25.97 2.95 10.06 89.30 27.90 32.40 7.31 33.76 6.73 63.10 39.20 12.65 55.41 52.83 35.89
 syntax 6 34.18 63.46 29.83 7.54 43.54 7.92 1.85 17.23 44.97 68.37 22.75 1.99 11.37 89.19 17.56 26.13 3.84 33.16 4.98 55.25 35.40 12.30 53.44 50.80 32.57
 CLIPF frequency 6 45.33 67.64 32.83 9.98 50.86 9.61 2.19 31.33 66.03 69.24 26.60 2.94 10.71 90.98 33.64 33.22 8.93 36.70 7.32 59.96 40.60 12.45 54.15 50.08 37.81
-CLIPA truncation 4 20.64 51.67 22.10 4.19 37.16 3.11 1.18 15.16 29.36 53.99 16.17 1.04 10.10 86.34 17.16 15.81 7.48 30.15 3.49 61.43 24.95 12.63 50.76 50.08 19.80
+FLIP truncation 4 20.64 51.67 22.10 4.19 37.16 3.11 1.18 15.16 29.36 53.99 16.17 1.04 10.10 86.34 17.16 15.81 7.48 30.15 3.49 61.43 24.95 12.63 50.76 50.08 19.80
 random 4 28.82 61.90 28.98 5.35 46.39 4.03 1.42 25.16 43.26 62.87 22.03 2.10 12.24 88.54 23.82 24.46 8.07 39.30 5.59 50.62 33.52 12.65 51.86 50.08 27.13
-block 4 28.08 60.63 27.10 5.09 45.61 3.82 2.43 23.24 36.63 64.15 21.80 2.29 8.41 87.04 24.44 23.95 7.55 43.05 5.51 55.80 32.62 9.49 53.55 50.08 26.66
+CLIPA block 4 28.08 60.63 27.10 5.09 45.61 3.82 2.43 23.24 36.63 64.15 21.80 2.29 8.41 87.04 24.44 23.95 7.55 43.05 5.51 55.80 32.62 9.49 53.55 50.08 26.66
 syntax 4 25.59 54.13 27.46 4.73 42.26 6.26 2.66 15.27 32.52 58.93 19.58 2.20 10.42 90.11 26.02 20.62 5.19 37.50 4.32 51.12 26.82 12.93 52.77 50.08 24.61
 CLIPF frequency 4 36.09 69.76 31.14 8.41 47.72 7.01 2.24 25.64 54.23 64.55 23.55 2.95 10.18 90.25 20.78 26.52 6.90 36.63 6.40 55.71 35.71 12.65 49.63 50.08 30.93
-6.4 Zero-shot Classification on More Datasets
-In Table 8, we present evaluation results on additional datasets.
+Table 6: Zero-shot accuracy on more classification datasets.
+Comparison of the zero-shot accuracy performance of CLIP, FLIP, CLIPA, and CLIPF on various datasets.
+The models are pre-trained on CC12M [3] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
+6.3 Zero-shot Image-Text Retrieval
+Tab.
+5 shows image-text retrieval results on Flickr30k [44] and COCO [24].
+In general, we see that text masking approaches are able to improve over FLIP.
+When using 16 text tokens, all strategies except the truncation method outperform FLIP, which is pre-trained on the full text.
+Here, we do not see consistent improvement over CLIP, although some cases come close.
+Note again that we are testing here only with 49 image tokens and 98 would fare better (see Tab.
+8 of the supplementary material), with additional computational cost.
+In contrast to the previous experiments, with image-text retrieval, we observe that CLIPF does not clearly dominate the other text masking strategies.
+The CLIPA random, CLIPA block, and CLIPF (frequency) strategies exhibit comparable performance in image and text retrieval across both datasets.
+The notable exception is the performance of CLIPF in the case that the text masking ratio is very high, i.e., only 4 input text tokens are used.
+Here, the benefit of CLIPF’s frequency-based approach is clearly evident.
+We point out that for image-text retrieval the performance of syntax masking is particularly low in comparison to the other forms of masking.
+Specifically, when using 8 and 6 text tokens, the performance gap between syntax masking and block, random, and frequency masking is approximately 9% on COCO and 6% on Flickr30k for image-text retrieval.
+This gap is wider than what was observed with zero-shot classification.
+We conjecture that the bad performance is due to the image-text retrieval task requiring understanding the overall meaning and logic of the text rather than just recognizing nouns or objects.
+For this reason, the fact that syntax masking retains a high number of nouns, as shown in  Sec.
+3, is particularly harmful.
+6.4 Zero-shot Classification on Additional Datasets
+In  Tab.
+6, we present evaluation results on additional datasets.
 With six text tokens, CLIPF consistently outperforms the original CLIP (no masking), FLIP (image masking with full-text), and other text masking strategies.
-Aligned with ImageNet findings, reducing text tokens from 16 to 8 significantly boosts performance.
+Aligned with ImageNet findings, reducing text tokens from 16 to 8 substantially boosts performance.
 Improvements include SUN397 [42] (+6.39%), DTD [6] (+9.31%), OxfordPets [31] (+4.64%), STL10 [7] (+9.31%), EuroSAT [11] (+6.60%), UCF101 [37] (+6.68%).
@@
 We point out that the Describable Textures Dataset (DTD)[6] contains 47 categories of texture patches, whose POS is mostly verb and that the OxfordPets dataset[31] also contains many class names that are not nouns.
-On these data sets the performance of syntax masking is particularly low.
+On these data sets, the performance of syntax masking is particularly low.
 Specifically, when using six text tokens, syntax masking underperforms CLIPF by 14.10% on DTD and 21.04% on OxfordPets.
-This low performance reflects that the VLM has overfitted on the large number of nouns retained by syntax masking.
-6.5 Scaling to the larger Dataset.
-Finally, we carry out an exploratory experiment to gain insight into different masking strategies with data sets of different sizes.
-We carry out the largest experiment that is possible given our computational capacity, which was training on the LAION-400M [17] data set, using 4 text tokens, a smaller image size (112x112) and 50% image masking.
-We trained for 16 epochs.
-Previous works have shown that both smaller image sizes and image masking are effective for accelerating CLIP pre-training [21, 8, 45].
-Results are shown in Fig.
-4 for both top-1 accuracy (left) and top-5 accuracy (right).
-We see from these graphs that using $<$ 10 epochs would not provide a complete picture on how the different masking approaches perform relative to each other.
-Overall, CLIPF, the frequency based approach dominates.
-However, syntax is performing better than it did on smaller data sets, possibly because the large amount of data is compensating for the overemphasis of syntax masking on nouns.
-It is important not to conclude on the basis of this experiment that syntax masking will continue to dominate block masking.
-Looking at the top-5 performance, we see that these two approaches near each other already in the 16 epochs we are able to compute.
-In sum, this experiment shows that CLIPF is a good performer as data scales up, but that the number of epochs is important in understanding the strength of different masking strategies.
-Figure 4: Zero-shot classification accuracy on ImageNet-1K over training epochs for CLIPF and CLIPA strategies when pre-training on LAION-400M datasets.
-We pre-trained the model using 25 image tokens and 4 text tokens for 16 epochs.
+This low performance reflects that the VLM has overfitted on the large number of nouns retained by syntax masking and mistakes in POS detection (which is challenging for OxfordPets).
+Note that, unlike the performance on ImageNet (see Figs.
+4 and 5 of the supplementary material), syntax masking on DTD remains worse than random, block, and frequency masking throughout all epochs.
 7 Conclusion and Outlook
-This paper has provided a systematic overview of text masking approaches for VLM training, which accelerate training and have the potential to also improve performance.
-In this paper, we demonstrate that training epochs have a substantial impact on the performance of different text masking strategies.
-We conclude, contrary to the existing state-of-the-art, CLIPA [12], that syntax masking is not the best performing text masking approach and that truncation masking can be avoided.
-We carry out an analysis of word frequency distributions that shows that frequency is what you need to think about when text masking the training data of VLMs.
-Specifically, both syntax and truncation have visible effects on the word frequency distribution, making clear why they do not perform well.
-We have introduced CLIPF, an approach for text masking based on word frequency that improves the performance of CLIP.
-Unlike syntax-based masking methods, CLIPF does not require part-of-speech (POS) information, which is time-consuming to process.
-CLIPF is also interesting because it controls frequency directly and in doing so seems to balance nouns with words in the “Other” POS categories, in contrast to other approaches.
-We carry out an extensive comparison among existing text matching strategies (CLIPA variants) and CLIPF that revals the strong performance of CLIPF in a variety of zero-shot classification tasks and when scaling to larger data sets.
-CLIPF demonstrates a notable ability to rapidly learn image and text representations when compared to a model that pre-trains on the full text and other text masking strategies.
-It also delivers strong performance in under highly aggressive masking (i.e., using only 4 tokens per training text).
-Moving forward, the insights gained from this paper about the importance of frequency can be used to further optimize CLIPF to directly optimize the word distribution allowing for more tuning control when trading off VLM performance and VLM training speech during text masking.
-Appendix A Text Masking Analysis
-A.1 Text Masking Cases
-As illustrated in Fig.
-5, there is a clear relationship between word masking probability and word frequency.
-Frequent words have a higher masking probability compared to infrequent ones.
-Additionally, a smaller threshold $t$ leads to a smaller difference between the masking probabilities of frequent and infrequent words.
-Therefore, it was necessary to choose a relatively larger threshold to ensure that both frequent and infrequent words are effectively masked.
-Figure 5: The curve of Formula 2. The x-axis is the word frequency $f(w_{i})$, and the y-axis is the $P(w_{i})$. The value of $t$ of formular 2 is set to $10^{-5}$, $10^{-6}$, $10^{-7}$,$10^{-8}$.
-Table  9 presents the potential text resulting from the text masking technique.
-Since truncation is fixed in each epoch, it may result in the loss of important information at the end of the text.
-Certain words tend to appear in specific positions within the text; for example, “the” and “a” are most likely to be the first words of a sentence.
-As shown in Fig.
-3, truncation retains more occurrences of “the” and “a” than other text masking strategies.
-In contrast, random and block strategies can generate different texts in each epoch for text data augmentation, but they may retain some words with a high frequency.
-Syntax masking retains the nouns in the text and remains the same in each epoch.
-However, this strategy may cause the model to overfit on the frequency of noun words.
-In contrast, CLIPF varies the text in each epoch, as words may be retained or removed according to their frequency.
-This strategy serves two primary purposes: it enhances text diversity and reduces the risk of overfitting to frequent words.
-Another advantage of CLIPF is that it can remove frequent prepositions that are less directly relevant to the objects in the image, such as “a,” “in,” “of,” and others.
-This helps the model focus on the most helpful aspects of the content.
-Table 10 shows the masking probabilities for certain words. High-frequency words such as “of" “the" “and" and “." are highly likely to be masked from the text.
-Table 9: Example from the CC12M dataset illustrating the effect of various text masking strategies, reducing text length to 6 words.
-The original caption is in the first row, followed by masked variants.
-Parameter $t$ in Equation 2 is set to $10^{-6}$.
+This paper has provided a systematic overview of text masking approaches for VLM training, which can boost performance and reduce computation.
+VLMs are important because of their wide use in practical applications, including improving Multimodal Large Language Models (MLLMs) [18, 25].
+We have carried out an analysis of how text masking impacts word frequency in the training data and discovered that text masking strategies perform better if they do not disproportionately impact word frequency.
+Our experiments also demonstrate that training epochs have a substantial impact on the performance of different text masking strategies.
+We conclude, contrary to the existing state-of-the-art, CLIPA [20], that syntax masking is not the best performing text masking approach.
+Further, this paper has introduced CLIPF, a strategy for text masking based on word frequency.
+Our experiment shows that CLIPF yields strong performance and shows the ability to outperform the original CLIP, while using less training resources.
+Unlike syntax-based masking methods, CLIPF does not require part-of-speech (POS) information, which consumes more resources to generate than frequency statistics.
+CLIPF is also interesting because it seems to balance nouns with words in the “Other” POS category, in contrast to syntax masking.
+We carry out a quantitative comparison among CLIPF and existing text matching strategies that reveals the strong performance of CLIPF in a large range of zero-shot classification tasks.
+CLIPF demonstrates a notable ability to learn image and text representations when compared to a model pre-trained on text masked with other existing text masking strategies.
+It also outperforms other approaches under highly aggressive masking (i.e., using only 4 tokens per training text).
+This paper has argued that frequency is what you need to think about when choosing a text masking strategy.
+Specifically, both truncation and syntax masking have a disproportionate impact on the frequencies of specific words, making clear why they do not perform well.
+Further, CLIPF, which exploits word-frequency directly, is a strong performer.
+These insights are a valuable guide for researchers and developers, who need to select an appropriate text masking strategy, while taking in to account available computational resources.
+Moving forward, CLIPF can be further optimized to more directly shape the word frequency distribution, providing even more control when trading off VLM performance and computation during text masking.
+Appendix A Details of Experimental Setup
+Our implementation follows CLIP [32], OpenCLIP [15], FLIP [22] and CLIPA [20]. In this section, we present the details of our experimental setup.
+Dataset Data set statistics are summarized in  Tab.
+7.
+For the majority of our experiments, we pre-train the models on Conceptual Captions 3M (CC3M) [36], Conceptual 12M (CC12M) [3].
+These datasets have been used by OpenCLIP, DeCLIP, SLIP, A-CLIP and BLIP [15, 21, 30, 43, 18] To reproduce CLIPA, we use LAION-400M [35], which is used by CLIPA.
+LAION-400M is approximately 133 times larger than CC3M and 32 times larger than CC12M, indicating that it contains a broader range of concepts and offers greater diversity compared to CC3M and CC12M.
+Note that the difference in caption length with have an impact on masking performance, although we do not directly measure the difference in this paper.
+Architecture For the image encoder, we use ViT-B/16 (86M parameters) [9] architectures with global average pooling and learnable positional embeddings.
+For the text encoder, we use a Transformer-based model [40] and byte-pair encoding with a 49K token vocabulary.
+Additionally, we run one experiment with a larger image encoder (VIT-L/16) with 303M parameters and report the results in the Appendix C of supplementary material.
+We chose these encoders because they are the largest ones available to us, given our current resource constraints.
+The input image size is 224 for all datasets.
+When using full text for training, the maximum context length is 32.
+Zero-padding is applied to input text that is shorter than the maximum token length of the model.
+We trained the models using 8 RTX A5000 GPUs with the same settings to ensure consistent conditions across all models.
+We experiment with different text token input sizes, namely, 32, 16, 8, 6, and 4 text tokens.
+As the input size gets smaller, we can increase the batch size, maximizing computational memory usage.
+For CC3M and CC12M, the batch sizes corresponding to the input sizes are: 664, 832, 896, 928, and 944.
+Note that CLIP (and therefore CLIPA and CLIPF) is trained using contrastive learning, which benefits from larger batch sizes.
+Note also that although we used different text masking with the same settings and batch sizes, syntax masking was slower than the other text masking strategies when conducting the experiments because POS processing for each word is time-consuming.
+We pre-train the model using the InfoNCE loss [39] with a learnable temperature parameter $\tau$ [5, 32]. To classify images, we calculate the cosine similarity between the image and text embeddings.
+Dataset Samples Total words Caption length
+CC3M $2.72\times 10^{6}$ $2.80\times 10^{7}$ $10.30\pm 4.72$
+CC12M $9.30\times 10^{6}$ $06\times 10^{8} 22.$$2.15\pm 17.20$
+LIAON-400M $2.98\times 10^{8}$ $3.71\times 10^{9}$ $12.51\pm 15.82$
+Table 7: Dataset statistics for pre-training datasets. Caption length refers to the number of words in the text.
+Config Pre-training Fine-tuning
+optimizer AdamW [27] AdamW [27]
+learning rate 1e-3 1e-5
+weight decay 0.2 0.2
+optimizer momentum $\beta_{1},\beta_{2}=0.9,0.98$ [4] $\beta_{1},\beta_{2}=0.9,0.98$ [4]
+learning rate schedule cosine decay [26] cosine decay [26]
+warmup steps 10k 10%
+epoch 30 1
+$t$ $10^{-6}$
+$\tau$ 0.07
+numerical precision amp amp
+RandomResizedCrop (50, 100) (50, 100)
+Table 8: Details of the pre-training and fine-tuning setups.
+Image masking Text masking Image tokens Text tokens Total Percentage
+0% 0.00% 196 32 228 100%
+75% 0.00% 49 32 81 35.53%
+50.00% 49 16 65 28.51%
+75.00% 49 8 57 25.00%
+81.25% 49 6 55 24.12%
+87.50% 49 4 53 23.25%
+Table 9: The number of image and text tokens that are processed during pre-training (CC3M and CC12M).
+Training Following CLIP, OpenCLIP [32, 15], we pre-train our model for 30 epochs on the CC3M and CC12M datasets.
+For the LAION-400M dataset, we pre-train the model for 16 epochs, extending CLIPA’s experiments, which were conducted for 6 epochs on the same dataset.
+Details of the pre-training configuration are given in Tab.
+8.
+[…]
+During text masking, we reduce the number of tokens from 32 to 16, 8, 6, and 4.
+To measure the training speed of CLIP, CLIPA, and CLIPF, we compare the number of text tokens processed by each model.
+As shown in  Tab.
+9, the number of text tokens processed by each model during pre-training is calculated based on different image and text masking ratios.
+When we pre-trained the model using image masking, the total number of tokens is 81 and the percentage of text tokens is 39.5%.
+When we apply 50% text masking, the total number of tokens is 65.
+Compared to training without text masking, this results in a speed increase of approximately 20%.
+Moreover, when we apply 87.5% text masking, the total number of tokens is 53, resulting in a speed increase of approximately 34% compared to training without text masking.
+Fine-tuning Following FLIP, and CLIPA, in order to bridge the distribution gap between pre-training and evaluation, we fine-tuned the model without images and text masking.
+Note that, fine-tuning the models without masking in masking work is focused on reducing the distribution gap and applied sparingly to avoid undoing the tradeoff.
+The details of the fine-tuning configuration are provided in  Tab.
+8.
+Figure 5: The curve of Equation 2. The x-axis is the word frequency $f(w_{i})$, and the y-axis is the $P(w_{i})$. The value of $t$ of Equation 2 is set to $10^{-5}$, $10^{-6}$, $10^{-7}$$10^{-8}$.
 Method Resulting Text
@@
 CLIPF siberian handsome man hugging smiling red
-Table 10: The probability of masking words in the text is calculated using the formula provided in Equation 2.
-The example is selected from CC12M [15].
-The value of $t$ of formular 2 is set to $10^{-6}$.
+Table 10: Example from the CC12M dataset illustrating the effect of various text masking strategies, reducing text length to 6 words.
+The original caption is in the first row, followed by masked variants.
+Parameter $t$ in Equation 2 is set to $10^{-6}$.
 Word Probability
@@
 and 0.991695
-siberian 0.75092
+siberian 0.750920
 dog 0.960531
 . 0.993678
-A.2 Word Category Distribution Across Different Text Masking Strategies
-As shown in Table 11, we provide the word type counts of Table 5.
-After applying text masking strategies such as truncation, random, block, syntax, and CLIPF, the word counts remain similar.
-However, SW-CLIP does not maximize the use of the input slots, utilizing only 85% of the words compared to other text masking strategies.
-Additionally, SW-CLIP masks a high percentage of other types of words, which may impact the model’s zero-shot ability.
-Table 11: Distribution of syntax counts and percentages before and after applying text masking. The dataset is CC12M and we retain 6 words for each text after applying text masking.
-Masking NN Count JJ VB OTHER NN (%) JJ (%) VB (%) OTHER (%) Total
+Table 11: The probability of masking words in the text is calculated using the formula provided in Equation 2.
+The example is selected from CC12M [3].
+The value of $t$ of Equation 2 is set to $10^{-6}$.
+Figure 6: The complete x-axis label for Fig.
+3.
+in Sec.
+3 We set the text length after text masking to 6.
+The x-axis represents the word index, which is sorted by counts of the original data, and the y-axis shows the word frequency.
+The dataset used is CC12M and the value of $t$ of Equation 2 is set to $10^{-6}$.
+We remove special characters from the vocabulary.
+Masking NN Count $J$ VB OTHER NN (%) JJ (%) VB (%) OTHER (%) Total
 Before masking 103,469,117 10,245,828 10,649,943 81,351,966 50.30% 4.98% 5.18% 39.55% 205,716,854
@@
 CLIPF 33,516,439 2,803,409 3,473,119 15,666,310 60.43% 5.06% 6.20% 28.24% 55,459,277
-Figure  6 shows the word cloud before and after applying text masking.
-In these visuals, the font size represents the relative frequency of each word.
-Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency.
-After applying frequency word masking, the word cloud in Table 6(b) shows a more balanced distribution of word frequencies compared to before the masking.
-The word “the" appears smaller than in other word clouds.
-The reduction in frequent words allows infrequent words to become more prominent.
-The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full-text.
-As depicted in Table 6(f), syntax masking tends to retain more nouns such as “person," “stock," and “illustration" that show a different pattern from the others.
-Fig.
-6 shows word clouds before and after applying text masking.
-In these visuals, the font size represents the relative frequency of each word.
-Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency.
-After applying word frequency masking, the word cloud in Fig.
-6(b) displays a more balanced distribution of word frequencies compared to before masking.
-The word “the” appears smaller than in other word clouds.
-The reduction in frequent words allows infrequent words to become more prominent.
-The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full text.
-As depicted in Fig.
-6(f), syntax masking tends to retain more nouns such as “person”, “stock”, and “illustration”, resulting in a different pattern from the others.
-Refer to caption Refer to caption
-(a) Before masking (b) Frequency
-Refer to caption Refer to caption
-(c) Truncation (d) Random
-Refer to caption Refer to caption
-(e) Block (f) Syntax
-Figure 6: The word cloud on the left displays the full text, while the one on the right shows the text after applying text masking, where we retain 6 words.
-The font size of the word represents the relative word frequency for each figure.
-The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$.
-A.3 Word Distribution Analysis for CLIPF
+Table 12: Distribution of syntax counts and percentages before and after applying text masking. The dataset is CC12M and we retain 6 words for each text after applying text masking.
+Evaluation setup Following CLIP [32], FLIP [22], and CLIPA [20], several classification benchmarks were used.
+Among these benchmarks is ImageNet-1K, a widely recognized dataset in computer vision.
+It is frequently used for image classification and VLM tasks and comprises 50K validation samples across 1K different classes.
+We filled each class into the templates provided by CLIP [32] to calculate the average of the text embeddings.
+We use the same evaluation settings as CLIP [32] to evaluate the other downstream datasets.
+Appendix B Text Masking Analysis
+B.1 Text Masking Cases
+As illustrated in  Fig.
+5, there is a clear relationship between word masking probability and word frequency.
+Frequent words have a higher masking probability compared to infrequent ones.
+Additionally, a smaller threshold $t$ leads to a smaller difference between the masking probabilities of frequent and infrequent words.
+Therefore, it was necessary to choose a relatively larger threshold to ensure that both frequent and infrequent words are effectively masked.
+Tab.
+10 presents the potential text resulting from the text masking technique.
+Since truncation is fixed in each epoch, it may result in the loss of important information at the end of the text.
+Certain words tend to appear in specific positions within the text; for example, “the” and “a” are most likely to be the first words of a sentence.
+As shown in  Fig.
+6, truncation retains more occurrences of “the” and “a” than other text masking strategies.
+In contrast, random and block strategies can generate different texts in each epoch for text data augmentation, but they may retain some words with a high frequency.
+Syntax masking retains the nouns in the text and remains the same in each epoch.
+However, this strategy may cause the model to overfit on the frequency of noun words.
+In contrast, CLIPF varies the text in each epoch, as words may be retained or removed according to their frequency.
+This strategy serves two primary purposes: it enhances text diversity and reduces the risk of overfitting to frequent words.
+Another advantage of CLIPF is that it can remove frequent prepositions that are less directly relevant to the objects in the image, such as “a,” “in,” “of,” and others.
+This helps the model focus on the most helpful aspects of the content.
+Tab.
+11 shows the masking probabilities for certain words.
+High-frequency words such as “of” “the” “and” and “.” are highly likely to be masked from the text.
+B.2 Word Category Distribution Across Different Text Masking Strategies
+As shown in  Tab.
+12, we present the word type counts corresponding to Tab.
+1 in the main paper.
+After applying text masking strategies such as truncation, random, block, syntax, and CLIPF, the word counts remain similar.
+However, SW-CLIP does not maximize the use of the input slots, utilizing only 85% of the words compared to other text masking strategies.
+Additionally, SW-CLIP masks a high percentage of other types of words, which may impact the model’s zero-shot ability.
+Figure 7: Frequency-based text masking strategies vary according to the number of text words used during pre-training. The dataset used is CC12M and the value of $t$ of Equation 2 is set to $10^{-6}$.
+Models Masking Image Tokens Text Tokens ViT-B/16
+IN-A IN-O IN-R IN-S IN-V2 ON
+CLIP 197 32 8.97 37.85 49.11 25.70 31.48 24.20
+CLIPF frequency 98 8 10.37 37.75 52.33 28.52 35.62 24.00
+Table 13: Zero-shot robustness evaluation.
+Comparison of the zero-shot accuracy performance of CLIP and CLIPF on various datasets when using more image tokens.
+The models are pre-trained on CC12M [3] for 30 epochs with image masking (50%) to speed up training and fine-tune the model an additional epoch without image and text masking.
+Models Masking Image Tokens Text Tokens Text Retrieval Image Retrieval
+Flickr30k COCO Flickr30k COCO
+R@1 R@1 R@1 R@1
+CLIP 197 32 62.62 35.54 45.42 24.22
+CLIPF frequency 98 8 63.11 37.14 46.84 24.89
+Table 14: Zero-shot Image-Text retrieval evaluation.
+We evaluated CLIP, CLIPF image-text retrieval performance on the COCO and Flickr30k datasets when using more image tokens.
+The models are pre-trained on CC12M [3] for 30 epochs with image masking (50%) to speed up training and fine-tune the model an additional epoch without image and text masking.
+B.3 Word Distribution Analysis for CLIPF
 In addition, we analyzed frequency text masking strategies that varied according to the number of words used, as shown in  Fig.
@@
 This configuration helps achieve a balanced word distribution, which is beneficial for pre-training VLMs.
-Figure 7: Frequency-based text masking strategies vary according to the number of text words used during pre-training. The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$.
-A.4 Word Distribution of CLIPF and SW-CLIP
-We also compare the word distributions of CLIPF and SW-CLIP after applying text masking.
-As illustrated in Fig.
-8, SW-CLIP masks more frequent words based on its threshold, preventing it from fully utilizing the input slots of the text encoder.
-Consequently, SW-CLIP exhibits a flatter word distribution compared to CLIPF.
-Figure 8: The figure illustrates the distribution of top-50 words in the text for SW-CLIP and CLIPF.
-Appendix B More Results
-More detailed results are presented in this section, including the results and learn curve of image-text retrieval.
-B.1 The details of Image-text Retrieval
-In Table 12, we show more details of Zero-shot Image-Text Retrieval on COCO and Flickr30k datasets.
+Models Masking Image Tokens Text Tokens Text Retrieval Image Retrieval
+Flickr30k COCO Flickr30k COCO
+R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
+CLIP 197 32 62.62 86.00 91.81 35.54 62.38 74.08 45.42 72.56 81.50 24.22 48.42 60.42
+FLIP 49 32 54.73 80.37 87.97 29.34 56.08 67.00 38.62 66.09 75.40 20.88 43.22 54.78
+FLIP truncation 16 44.67 73.08 81.85 25.54 51.90 64.64 34.99 61.05 70.85 19.64 41.91 53.51
+random 58.48 84.62 90.53 32.36 58.76 69.98 43.61 70.67 80.04 22.81 46.00 57.71
+CLIPA block 56.51 81.26 89.05 30.82 58.32 70.38 44.06 71.20 80.10 22.66 46.24 58.06
+syntax 54.54 81.07 88.36 29.60 56.52 68.54 41.16 68.32 77.51 21.40 44.82 56.56
+CLIPF frequency 57.89 84.62 90.04 31.52 58.38 70.30 42.72 69.61 78.60 22.57 46.15 57.95
+FLIP truncation 8 30.47 59.47 70.51 16.92 38.62 51.34 23.96 47.29 57.85 12.78 31.66 42.92
+random 58.58 84.52 91.81 36.24 62.16 72.90 43.79 70.89 80.14 23.16 46.74 58.86
+CLIPA block 60.06 85.01 92.11 35.88 62.58 73.74 45.98 73.23 82.49 24.65 48.81 60.70
+syntax 50.30 78.30 86.98 29.64 55.42 67.18 38.46 66.77 77.36 20.28 43.78 55.40
+CLIPF frequency 58.68 85.10 91.51 34.74 61.38 71.88 44.57 72.58 81.92 23.30 47.50 59.24
+FLIP truncation 6 32.25 61.05 72.98 16.18 39.52 52.48 22.39 47.46 59.39 11.93 30.03 40.86
+random 55.23 82.54 89.25 32.58 57.54 68.68 42.47 70.12 79.43 21.34 44.79 56.26
+CLIPA block 54.44 79.68 88.66 33.12 58.96 70.64 41.68 69.13 79.33 21.73 45.23 57.29
+syntax 46.15 76.04 84.81 26.78 52.56 64.62 34.08 61.76 73.25 17.99 40.22 52.05
+CLIPF frequency 56.02 82.05 88.56 32.32 58.58 70.00 41.05 69.17 78.88 21.28 44.55 56.05
+FLIP truncation 4 30.97 56.51 67.06 17.44 38.48 49.76 21.09 43.77 55.70 11.07 27.98 38.54
+random 41.62 69.53 80.18 24.14 48.32 60.14 30.26 56.92 68.42 15.63 35.72 47.30
+CLIPA block 40.83 69.03 79.09 23.56 47.94 59.20 30.77 57.85 69.29 14.98 35.64 47.35
+syntax 38.66 65.88 75.74 21.68 44.38 56.60 26.08 52.80 64.32 14.14 33.71 44.98
+CLIPF frequency 45.07 72.49 81.95 25.60 49.98 61.96 31.72 59.35 71.66 15.96 36.58 48.32
 Table 15: Zero-shot Image-Text Retrieval, we evaluated CLIP, FLIP, and CLIPF image-text retrieval performance on COCO and Flickr30k datasets.
 The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
-Models Masking Text Tokens Text Retrieval Image Retrieval
-Flickr30k COCO Flickr30k COCO
-R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
-CLIP 32 62.62 86.00 91.81 35.54 62.38 74.08 45.42 72.56 81.50 24.22 48.42 60.42
-FLIP 32 54.73 80.37 87.97 29.34 56.08 67.00 38.62 66.09 75.40 20.88 43.22 54.78
-CLIPA truncation 16 44.67 73.08 81.85 25.54 51.90 64.64 34.99 61.05 70.85 19.64 41.91 53.51
-random 16 58.48 84.62 90.53 32.36 58.76 69.98 43.61 70.67 80.04 22.81 46.00 57.71
-block 16 56.51 81.26 89.05 30.82 58.32 70.38 44.06 71.20 80.10 22.66 46.24 58.06
-syntax 16 54.54 81.07 88.36 29.60 56.52 68.54 41.16 68.32 77.51 21.40 44.82 56.56
-CLIPF frequency 16 57.89 84.62 90.04 31.52 58.38 70.30 42.72 69.61 78.60 22.57 46.15 57.95
-CLIPA truncation 8 30.47 59.47 70.51 16.92 38.62 51.34 23.96 47.29 57.85 12.78 31.66 42.92
-random 8 58.58 84.52 91.81 36.24 62.16 72.90 43.79 70.89 80.14 23.16 46.74 58.86
-block 8 60.06 85.01 92.11 35.88 62.58 73.74 45.98 73.23 82.49 24.65 48.81 60.70
-syntax 8 50.30 78.30 86.98 29.64 55.42 67.18 38.46 66.77 77.36 20.28 43.78 55.40
-CLIPF frequency 8 58.68 85.10 91.51 34.74 61.38 71.88 44.57 72.58 81.92 23.30 47.50 59.24
-CLIPA truncation 6 32.25 61.05 72.98 16.18 39.52 52.48 22.39 47.46 59.39 11.93 30.03 40.86
-random 6 55.23 82.54 89.25 32.58 57.54 68.68 42.47 70.12 79.43 21.34 44.79 56.26
-block 6 54.44 79.68 88.66 33.12 58.96 70.64 41.68 69.13 79.33 21.73 45.23 57.29
-syntax 6 46.15 76.04 84.81 26.78 52.56 64.62 34.08 61.76 73.25 17.99 40.22 52.05
-CLIPF frequency 6 56.02 82.05 88.56 32.32 58.58 70.00 41.05 69.17 78.88 21.28 44.55 56.05
-CLIPA truncation 4 30.97 56.51 67.06 17.44 38.48 49.76 21.09 43.77 55.70 11.07 27.98 38.54
-random 4 41.62 69.53 80.18 24.14 48.32 60.14 30.26 56.92 68.42 15.63 35.72 47.30
-block 4 40.83 69.03 79.09 23.56 47.94 59.20 30.77 57.85 69.29 14.98 35.64 47.35
-syntax 4 38.66 65.88 75.74 21.68 44.38 56.60 26.08 52.80 64.32 14.14 33.71 44.98
-CLIPF frequency 4 45.07 72.49 81.95 25.60 49.98 61.96 31.72 59.35 71.66 15.96 36.58 48.32
-B.2 Learn Curved for Image-text Retrieval
-We also investigate the zero-shot image-text retrieval performance over training epochs on the COCO and Flickr30k datasets.
-It also shows that the number of training epochs has a substantial impact on the performance, as illustrated in Fig.
-9.
-And the performance gap between syntax masking and the other text masking on the zero-shot image-text retrieval task is larger than zero-shot classification task.
-This is because the zero-shot image-text retrieval task requires the model not only to understand the objects/nouns, but also to understand the logic of the text.
-Moreover, we present the zero-shot image-text retrieval performance across training epochs for models pre-trained on the LAION-400M dataset.
-As illustrated in Fig.
-10, the performance patterns of models trained on LAION-400M are similar to those of previous models that pre-trained on the CC12M dataset.
-Specifically, CLIPF consistently outperforms syntax masking.
-Block masking achieves comparable performance to syntax masking and even surpasses it on the image-to-text retrieval task using the COCO dataset at epoch 16th.
-Figure 9: Zero-shot Image-Text Retrieval accuracy on COCO and Flick30k over training epochs for CLIPF and CLIPA strategies.
-The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model for an additional epoch without image and text masking.
-Figure 10: Zero-shot Image-Text Retrieval accuracy on COCO and Flick30k over training epochs for CLIPF and CLIPA strategies.
-The backbone of the image encoder is ViT-B/16, and the model is pre-trained on LAION-400M for 16 epochs with 25 image tokens and 4 text tokens to speed up training and fine-tune the model for 0.4 epoch with 224x224 image size and 32 text tokens.
-Additionally, Fig.
-11 illustrates the validation loss across pre-training epochs.
-It is evident that the validation loss for truncation and syntax masking methods begins to increase earlier compared to other text masking methods, suggesting that these methods may overfit to the data sooner.
-Figure 11: The validation loss curve of COCO and Flick30k datasets over the pre-training epochs.
-B.3 Learn Curved for DTD and OxfordPets Datasets
-The learning curves for the DTD [40] and OxfordPets [41] datasets are presented because not all class names in these datasets are nouns.
-As shown in Table 13, the performance of syntax masking on the DTD dataset is consistently lower than that of random, block, and frequency masking across all epochs, rather than initially outperforming them and declining at later stages in the ImageNet dataset, as shown in Fig.
-12.
-Moreover, the performance of syntax masking is very poor and almost close to truncation masking.
-OxfordPets is a dataset containing incomplete noun class names.
-In the early training epochs, syntax masking performs similarly to frequency masking and still outperforms truncation, block, and random masking.
-However, in the later stages of training, syntax masking performs much worse than the other text masking strategies.
-Figure 12: Zero-shot classification accuracy on ImageNet-1K dataset over training epochs for CLIPF and CLIPA strategies.
+Models Datasets Image Tokens Text Tokens ViT-B/16
+pre-train fine-tune
+SW-CLIP CC3M 49 6 14.9 16.8
+CLIPF 14.4 18.2
+SW-CLIP CC12M 49 8 35.9 38.5
+CLIPF 36.6 39.3
+Table 16: Comparison of the zero-shot accuracy performance of SW-CLIP and CLIPF on ImageNet-1k.
+The models are pre-trained on CC12M [3] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
+Appendix C More Results
+More detailed results are presented in this section, including the results and learn curve of image-text retrieval.
+C.1 More image tokens
+As shown in Tab.
+13 and Tab.
+14, CLIPF achieves better zero-shot robustness in 4 out of 6 datasets and image–text retrieval performance in both Flick30k and COCO datasets than the original CLIP, even though it only uses 50% of the image tokens and 25% of the text tokens.
+C.2 The details of Image-text Retrieval
+In Tab. 15, we show more details of Zero-shot Image-Text Retrieval on COCO and Flickr30k datasets.
+Figure 8: Zero-shot classification accuracy on DTD dataset over training epochs for CLIPF and CLIPA strategies.
 The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
-Figure 13: Zero-shot classification accuracy on DTD dataset over training epochs for CLIPF and CLIPA strategies.
+Figure 9: Zero-shot classification accuracy on ImageNet-1K dataset over training epochs for CLIPF and CLIPA strategies after fine-tuning.
 The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
@@
 The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
-B.4 Learn Curved before fine-tuning
-In Fig.
-1, we show the learning curve after fine-tuning.
-Moreover, we show the learning curve before fine-tuning in Fig.
-15 (30 epochs) and Fig.
-16 (45 epochs).
-They have a similar pattern to the models before fine-tuning for every text masking strategy.
-Figure 15: Zero-shot classification accuracy on ImageNet-1K over the epochs for the different text masking strategies before fine-tuning.
-The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training.
-Figure 16: Zero-shot classification accuracy on ImageNet-1K over the 45 epochs for the different text masking strategies before fine-tuning.
-The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 45 epochs with image and text masking (75%) to speed up training.
-B.5 Fine-tuning the models with More Epochs
-The purpose of fine-tuning the models without image and text masking is to reduce the distribution gap between pre-training and evaluation.
-As demonstrated in FLIP and CLIPA, a small number of fine-tuning steps is usually sufficient to bridge this gap [8, 12].
-We conducted additional experiments using two fine-tuning epochs, shown in Tab 13, to demonstrate that increasing the number of fine-tuning epochs does not substantially improve performance but instead leads to higher computational costs.
-Table 13: Comparison of CLIPF, CLIP, and CLIPA for zero-shot classification on ImageNet-1K.
-Reported as top-1 accuracy.
-All models use a ViT-B/16 image encoder backbone pre-trained on CC12M for 30 epochs with 75% image masking for speedup (pre-train).
-Then, we fine-tune the model with one epoch (fine-tune-1) and two epochs (fine-tune-2) without image and text masking.
-Models Masking Image Tokens Text Tokens CC12M
-pre-train fine-tune-1 fine-tune-2
-CLIP 197 32 36.6
-FLIP 49 32 32.0 33.7 34.1
-CLIPA truncation 49 8 25.4 28.4 28.5
-random 34.5 36.9 37.1
-block 35.5 37.9 38.0
-syntax 28.5 35.0 35.5
-CLIPF frequency 36.6 39.3 39.4
-B.6 Comparison with SW-CLIP
+Models Masking Image Tokens Text Tokens ViT-L/16
+pre-train fine-tune
+FLIP truncation 49 8 28.4 31.7
+random 36.0 38.3
+CLIPA block 36.8 39.5
+syntax 32.4 37.5
+CLIPF frequency 38.0 40.2
+Table 17: Comparison of the ImageNet-1k zero-shot accuracy performance of ViT-L/16 with different text masking strategies.
+The models are pre-trained on CC12M [3] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
+C.3 Learn Curved for DTD and OxfordPets Datasets
+The learning curves for the DTD [6] and OxfordPets [31] datasets are presented because not all class names in these datasets are nouns.
+As shown in Fig.
+8, the performance of syntax masking on the DTD dataset is consistently lower than that of random, block, and frequency masking across all epochs, rather than initially outperforming them and declining at later stages in the ImageNet dataset, as shown in  Fig.
+9.
+Moreover, the performance of syntax masking is very poor and almost close to truncation masking.
+OxfordPets is a dataset containing incomplete noun class names.
+As shown in Fig.
+10, in the early training epochs, syntax masking performs similarly to frequency masking and still outperforms truncation, block, and random masking.
+However, in the later stages of training, syntax masking performs much worse than the other text masking strategies.
+C.4 Comparison with SW-CLIP
 In this section, we provide a comparison of CLIPF with SW-CLIP [23], which also uses frequency-based sampling but imposes a threshold on the frequency score.
@@
 Subsequently, we fine-tuned both models with an additional epoch without any image or text masking.
-The results in Table 14 show that after fine-tuning CLIPF outperforms SW-CLIP.
-Table 14: Comparison of the zero-shot accuracy performance of SW-CLIP and CLIPF on ImageNet-1k.
-The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
-Models Datasets Image Tokens Text Tokens ViT-B/16
-pre-train fine-tune
-SW-CLIP CC3M 49 6 14.9 16.8
-CLIPF 14.4 18.2
-SW-CLIP CC12M 49 8 35.9 38.5
-CLIPF 36.6 39.3
-B.7 Applying text masking on SigLIP
+The results in Tab.
+16 show that after fine-tuning CLIPF outperforms SW-CLIP.
+C.5 Large Image Encoder Architecture
+We carried out an extra experiment which is apply different text masking strategies on larger architectures which is ViT-L/16.
+The behavior of different architectures is very consistent when applying different text masking strategies, as shown in Tab.
+17.
+C.6 Applying text masking on SigLIP
 We also applied the text masking strategies to SigLIP; the results are presented in  Tab.
-15.
+18.
 CLIPF still achieves superior performance compared to other text masking strategies and without text masking.
 This further supports our conclusion that frequency-based masking is the more effective strategy.
-Table 15: We apply text masking (75%) on SigLIP to pre-train on CC12M for 30 epochs with 75% image masking for speedup pre-training. Reporting the zero-shot classification accuracy on ImageNet-1K.
 Models Masking Image Tokens Text Tokens ViT-B/16
@@
 SigLIP frequency 32.4 34.8
-Appendix C Ablation
-Table 16: We pre-train CLIPF on CC12M dataset across different thresholds in Equation 2.
+Table 18: Comparison of the ImageNet-1k zero-shot accuracy performance of SigLIP with different text masking strategies.
 The models are pre-trained on CC12M [3] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
+Appendix D Ablation
 model Image tokens Text tokens Threshold ViT-B/16
@@
 1e-7 36.1 38.6
-C.1 Threshold Analysis
-To investigate the impact of the threshold $t$ in Equation 2 on model performance, we pre-trained models with varying thresholds.
-As shown in Table 16, thresholds of $51e-$ 5, $61e-$ 6, and $71e-$ 7 yield comparable results, all outperforming the other text masking strategies.
-This indicates that CLIPF is relatively insensitive to small variations in thresholds, which are intended to maintain the word masking probability within the range of 0 to 1.
-However, using too small a threshold reduces the differences in text masking probabilities; therefore, we do not recommend using an excessively low threshold.
-Table 17: Comparison of CLIPF pre-trained with token masking and word masking on ImageNet-1K for zero-shot classification.
+Table 19: We pre-train CLIPF on the CC12M dataset across different thresholds in Equation 2.
 The models are pre-trained on CC12M [3] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
@@
 frequency-word 49 8 36.6 39.3
-C.2 Word and Token Analysis
+Table 20: Comparison of CLIPF pre-trained with token masking and word masking on ImageNet-1K for zero-shot classification.
+The models are pre-trained on CC12M [3] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
+Model Masking Time (minutes)
+FLIP trancation 39.81
+random 48.43
+CLIPA block 40.81
+syntax 217.76
+CLIPF frequency 84.64
+Table 21: Processing times for the CC12M dataset by using different masking strategies.
+D.1 Threshold Analysis
+To investigate the impact of the threshold $t$ in Equation 2 on model performance, we pre-trained models with varying thresholds.
+As shown in Tab.
+19, thresholds of 1e51e-5, 1e61e-6, and 1e71e-7 yield comparable results, all outperforming the other text masking strategies.
+This indicates that CLIPF is relatively insensitive to small variations in thresholds, which are intended to maintain the word masking probability within the range of 0 to 1.
+However, using too small a threshold reduces the differences in text masking probabilities; therefore, we do not recommend using an excessively low threshold.
+D.2 Word and Token Analysis
 Since Open_CLIP encodes text using byte-pair encoding (BPE) [15], some words are represented by multiple tokens.
 Therefore, in this experiment, we calculate the masking probability based on token frequency rather than whole words.
-Table 17 compares the performance of both approaches.
+Tab.
+20 compares the performance of both approaches.
 Although token frequency masking may disrupt the original word structure and thus achieve lower performance compared to word frequency masking, it still slightly outperforms other text masking strategies.
+Appendix E Computational overhead
+We carried out a simple measurement of how long it takes to tokenize CC12M using one thread.
+As shown in  Tab.
+21, syntax masking required 2.5 times the number of minutes of CLIPF (218 vs.
+85) and about 5.4 times that of truncation, random, and block masking.